### v1

# Bias-Corrected Data Synthesis for Imbalanced Learning

Abstract
Imbalanced data, where the positive samples represent only a small proportion compared to the negative samples, makes it challenging for classification problems to balance the false positive and false negative rates. A common approach to addressing the challenge involves generating synthetic data for the minority group and then training classification models with both observed and synthetic data. However, since the synthetic data depends on the observed data and fails to replicate the original data distribution accurately, prediction accuracy is reduced when the synthetic data is naively treated as the true data. In this paper, we address the bias introduced by synthetic data and provide consistent estimators for this bias by borrowing information from the majority group. We propose a bias correction procedure to mitigate the adverse effects of synthetic data, enhancing prediction accuracy while avoiding overfitting. This procedure is extended to broader scenarios with imbalanced data, such as imbalanced multi-task learning and causal inference. Theoretical properties, including bounds on bias estimation errors and improvements in prediction accuracy, are provided. Simulation results and data analysis on handwritten digit datasets demonstrate the effectiveness of our method.
Keywords: Bias correction; imbalanced classification; oversampling; prediction accuracy; synthetic data.
1 Introduction
1.1 Background
Imbalanced classification is a fundamental challenge in modern machine learning, arising when the number of observations in one class significantly exceeds that in another class. This issue is prevalent in diverse applications, including detecting rare diseases in medical diagnosis (Rajkomar et al., 2019; Faviez et al., 2020), fraud detection (Subudhi and Panigrahi, 2018), anomaly detection in industrial systems (Kong et al., 2020), and cybersecurity (Sarker, 2019). Traditional classification algorithms often perform poorly under such an imbalance, as they tend to be biased towards the majority class, leading to suboptimal sensitivity and an increasing risk of overlooking critical minority instances.
A common strategy for addressing such challenges in imbalanced classification is data augmentation, which aims to rebalance the samples in different classes by artificially modifying or expanding the training dataset. Resampling-based approaches include undersampling – removing samples from the majority class, and oversampling – expanding the minority class. Undersampling techniques, such as Tomek’s links (Tomek, 1976) and cluster centroid (Lemaître et al., 2017), often suffer from information loss due to discarding potentially informative majority samples. In contrast, oversampling is typically preferred, and various methods have been proposed to enrich the minority class. The reweighting procedure, which assigns higher weights to the minority samples, is equivalent to oversampling by replicating the minority samples. While bootstrap (Efron and Tibshirani, 1994) is a widely used resampling method in statistics, its naive application in oversampling may be sensitive to outliers and may introduce variance inflation. Among oversampling methods, the Synthetic Minority Oversampling TEchnique (SMOTE, Chawla et al. (2002)) has been especially influential. SMOTE generates synthetic samples by interpolating between minority samples and has inspired numerous variants, such as Borderline-SMOTE (Han et al., 2005), ADASYN (He et al., 2008), and safe-level-SMOTE (Bunkhumpornpat et al., 2009), which aim to better capture the geometry of the data and concentrate synthetic sample generation near the decision boundary where classification is challenging. For a comprehensive review of resampling techniques in imbalanced settings, see Mohammed et al. (2020).
Beyond empirical success, recent theoretical studies have examined the statistical properties of synthetic procedures and their impact on classification risk. For example, Elreedy et al. (2024) and Sakho et al. (2024) separately derive the probability distribution of SMOTE-generated synthetic samples, with the latter further proving that the synthetic density function vanishes near the boundary of the minority support. Another widely used augmentation method is Mixup (Zhang et al., 2017), which generates new samples by convex combinations of covariates and their labels. Theoretical results for Mixup include robustness against adversarial attacks and improved generalization by reducing overfitting (Zhang et al., 2020), as well as conditions under which Mixup helps reduce calibration errors (Zhang et al., 2022; Naeini et al., 2015).
Since synthetic samples are typically highly dependent on the original training data, it is crucial to carefully handle such a dependence structure. Tian and Shen (2025) propose a partition-based framework in which one subset of data is used to generate synthetic samples, and the other independent subset is used for training. Nevertheless, a fundamental question remains unresolved: Under what conditions do synthetic procedures improve classification, and how can their potential adverse effects be avoided?
1.2 Our Contribution
We summarize our contributions as follows:
Bias-corrected synthetic data augmentation for imbalanced classification. We develop a bias correction methodology that effectively estimates and adjusts for the discrepancy between the synthetic distribution and the true distribution. By borrowing information from the majority class, our procedure builds a bridge between the observed data and the otherwise unobservable bias in the minority class. Since the minority bias induced by synthetic data is non-negligible, our procedure effectively reduces the bias by an explicit correction term. Theoretically, this bias correction procedure results in improved performance for suboptimal synthetic generators, as confirmed by both simulations and data analysis.
Theoretical guarantees and error bounds. We provide non-asymptotic error bounds for estimators based on raw data, synthetic augmentation, and bias correction methodology. These results identify the regimes where bias correction yields substantial improvement and clarify the trade-offs between variance reduction and bias inflation under different levels of imbalance. Our theoretical results answer the questions of when synthetic augmentation alone suffices and when bias correction is indispensable.
Unified framework with practical validation. We design a general framework that integrates bias correction with diverse synthetic generators, including Gaussian mixture, perturbed sampling, and SMOTE. Through extensive simulations and real-world data analysis, we demonstrate that the proposed method consistently enhances both predictive accuracy and parameter estimation, offering robustness across different imbalance ratios and model architectures.
2 Methodology
We begin by introducing the setting for binary classification with imbalanced data. Suppose that the training data consist of $n$ independent and identically distributed (i.i.d.) samples $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, where $\bm{X}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional covariate vector and $Y_{i}\in\{0,1\}$ represents the class label. Assume that $Y_{i}$’s follow a Bernoulli distribution with $\pi_{1}=\mathbb{P}(Y_{i}=1)$ and $\pi_{0}=\mathbb{P}(Y_{i}=0)$. In the imbalanced setting, we assume $0<\pi_{1}\leq 1/2$, so that the class $Y=1$ is underrepresented. For convenience, we refer to $Y=1$ as the minority class and $Y=0$ as the majority class. Let $n_{1}=\sum_{i=1}^{n}Y_{i}$ and $n_{0}=n-n_{1}$ denote the respective sample sizes, with $n_{1}\ll n_{0}$ with high probability. Without loss of generality, we assume that the samples are ordered such that $Y_{1}=\dots=Y_{n_{1}}=1$ and $Y_{n_{1}+1}=\dots=Y_{n}=0$. We also assume that $\bm{X}\mid(Y=1)\sim\mathcal{P}_{1}$ and $\bm{X}\mid(Y=0)\sim\mathcal{P}_{0}$, where $\mathcal{P}_{1},\mathcal{P}_{0}$ represent the class-conditional distributions.
2.1 Bias Correction with Synthetic Data
Our goal is to build a model to efficiently predict $Y_{n+1}$ based on the new covariate vector $\bm{X}_{n+1}$. For probability prediction function $f\in\mathcal{F}$, where $\mathcal{F}=\{f:\mathbb{R}^{d}\to[0,1]\}$, denote the corresponding loss function as a binary cross-entropy loss, for example, $\ell_{f}(\bm{X},Y)=-Y\log f(\bm{X})-(1-Y)\log(1-f(\bm{X}))$. With the raw data $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, we can just train the prediction function by minimizing the empirical loss function $L^{\rm raw}$:
$L^{\rm raw}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i}).$ (2.1)
In the case that $n_{1}\ll n_{0}$, a trivial guess that all samples are from the majority group will result in accuracy as high as $n_{0}/n$, which is close to $1$, but it does not provide information from the data. To deal with this problem, an intuitive way is to make the data balanced by adding synthetic data samples to the minority group. Assume that we have $\tilde{n}_{1}$ synthetic samples for the minority group: $(\tilde{\bm{X}}^{(1)}_{i},\tilde{Y}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$, where $\tilde{Y}_{i}^{(1)}=1$ for $i=1,\ldots,\tilde{n}_{1}$. By equally treating the synthetic and raw samples, we can run the algorithm by minimizing the synthetic-augmented loss function $L^{\rm syn}$:
$L^{\rm syn}(f)=\frac{1}{n+\tilde{n}_{1}}\bigg(\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})\bigg).$ (2.2)
See Figure 1 for the illustration of the imbalanced learning based on raw data and synthetic augmentation. By introducing $\tilde{n}_{1}$ synthetic samples from the minority group, $L^{\rm syn}$ is a loss function from a “balanced” dataset, especially compared with $L^{\rm raw}$. While $L^{\rm syn}$ helps improve the prediction accuracy, a concern arises when the synthetic data fails to exactly recover the distribution of the minority group $\mathcal{P}_{1}$. This will cause a bias between the loss functions of data from the true distribution $\mathcal{P}_{1}$ to the synthetic distribution $\tilde{\mathcal{P}}_{1}$ as follows,
$\Delta_{1}=\mathbb{E}_{\bm{X}\sim\mathcal{P}_{1}}\{\ell_{f}(\bm{X},1)\}-\mathbb{E}_{\tilde{\bm{X}}\sim\tilde{\mathcal{P}}_{1}}\{\ell_{f}(\tilde{\bm{X}},1)\}.$ (2.3)
Figure 1: A pictorial illustration of imbalanced learning based on raw data, synthetic augmentation and bias correction.
Consider $n_{1}^{*}$ which satisfies $n_{1}+n_{1}^{*}\approx n_{0}$. Imagine that we have $n_{1}^{*}$ unobserved samples from the minority group $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ where $Y_{i}^{*}=1$ and $\bm{X}_{i}^{*}\mid(Y_{i}^{*}=1)\sim\mathcal{P}_{1}$. By introducing the unobserved minority samples, the total dataset $(\bm{X}_{i},Y_{i})_{i=1}^{n}$ and $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ is roughly balanced. Denote the sample bias caused by the synthetic data as
$\hat{\Delta}_{1}=\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)}).$ (2.4)
Suppose we have the balanced data including observed dataset $(\bm{X}_{i},Y_{i})_{i=1}^{n}$ and unobserved dataset $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$. The empirical loss function for the balanced dataset is
$L^{\rm bal}(f)=$ $\frac{1}{n+n_{1}^{*}}\left\{\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})\right\}$
$=$ $\frac{1}{n+n_{1}^{*}}\left\{\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+n_{1}^{*}\cdot\left(\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\hat{\Delta}_{1}\right)\right\}.$ (2.5)
Since $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ are not observable, it is impossible to calculate $\hat{\Delta}_{1}$ with the observed data. We can estimate the bias from the available data in the majority group as follows. First, randomly partition the majority indices into generation subgroup $\mathcal{S}_{0g}$ and correction subgroup $\mathcal{S}_{0c}$ with corresponding sizes $n_{0g}$ and $n_{0c}$, respectively. Next, using samples from the generation subgroup $(\bm{X}_{i},Y_{i})_{i\in\mathcal{S}_{0g}}$, generate $\tilde{n}_{0}$ synthetic samples $(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ by the same synthetic generator, where $\tilde{Y}_{i}^{(0)}=0$ for $i=1,\ldots,\tilde{n}_{0}$. Then consider the population bias of the loss function from the true majority distribution $\mathcal{P}_{0}$ to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ by
$\Delta_{0}=\mathbb{E}_{\bm{X}\sim\mathcal{P}_{0}}\{\ell_{f}(\bm{X},0)\}-\mathbb{E}_{\tilde{\bm{X}}\sim\tilde{\mathcal{P}}_{0}}\{\ell_{f}(\tilde{\bm{X}},0)\}.$ (2.6)
Finally, obtain the sample loss bias for the majority group using the majority synthetic samples and the correction subsamples by
$\hat{\Delta}_{0}=\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},Y_{i})-\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)}).$ (2.7)
Note that all elements for calculating $\hat{\Delta}_{0}$ are available and $\hat{\Delta}_{0}$ is constructed the same way as $\hat{\Delta}_{1}$. Suppose that the transformation from $\mathcal{P}_{0}$ to $\mathcal{P}_{1}$ is captured by a measurable function $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$, and assume that the transformation property is well maintained by the corresponding synthetic distributions. This assumption is formally elaborated in Section 3.1. Given the above properties, it is possible to estimate the unobservable $\hat{\Delta}_{1}$ using the data-driven estimator $\hat{\Delta}_{0}$. Thus, by modifying the balanced loss $L^{\rm bal}(f)$ in (2.5), we propose the following bias-correction loss function $L^{\rm bc}(f)$:
$\begin{split}L^{\rm bc}(f)=\frac{1}{n+\tilde{n}_{1}}\bigg[\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\tilde{n}_{1}\cdot\bigg\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\hat{\Delta}_{0}\bigg\}\bigg].\end{split}$ (2.8)
The majority bias correction term $\hat{\Delta}_{0}$ captures the loss function bias induced by the discrepancy between the true and synthetic distributions from the majority group. Under mild assumptions, $\hat{\Delta}_{0}$ is a good representation of $\hat{\Delta}_{1}$, the loss function bias from the minority group, up to a fixed bias-transfer error and sampling fluctuations, as illustrated in Section 3.1. With this property, the term $\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\hat{\Delta}_{0}$ is regarded as the average loss from the unobserved minority samples after correcting the synthetic bias. Consequently, the bias-corrected loss $L^{\rm bc}$ represents a valid average loss from a roughly balanced dataset. Finally, we can find the prediction function by minimizing the bias-corrected loss:
$\hat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$ (2.9)
The process of bias correction for imbalanced classification is summarized in Algorithm 1. See Figure 1 for an illustration.
Algorithm 1 Bias Correction for Imbalanced Classification
0: Imbalanced data $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, prediction function class $\mathcal{F}$, loss function $\ell_{f}$, synthetic generator $\mathcal{G}$, minority synthetic size $\tilde{n}_{1}$, majority synthetic size $\tilde{n}_{0}$ and generation size $n_{0g}$.
1: Minority augmentation: Generate $\tilde{n}_{1}$ synthetic minority samples
$(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}\leftarrow\mathcal{G}\big((\bm{X}_{i})_{Y_{i}=1}\big).$
2: Partition the majority index set $\mathcal{S}_{0}=\{i:Y_{i}=0\}$ into a generation set $\mathcal{S}_{0g}$ and a correction set $\mathcal{S}_{0c}$ with corresponding sizes $n_{0g}$ and $n_{0c}=n_{0}-n_{0g}$.
3: Generate $\tilde{n}_{0}$ synthetic majority samples
$(\tilde{\bm{X}}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}\leftarrow\mathcal{G}\big((\bm{X}_{i})_{i\in\mathcal{S}_{0g}}\big).$
4: Compute the empirical majority bias $\hat{\Delta}_{0}$ according to Equation (2.7).
5: Form the bias-corrected loss $L^{\rm bc}(f)$ as defined in Equation (2.8).
6: Obtain the predictor by
$\hat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$
6: The prediction function $\hat{f}^{\rm bc}:\mathbb{R}^{d}\to(0,1)$.
2.2 Multi-Task Imbalanced Learning
In this subsection, we focus on applying bias correction techniques to datasets involving multiple related tasks, a scenario commonly addressed by multi-task learning (MTL). MTL is a machine learning paradigm where multiple tasks are learned simultaneously, enabling the model to leverage shared information and learn a common robust representation (Caruana, 1997; Zhang and Yang, 2021). For example, in genomic studies, researchers analyze gene expression data from different regional populations to identify genetic markers for specific diseases such as Alzheimer’s disease (Zhang and Shen, 2011). In this case, each regional population is regarded as a separate learning task. While the goal of identifying Alzheimer’s disease is shared, the genetic and environmental differences between populations lead to unique data distributions. This makes it necessary to utilize a multi-task learning framework to leverage the common structure. However, the number of individuals with the disease is typically smaller than the number of healthy individuals, creating a within-task imbalance problem (Wu et al., 2018; Guo et al., 2025). We apply the bias correction procedure to such imbalanced MTL problems to improve the predictive performance by leveraging information from all tasks.
Consider datasets from $K$ learning tasks and for each task $k=1,\ldots,K$, there are $n_{k}$ samples of covariates $\bm{X}_{ki}\in\mathbb{R}^{d}$ and class labels $Y_{ki}\in\{0,1\}$ for $i=1,\ldots,n_{k}$. Under the imbalanced setting, the class labels are imbalanced within each task such that the marginal probability $\pi_{k1}=\mathbb{P}(Y_{ki}=1)<1/2$. The dependence structure of the class labels on the covariates is captured by the following Bernoulli model:
$\mathbb{P}(Y_{ki}=1\mid\bm{X}_{ki}=\bm{x})=\sigma(\bm{x}^{\top}\bm{B}\bm{\alpha}_{k})\quad\text{ for }i=1,\ldots,n_{k},$ (2.10)
where $\sigma(t)=1/(1+e^{-t})$ denotes the logistic function, $\bm{B}\in\mathbb{R}^{d\times r}$ denotes the shared coefficient matrix across $K$ tasks and $\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K}\in\mathbb{R}^{r}$ are task-specific. Denote $\bm{M}=\bm{B}(\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K})\in\mathbb{R}^{d\times K}$ as the unknown coefficient matrix with ${\rm rank}(\bm{M})=r$. Consider the task-specific coefficient vector $\bm{\beta}_{k}=\bm{B}\bm{\alpha}_{k}\in\mathbb{R}^{d}$ as the $k$th column of $\bm{M}$ for $k=1,\ldots,K$. Our goal is to learn the left singular vector space of the shared matrix $\bm{B}$.
For any $\bm{\beta}\in\mathbb{R}^{d}$, the prediction function is provided by $f(\bm{x})=\sigma(\bm{x}^{\top}\bm{\beta})$ and denote the loss function as
$\ell_{f}(\bm{x},y)=\ell(\bm{x},y;\bm{\beta})=-y\log(\sigma(\bm{x}^{\top}\bm{\beta}))-(1-y)\log(1-\sigma(\bm{x}^{\top}\bm{\beta})).$
For each task, we obtain the estimation $\hat{\bm{\beta}}_{k}^{\rm raw}$ from $(\bm{X}_{ki},Y_{ki})_{i=1}^{n_{k}}$ by minimizing the loss function from the raw data $L^{\rm raw}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta}).$ Consider the minority synthetic samples $(\tilde{\bm{X}}_{ki}^{(1)})_{i=1}^{\tilde{n}_{k1}}$ and majority synthetic samples $(\tilde{\bm{X}}_{ki}^{(0)})_{i=1}^{\tilde{n}_{k0}}$. We can also obtain synthetic and bias-corrected loss functions by
$L^{\rm syn}(\bm{\beta})=$ $\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\sum_{i=1}^{\tilde{n}_{k1}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})\right],$
$L^{\rm bc}(\bm{\beta})=$ $\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\tilde{n}_{k1}\cdot\left\{\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})+\hat{\Delta}_{k0}\right\}\right],$
where $\hat{\Delta}_{k0}$ denotes the bias correction term from the majority samples in the correction set $\mathcal{S}_{kc}$ and majority synthetic samples:
$\hat{\Delta}_{k0}=$ $\frac{1}{|\mathcal{S}_{kc}|}\sum_{i\in\mathcal{S}_{kc}}\ell(\bm{X}_{ki},0;\bm{\beta})-\frac{1}{\tilde{n}_{k0}}\sum_{i=1}^{\tilde{n}_{k0}}\ell(\tilde{\bm{X}}_{ki}^{(0)},0;\bm{\beta}).$
Denote $\hat{\bm{\beta}}_{k}^{\rm syn}$ and $\hat{\bm{\beta}}_{k}^{\rm bc}$ as the minimizers of the synthetic loss $L^{\rm syn}$ and bias-corrected loss $L^{\rm bc}$, respectively.
Next, we consider the estimation of the left singular matrix $\bm{U}$ from any coefficient estimators $(\hat{\bm{\beta}}_{k})_{k=1}^{K}$. First, collect all estimators into $\hat{\bm{M}}=(\hat{\bm{\beta}}_{1},\dots,\hat{\bm{\beta}}_{K})\in\mathbb{R}^{d\times K}$. Next, conduct eigendecomposition of $\hat{\bm{M}}\hat{\bm{M}}^{T}$ such that $\hat{\bm{M}}\hat{\bm{M}}^{T}=\hat{\bm{U}}^{\prime}\hat{\bm{\Lambda}}\hat{\bm{U}}^{\prime\top}$, where $\hat{\bm{U}}^{\prime}\in\mathbb{R}^{d\times d}$ is an orthonormal eigenvector matrix satisfying $(\hat{\bm{U}^{\prime}})^{\top}\hat{\bm{U}}^{\prime}=\bm{I}_{d}$ and $\hat{\bm{\Lambda}}={\rm diag}(\hat{\lambda}_{1},\dots,\hat{\lambda}_{d})$ is a diagonal eigenvalue matrix with decreasing eigenvalues $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{d}$. The rank of the latent embedding matrix is estimated by maximizing the eigenvalue ratio such that $\hat{r}=\arg\max_{1\leq r\leq d_{-}}\hat{\lambda}_{r}/\hat{\lambda}_{r+1}$, where $d_{-}<d$ is a constant to avoid the case of extremely small eigenvalues. Finally, take the first $\hat{r}$ columns of $\hat{\bm{U}}^{\prime}$ to obtain the estimated shared embedding matrix $\hat{\bm{U}}=\hat{\bm{U}}^{\prime}_{1:\hat{r}}$. By substituting the estimator $\hat{\bm{\beta}}_{k}$ by the raw estimator $\hat{\bm{\beta}}_{k}^{\rm raw}$, synthetic estimator $\hat{\bm{\beta}}_{k}^{\rm syn}$ and the bias-corrected estimator $\hat{\bm{\beta}}_{k}^{\rm bc}$, we are able to obtain the corresponding latent embedding matrix estimators $\hat{\bm{U}}^{\rm raw}$, $\hat{\bm{U}}^{\rm syn}$ and $\hat{\bm{U}}^{\rm bc}$.
Suppose we are then provided with samples from a new task $(\bm{X}_{K+1,i},Y_{K+1,i})_{i=1}^{n_{K+1}}$ from the following Bernoulli model with the same shared structure:
$\mathbb{P}(Y_{K+1,i}=1\mid\bm{X}_{K+1,i}=\bm{x})=\sigma(\bm{x}^{\top}\bm{B}\bm{\alpha}_{K+1}).$
The estimated shared embedding matrix $\hat{\bm{U}}$ helps us to estimate the coefficient in a lower dimension $\hat{r}$ rather than $d$. To obtain the estimation, we can first project the covariates into a lower-dimensional embedding subspace by $\hat{\bm{Z}}_{K+1,i}=\hat{\bm{U}}^{\top}\bm{X}_{K+1,i}$ for each $i=1,\dots,n_{K+1}$. Next, obtain $\hat{\bm{\theta}}_{K+1}$ which minimizes the loss function, for example, $L^{\rm raw}(\bm{\theta})$ on the dataset $(\hat{\bm{Z}}_{K+1,i},Y_{K+1,i})_{i=1}^{n_{K+1}}$. Note that the empirical loss function $L^{\rm raw}$ can be replaced by $L^{\rm syn}$ and $L^{\rm bc}$ depending on the imbalance of task $K+1$. Next, project $\hat{\bm{\theta}}_{K+1}$ back to the coefficient space by $\hat{\bm{\beta}}_{K+1}=\hat{\bm{U}}\hat{\bm{\theta}}_{K+1}$ and obtain the prediction function $\hat{f}_{K+1}(\bm{x})=\sigma(\bm{x}^{\top}\hat{\bm{\beta}}_{K+1})$. By substituting $\hat{\bm{U}}$ with the above $\hat{\bm{U}}^{\rm raw}$, $\hat{\bm{U}}^{\rm syn}$ and $\hat{\bm{U}}^{\rm bc}$, we can obtain the corresponding coefficient estimators and prediction function. In Section 3.2, we provide the theoretical results of the coefficient and embedding matrix estimations of the three methods and show the conditions under which the bias correction procedure outperforms the synthetic procedure.
2.3 Average Treatment Effect Estimation
Our proposed methodology has an application to average treatment effect (ATE) estimation, one fundamental problem in causal inference (Rubin, 1974; Lunceford and Davidian, 2004). In this context, imbalanced data often arises when the number of individuals receiving the treatment is significantly smaller than the number of individuals receiving the control. This imbalance makes the estimation of the ATE, more specifically, the expected outcome for the minority group, less reliable due to the limited sample size. This directly degrades the credibility and robustness of the final ATE estimate. Consequently, addressing this imbalanced data is essential for accurate causal inference.
Suppose $Y(1)$ and $Y(0)$ are the potential responses under treatment $Z=1$ and control $Z=0$. The observed response is a function of the potential responses and the treatment indicator:
$Y=ZY(1)+(1-Z)Y(0).$
Then the ATE is defined as
$\tau=\mathbb{E}\{Y(1)\}-\mathbb{E}\{Y(0)\}.$
Consider i.i.d. observations $(\bm{X}_{i},Y_{i},Z_{i})_{i=1}^{n}$, where
$Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0),\qquad i=1,\ldots,n,$
$\bm{X}_{i}\in\mathbb{R}^{d}$ is a vector of covariates, and $Z_{i}\in\{0,1\}$ is the treatment indicator. Suppose there are $n_{1}$ and $n_{0}$ samples from the treatment group and control group, respectively, and define the treated and control covariate indices as $\mathcal{S}_{1}=\{i:Z_{i}=1\}$ and $\mathcal{S}_{0}=\{i:Z_{i}=0\}$. In the imbalanced case where $n_{1}<n_{0}$, we aim to estimate the ATE augmented with synthetic data by the bias correction approach.
Consider the propensity score (Rosenbaum and Rubin, 1983), which is defined as the conditional probability of a sample receiving treatment given the corresponding covariate $\bm{X}_{i}=\bm{x}$: $e^{*}(\bm{x})=\mathbb{P}(Z_{i}=1\mid\bm{X}_{i}=\bm{x}).$ For the propensity score, suppose we have an estimating model denoted as $e(\bm{x})$. Similarly, for the conditional means of the responses given the covariates $\mu_{1}^{*}(\bm{x})=\mathbb{E}\{Y_{i}(1)\mid\bm{X}_{i}=\bm{x}\}$ under treatment and $\mu_{0}^{*}(\bm{x})=\mathbb{E}\{Y_{i}(0)\mid\bm{X}_{i}=\bm{x}\}$ under control, suppose we have the estimating models $\mu_{1}(\bm{x})$ and $\mu_{0}(\bm{x})$, respectively. We consider the augmented inverse propensity weighting (AIPW) estimators (Rubin, 1978; Glynn and Quinn, 2010):
$\hat{\mu}_{1}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}}{e(\bm{X}_{i})}+\mu_{1}(\bm{X}_{i})\right],$ (2.11)
$\hat{\mu}_{0}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}}{1-e(\bm{X}_{i})}+\mu_{0}(\bm{X}_{i})\right],$ (2.12)
$\hat{\tau}^{\rm AIPW}=$ $\hat{\mu}_{1}^{\rm AIPW}-\hat{\mu}_{0}^{\rm AIPW}.$ (2.13)
With the observations $(\bm{X}_{i},Y_{i},Z_{i})_{i=1}^{n}$, we can first fit separate regression models of the responses on the covariates and obtain the coefficient estimators $\hat{\bm{\beta}}_{1}$ and $\hat{\bm{\beta}}_{0}$ under treatment and control, respectively. Let $\hat{\mu}_{1}(\bm{x})=\bm{x}^{\top}\hat{\bm{\beta}}_{1}$ and $\hat{\mu}_{0}(\bm{x})=\bm{x}^{\top}\hat{\bm{\beta}}_{0}$ be the estimated responses under treatment and control, respectively.
Next, we consider the propensity score estimation. Suppose we are interested in a loss function $\ell_{f}$, e.g., logistic loss, for a prediction function $f:\mathbb{R}^{d}\to[0,1]$. With the raw data, we can obtain the propensity score estimation $\hat{f}^{\rm raw}$ by minimizing the empirical loss function $L^{\rm raw}$ in (2.1). Assume that there are $\tilde{n}_{1}$ synthetic minority covariate samples $\tilde{\bm{X}}_{i}^{(1)}$, then we can obtain the synthetic-augmented propensity score estimation $\hat{f}^{\rm syn}$ by minimizing the $L^{\rm syn}$ in (2.2). Partition the control index set into a generation set $\mathcal{S}_{0g}$ and a correction set $\mathcal{S}_{0c}$. Applying the same synthetic generator to obtain $\tilde{n}_{0}$ control synthetic covariates $\tilde{\bm{X}}_{i}^{(0)}$ from the generation set, we can obtain the bias-corrected propensity score estimation $\hat{f}^{\rm bc}$ by minimizing $L^{\rm bc}$ in (2.8).
Plugging the treatment and control models $\hat{\mu}_{1}(\bm{x})$ and $\hat{\mu}_{0}(\bm{x})$ as well as the propensity score estimation $\hat{e}(\bm{x})\in\{\hat{f}^{\rm raw}(\bm{x}),\hat{f}^{\rm syn}(\bm{x}),\hat{f}^{\rm bc}(\bm{x})\}$ into (2.13), the AIPW estimators $\hat{\tau}^{\rm raw},\hat{\tau}^{\rm syn}$ and $\hat{\tau}^{\rm bc}$ can be derived corresponding to the propensity score estimators.
3 Theoretical Properties
3.1 Bias Correction for Risk Functions
In this subsection, we first propose an upper bound for the difference between the minority and majority bias correction terms, $|\hat{\Delta}_{1}-\hat{\Delta}_{0}|$. This result provides a theoretical guarantee for the construction of the bias-corrected loss function $L^{\rm bc}$ in (2.8). We then derive a lower bound for the minority bias correction term $\hat{\Delta}_{1}$ of SMOTE, thereby illustrating that treating synthetic samples as real data may introduce bias. Finally, we present the properties of the bias-corrected predictor with respect to the balanced population risk function, demonstrating that the proposed procedure effectively leverages synthetic samples while ensuring strong performance on imbalanced data.
Assumption 1.
Consider the support of covariates $\mathcal{X}\subseteq\mathbb{R}^{d}$. Suppose the following conditions are satisfied:
  1. (A1)
    Distribution transformation. There exists a measurable function $T:\mathcal{X}\to\mathcal{X}$ and a constant $\varepsilon_{T}$ such that
    $\mathcal{P}_{1}=(\mathcal{P}_{0})_{\#T},\quad\mathcal{W}_{1}(\tilde{\mathcal{P}}_{1},(\tilde{\mathcal{P}}_{0})_{\#T})\leq\varepsilon_{T},$
    where $(\cdot)_{\#T}$ denotes the distribution transformation pushforward by $T$. For instance, if $\bm{X}\sim\mathcal{P}_{0}$, then $T(\bm{X})\sim(\mathcal{P}_{0})_{\#T}$.
  2. (A2)
    Lipschitz smoothness of the transformed loss. There exist constants $L_{\ell},L_{T}>0$ such that for all $\bm{x}_{1},\bm{x}_{2}\in\mathcal{X}$ and $y\in\{0,1\}$,
    $|\ell_{f}(T(\bm{x}_{1}),y)-\ell_{f}(T(\bm{x}_{2}),y)|\leq$ $L_{T}\|\bm{x}_{1}-\bm{x}_{2}\|_{2},$
    $|\ell_{f}(\bm{x}_{1},y)-\ell_{f}(\bm{x}_{2},y)|\leq$ $L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|_{2}.$
  3. (A3)
    Transformation bound for the loss function. There exists $\varepsilon_{h}$ such that
    $|\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)|\leq\varepsilon_{h},\quad\text{ and }\quad|\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\bm{X},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\bm{X},0)|\leq\varepsilon_{h}.$
By Assumption 1 (A1), $T$ represents the transformation from the majority distribution to the minority distribution: if $\bm{X}\sim\mathcal{P}_{0}$, then $T(\bm{X})\sim\mathcal{P}_{1}$. For synthetic distributions, the pair $(\tilde{\mathcal{P}}_{0},\tilde{\mathcal{P}}_{1})$ approximately preserves the same relation: $\tilde{\mathcal{P}}_{1}$ is close in $\mathcal{W}_{1}$ distance to $(\tilde{\mathcal{P}}_{0})_{\#T}$ up to $\varepsilon_{T}$. This ensures that the synthetic generator retains the transformation structure. (A2) ensures the Lipschitz continuity of the loss and the transformed loss. Small covariate perturbations change the loss after transformation $T$ by at most $L_{T}$ times the perturbation size. The stability guarantees that small input fluctuations do not lead to extremely large changes in loss. (A3) further bounds the discrepancy between the minority loss and the majority loss. This condition requires that the loss function $\ell_{f}$ is roughly symmetric in expectation under the majority distribution $\mathcal{P}_{0}$ and the minority distribution $\mathcal{P}_{1}$ up to an error bound $\varepsilon_{h}$. This property also extends to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ and its transformation $(\tilde{\mathcal{P}}_{0})_{\#T}$. It ensures that the transformation does not introduce excessive bias into loss evaluation across the two distributions. Under these assumptions, we now establish a high-probability upper bound for $|\hat{\Delta}_{1}-\hat{\Delta}_{0}|$.
Proposition 1.
Suppose Assumption 1 holds and there exists $0<c_{1}<c_{2}<1/2$ such that $c_{1}\leq n_{1}/n\leq c_{2}$. Then for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}-\hat{\Delta}_{0}|\leq 2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}+\sqrt{\frac{\log(8/\alpha)}{2}}\left(\frac{1}{\sqrt{n_{1}^{*}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right).$
Specifically, when $n_{1}^{*}=O(n_{0}-n_{1})$, $\tilde{n}_{1}=n_{1}^{*}=O(n_{0}-n_{1})$, $n_{0c}=n_{1}^{*}=O(n_{0}-n_{1})$ and $\tilde{n}_{0}=n_{1}^{*}=O(n_{0}-n_{1})$, the result can be written as follows: For any $\alpha>0$ and some constant $C>0$, with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}-\hat{\Delta}_{0}|\leq 2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}+C\sqrt{\frac{\log(8/\alpha)}{n_{0}-n_{1}}}.$
Proposition 1 provides an upper bound for the difference between the majority and minority groups. This upper bound does not directly depend on how accurately the synthetic distribution $\tilde{\mathcal{P}}_{0}$ recovers the true distribution $\mathcal{P}_{0}$. Instead, the discrepancy is controlled through three components: the loss gap in expectation $\varepsilon_{h}$, the transformation approximation error $\varepsilon_{T}$, and the perturbation fluctuation in the order of $1/\sqrt{n}$. This implies that even when the synthetic generator produces samples that poorly approximate the true minority distribution, the bias correction step is nevertheless able to keep the additional error within a well-defined bound. In practice, this means that the procedure remains stable and effective even for poorly performed synthetic generators, ensuring the reliability of the bias-corrected estimator.
A key challenge with synthetic oversampling methods such as SMOTE is that the synthetic distribution $\tilde{\mathcal{P}}_{1}$ of the minority class does not perfectly match the true distribution $\mathcal{P}_{1}$. This discrepancy inevitably introduces a bias in the empirical risk. Here we show two complementary results: (i) SMOTE introduces a non-negligible bias in the minority class, for which we establish a population and empirical lower bound; and (ii) by applying a bias-correction procedure, we can upper bound the corrected error in terms of the distribution discrepancy between $\mathcal{P}_{1},\mathcal{P}_{0}$ and their synthetic counterparts.
Assumption 2.
  1. (A1)
    The loss $\ell_{f}(\bm{x},y)$ is Lipschitz in $\bm{x}$ with constant $L_{\ell}$ uniformly over $y\in\{0,1\}$, i.e., $|\ell_{f}(\bm{x}_{1},y)-\ell_{f}(\bm{x}_{2},y)|\leq L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|$ for any $\bm{x}_{1},\bm{x}_{2}\in\mathbb{R}^{d}$ and $y\in\{0,1\}$.
  2. (A2)
    $\mathcal{P}_{1}$ is supported on a bounded set $B(0,R)\subseteq\mathbb{R}^{d}$ and has a density $f_{1}$ satisfying $0<C_{1}\leq f_{1}(\bm{x})\leq C_{2}<\infty$ for all $\bm{x}\in{\rm supp}(\mathcal{P}_{1})$.
  3. (A3)
    There exists a constant $C_{3}>0$ such that for any $\bm{x}_{1}\in{\rm supp}(\mathcal{P}_{1})$, any of its $K$ nearest neighbors $\bm{x}_{2}$, and any $u\in[0,1]$,
    $|\ell_{f}(\bm{x}_{1}+u(\bm{x}_{2}-\bm{x}_{1}),1)-\ell_{f}(\bm{x}_{1},1)|\geq C_{3}u\|\bm{x}_{2}-\bm{x}_{1}\|_{2}.$
Assumption 2 (A1) guarantees that the loss function $\ell_{f}$ is Lipschitz continuous with respect to $\bm{x}$ uniformly over $y\in\{0,1\}$. Assumption 2 (A2) ensures that the minority distribution has bounded support and a bounded density function on its support. Assumption 2 (A3) holds, for example, if $\ell_{f}(\cdot,1)$ is differentiable on ${\rm supp}(\mathcal{P}_{1})$ and its gradient satisfies $(\bm{x}_{2}-\bm{x}_{1})^{\top}\nabla\ell_{f}(\bm{z},1)\geq C_{3}\|\bm{x}_{2}-\bm{x}_{1}\|_{2}$ for any $\bm{z}$ on the segment between $\bm{x}_{1}$ and $\bm{x}_{2}$.
The following theorem shows that when the minority synthetic samples are generated by SMOTE, the induced bias cannot vanish too quickly. Specifically, there is a lower bound that scales with $(K/n_{1})^{1/d}$, reflecting the discrepancy between $\mathcal{P}_{1}$ and $\tilde{\mathcal{P}}_{1}$.
Proposition 2.
Suppose Assumption 2 holds, and the synthetic minority samples are i.i.d. generated from $\tilde{\mathcal{P}}_{1}$ via SMOTE with parameter $K$. Then there exists a constant $c_{1}>0$, depending on $(d,C_{1},C_{2},C_{3},R)$, such that for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}|\geq c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}-\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}-\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}.$
Proposition 2 demonstrates that the bias introduced into the loss function by the synthetic samples is statistically non-negligible. This bias prevents the synthetic augmented loss function (2.2) from serving as a close substitution for the balanced loss function (2.5). This discrepancy between these loss functions leads to a noticeable difference between their corresponding minimizers, which will potentially reduce the performance of the trained classifier.
Beyond quantifying the bias induced by synthetic sampling, an important question is how such bias affects the learning procedure itself. The natural target is the population balanced risk, which is defined as
$L^{*}(f)=\frac{1}{2}\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]+\frac{1}{2}\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)].$ (3.1)
The population balanced risk function $L^{*}$ represents the optimal case where the two classes are balanced with equal probability, which eliminates the effects of overfitting from imbalanced data. Denote the corresponding population balanced risk minimizer as
$f^{*}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{*}(f),$ (3.2)
where $\mathcal{F}=\{f:\mathcal{X}\to(0,1)\}$ represents the prediction function class. In practice, however, one does not have access to the population balanced loss $L^{*}$, but instead minimizes an empirical loss. With synthetic samples generated and the bias correction procedure, we have the empirical minimizer of the bias-corrected loss function (2.8):
$\hat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$
Next, we investigate the following question: how close is the empirical bias-corrected minimizer $\hat{f}^{\rm bc}$ to the population balanced minimizer $f^{*}$?
To answer this question, we consider a uniform assumption, which requires that the gap between the bias terms for the minority and majority groups remains controlled. Suppose that Assumption 1 is satisfied for all prediction functions $f\in\mathcal{F}$. The basic idea indicates that the bias observed in the minority group can be transferred to the majority group up to a controlled error $\varepsilon_{\rm BT}$, where
$\varepsilon_{\rm BT}=2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}.$
Intuitively, if $\varepsilon_{\rm BT}$ is small, then correcting for the minority bias using information from the majority group is reliable.
The following theorem then provides an upper bound on the excess population risk of the bias-corrected minimizer.
Theorem 3.1.
Suppose Assumption 1 is satisfied for all $f\in\mathcal{F}$ and $\tilde{n}_{1}/(n_{0}-n_{1})\to 1$. Then for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$L^{*}(\hat{f}^{\rm bc})-L^{*}(f^{*})\leq$ $\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\varepsilon_{\rm BT}$
$+\sqrt{\frac{\log(10/\alpha)}{2}}\left\{\frac{\pi_{1}/\pi_{0}}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\left(\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}\right)\right\}.$
The bound in Theorem 3.1 shows that the excess population risk of the bias-corrected estimator is controlled by two types of terms:
  1. (i)
    A bias transfer term, $\varepsilon_{\rm BT}$, which measures the worst-case mismatch of the bias correction from the majority to the minority group.
  2. (ii)
    Complexity terms with the order of inverse square root of sample sizes, which quantify the statistical fluctuations from randomness of finite samples in each component dataset.
Thus, bias correction ensures that even though synthetic oversampling introduces a nontrivial distributional bias, the bias-corrected empirical risk minimizer achieves population risk close to the optimal $f^{*}$, up to statistical and transfer errors.
3.2 Bias Correction for Multi-Task Learning
In transfer learning with multiple sources, samples in each source $k=1,\ldots,K$ are independently drawn from the corresponding distribution and regression parameter $\bm{\beta}_{k}$. We study how synthetic augmentation and bias correction affect the accuracy of parameter estimation across tasks, and how these errors accumulate to the shared low-rank structure.
For each source $k$, we observe $n_{k}$ i.i.d. samples $(\bm{X}_{ki},Y_{ki})_{i=1}^{n_{k}}$ with $Y_{ki}\in\{0,1\}$ following the logistic model:
$\mathbb{P}(Y_{ki}=1\mid\bm{X}_{ki}=\bm{x})=\sigma(\bm{x}^{\top}\bm{\beta}_{k}),\quad\text{ where }\sigma(t)=1/(1+\exp(-t)).$
We impose the following assumptions.
Assumption 3.
For each source index $k\in\{1,\ldots,K\}$:
  1. (A1)
    There exists $R>0$ such that $\|\bm{X}_{ki}\|_{2}\leq R$ almost surely.
  2. (A2)
    The Fisher information at $\bm{\beta}_{k}$ is uniformly positive definite with
    $\underline{\kappa}_{k}\bm{I}_{d}\preceq\bm{H}_{k}:=\mathbb{E}[\sigma^{\prime}(\bm{X}_{k1}^{\top}\bm{\beta}_{k})\bm{X}_{k1}\bm{X}_{k1}^{\top}]\preceq\bar{\kappa}_{k}\bm{I}_{d}\quad\text{ for some constants }0<\underline{\kappa}_{k}\leq\bar{\kappa}_{k}<\infty.$
  3. (A3)
    The synthetic samples $\tilde{\bm{X}}^{(1)}_{ki}$ and $\tilde{\bm{X}}_{ki}^{(0)}$ are generated by a fixed mechanism i.i.d. conditional on the training data.
  4. (A4)
    For $y\in\{0,1\}$, the gradient of the loss function $g_{k,y}(\bm{x})=\nabla_{\bm{\beta}}\ell(\bm{x},y;\bm{\beta}_{k})$ is $L_{k}$-Lipschitz on ${\rm supp}(\mathcal{P}_{k,t})\cup{\rm supp}(\tilde{\mathcal{P}}_{k,t})$, where $\ell$ represents the logistic loss function.
The following theorem establishes nonasymptotic bounds for three types of estimations: the raw MLE $\hat{\bm{\beta}}_{k}$, the synthetic augmented estimator $\tilde{\bm{\beta}}_{k}$, and the bias-corrected estimator $\tilde{\bm{\beta}}_{k}^{\rm bc}$. It also quantifies how these parameter errors propagate to the estimation of the shared low-rank structure.
Theorem 3.2.
Under Assumption 3, for any $\alpha\in(0,1)$, there exist constants $C_{1},C_{2},C_{3}>0$ such that, with probability at least $1-\alpha$, the following properties hold simultaneously for each source $k$:
${\rm(i)}\quad$ $C_{1}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\max}(\bm{H}_{k})}\right)\leq\|\hat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}\leq C_{1}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\min}(\bm{H}_{k})}\right),$
${\rm(ii)}\quad$ $\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\leq\frac{L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})+C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
$\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\geq\frac{L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})-C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
${\rm(iii)}\quad$ $\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}\left(\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{{\rm BT}}+\varepsilon_{{\rm sampling},k}\right),$
where
$\varepsilon_{{\rm sampling},k}=$ $C_{3}R\sqrt{\log(10d/\alpha)}\left\{\frac{\pi_{1}/(2\pi_{0})}{\sqrt{n_{k1}}}+\frac{1}{2\sqrt{n_{k0}}}+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\left(\frac{1}{\sqrt{n_{k0,c}}}+\frac{1}{\sqrt{\tilde{n}_{k0}}}+\frac{1}{\sqrt{\tilde{n}_{k1}}}\right)\right\}.$
Furthermore, consider the $d\times K$ true and estimated matrices as $\bm{M}=[\bm{\beta}_{1},\cdots,\bm{\beta}_{K}]$, $\hat{\bm{M}}^{\rm raw}=[\hat{\bm{\beta}}_{1}^{\rm raw},\cdots,\hat{\bm{\beta}}_{K}^{\rm raw}]$, $\hat{\bm{M}}^{\rm syn}=[\hat{\bm{\beta}}_{1}^{\rm syn},\cdots,\hat{\bm{\beta}}_{K}^{\rm syn}]$ and $\hat{\bm{M}}^{\rm bc}=[\hat{\bm{\beta}}_{1}^{(\rm bc)},\cdots,\hat{\bm{\beta}}_{K}^{(\rm bc)}]$. Let $\bm{U}$ be the matrix of leading $r$ left singular vectors of $\bm{M}$ and $D=\sigma_{r}(\bm{M})-\sigma_{r+1}(\bm{M})$ the spectral gap. Then, with the same probability,
${\rm(iv)}\quad$ $\|\sin\Theta(\hat{\bm{U}}^{\rm raw},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\hat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(v)}\quad$ $\|\sin\Theta(\hat{\bm{U}}^{\rm syn},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(vi)}\quad$ $\|\sin\Theta(\hat{\bm{U}}^{\rm bc},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2}.$
The bounds illustrate the trade-offs among three estimators:
  1. (i)
    The raw MLE $\hat{\bm{\beta}}_{k}^{\rm raw}$ converges at the standard parametric rate $n_{k}^{-1/2}$.
  2. (ii)
    The synthetic estimator $\hat{\bm{\beta}}^{\rm syn}$ inherits an additional bias, which reflects the distributional mismatch between $\mathcal{P}_{k,0}$ and $\tilde{\mathcal{P}}_{k,0}$.
  3. (iii)
    The bias-corrected estimator $\hat{\bm{\beta}}_{k}^{\rm bc}$ removes the bias at the cost of extra sampling fluctuations and the residual class-difference bias $\varepsilon_{\rm BT}$.
Finally, the subspace error bounds (iv)-(vi) show how these parameter errors accumulate in estimating the shared low-rank structure, with stability governed by the spectral gap $D$.
Remark 1.
When
$\varepsilon_{\rm BT}\leq\frac{2\pi_{0}}{\pi_{0}-\pi_{1}}\cdot\frac{\underline{\kappa}_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0}),$ (3.3)
the bias correction parameter estimator $\hat{\bm{\beta}}_{k}^{\rm bc}$ has smaller errors than the synthetic augmented parameter estimator $\hat{\bm{\beta}}^{\rm syn}$. In contrast, when (3.3) does not hold, the advantage of the bias correction procedure is not guaranteed. In addition, when $\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})=O_{p}(n_{k}^{-1/2})$, the lower bound for the synthetic-augmented estimator error $\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|$ in Theorem 3.2 becomes negative and thus can be replaced by zero. Theorem 3.2 shows that the bias correction performs well, especially for “bad” synthetic generators.
3.3 Average Treatment Effect Estimation
Let $W=(\bm{X},Y,Z)$ be the observed dataset with covariates $\bm{X}\in\mathbb{R}^{d}$, the treatment/control indicator $Z\in\{0,1\}$ and observed response Y=ZY(1)+(1Z)Y(0)Y=ZY(1)+(1-Z)Y(0). The average treatment effect is $\tau=\mathbb{E}[Y(1)]-\mathbb{E}[Y(0)]$. Denote the conditional treatment and control responses as $\mu_{1}^{*}(\bm{x})=\mathbb{E}[Y(1)\mid\bm{X}=\bm{x}]$ and $\mu_{0}^{*}(\bm{x})=\mathbb{E}[Y(0)\mid\bm{X}=\bm{x}]$, respectively. Let $e^{*}(\bm{x})=\mathbb{P}(Z=1\mid\bm{X}=\bm{x})$ be the propensity score. For any arbitrary functions $\mu_{1}(\cdot)$, $\mu_{0}(\cdot)$ and $e(\cdot)$, define
$\psi(W;\mu_{1},\mu_{0},e)=\left\{\mu_{1}(\bm{X})+\frac{Z(Y-\mu_{1}(\bm{X}))}{e(\bm{X})}\right\}-\left\{\mu_{0}(\bm{X})+\frac{(1-Z)(Y-\mu_{0}(\bm{X}))}{1-e(\bm{X})}\right\}.$
Then the augmented inverse propensity weighting estimator of ATE given $\hat{\mu}_{1}$, $\hat{\mu}_{0}$ and $\hat{e}$ is provided by
[…]
Assumption 4.
  1. (A1)
    Identifiability. $(Y(1),Y(0))$ is independent of $Z$ conditional on $\bm{X}$ and $\eta\leq e^{*}(\bm{X})\leq 1-\eta$ almost surely for some $0<\eta<1$.
  2. (A2)
    Bounded moments. $|Y|\leq M$ and $\|\bm{X}\|_{2}\leq R$ for some constants $M,R>0$.
Theorem 3.3.
Under Assumption 4, denote $r_{a}=(\mathbb{E}|\hat{\mu}_{a}(\bm{X})-\mu_{a}^{*}(\bm{X})|^{2})^{1/2}$ for $a\in\{0,1\}$ and $r_{e}=(\mathbb{E}|\hat{e}(\bm{X})-e^{*}(\bm{X})|^{2})^{1/2}$. There exist constants $C_{0},C_{1}>0$ depending only on $\eta$ and $M$, such that with probability at least $1-\alpha$,
$|\hat{\tau}^{\rm AIPW}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}}{\eta}(r_{1}+r_{0})r_{e}+C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{e}+r_{1}r_{0}\big).$
Corollary 1.
Under Assumptions 3 and 4, suppose the propensity score estimation is obtained from the bias-corrected coefficient estimator $\hat{e}^{\rm bc}(\bm{x})=\sigma(\bm{x}^{\top}\hat{\bm{\beta}}^{\rm bc})$, then the corresponding ATE estimator satisfies
$|\hat{\tau}^{\rm AIPW,bc}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}R}{\eta}(r_{1}+r_{0})r_{\beta}+C_{0}R\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{\beta}+r_{1}r_{0}\big),$
where $r_{\beta}$ represents the error bound for $\hat{\bm{\beta}}^{\rm bc}$:
$r_{\beta}=\frac{1}{\underline{\kappa}_{k}}\left(\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{{\rm BT}}+\varepsilon_{{\rm sampling}}\right).$
Suppose that the treatment and control effect estimations $\hat{\mu}_{1}(\cdot)$ and $\hat{\mu}_{0}(\cdot)$ are obtained from the raw data in the treatment and control groups, respectively. For example, when the model is correctly specified for linear regression, the corresponding errors scale with the sample sizes, i.e., $r_{a}=O(n_{a}^{-1/2})$ for $a\in\{0,1\}$. Then according to Theorem 3.3, how close $\hat{\tau}^{\rm AIPW}$ is to the true ATE $\tau$ depends on the propensity score estimation error $r_{e}$. Since the propensity score estimation is obtained from the coefficient estimation $\hat{\bm{\beta}}$, we can have the following result: when (3.3) holds, the ATE estimation from the bias correction procedure $\hat{e}^{\rm bc}(\bm{X})=\sigma(\bm{X}^{\top}\hat{\bm{\beta}}^{\rm bc})$ has a smaller error than the synthetic augmented estimation from $\hat{e}^{\rm syn}(\bm{X})=\sigma(\bm{X}^{\top}\hat{\bm{\beta}}^{\rm syn})$. In contrast, when (3.3) does not hold, the advantage of the bias correction procedure is not guaranteed.
4 Simulation Studies
4.1 Mean Shift Model
The simulation investigates the performance of imbalanced classification on synthetic augmented data with and without bias correction. We consider binary classification with data generated as follows. First, generate $Y_{i}$ i.i.d. from the Bernoulli distribution with parameter $\pi_{1}$. Second, generate each element of the covariates $\bm{X}_{i}$ from three distributions ($t(2)$, $\mathcal{N}(\cdot,1)$, and ${\rm Logistic}(\cdot,5)$). For the Mean shift model, the minority distribution is transferred from the majority distribution by adding a constant distribution shift vector $\bm{\mu}$. We split the data randomly into the training set, validation set, and test set with probabilities $60\%$, $20\%$ and $20\%$, respectively. Without loss of generality, we reorder the samples such that $Y_{i}=1$ for $i=1,\ldots,n_{1}$ and $Y_{i}=0$ for $i=n_{1}+1,\ldots,n$, where $n_{1}$ denotes the minority sample size.
The synthetic data are generated from the SMOTE algorithm for the first setting. Specifically, the minority synthetic samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$ are generated from all minority covariates $(\bm{X}_{i})_{i=1}^{n_{1}}$. The majority synthetic samples $(\tilde{\bm{X}}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ are generated from a subset of the majority covariates of size $n_{1}$, denoted as $(\bm{X}_{i})_{i=n_{1}+1}^{2n_{1}}$.
Consider the loss function $\ell(\bm{x},y;\bm{\beta})=-y\cdot\bm{x}^{\top}\bm{\beta}+\log(1+\exp(\bm{x}^{\top}\bm{\beta}))$. For the three methods – using raw data, synthetic-augmented data, and synthetic-augmented data with bias correction – we train models by minimizing the respective loss functions given in Equations (2.1), (2.2), and (2.8) for 100 epochs. The entire simulation procedure is repeated over 100 replicates to reduce the impact of randomness. The resulting evaluation metrics for the three methods are summarized in Table 1.
Table 1: Performance metrics (recall, precision, F1-score) evaluated on raw and synthetic-augmented data (with/without bias correction) across varying distributions based on the mean shift model. The results are based on 100 simulations and bold values indicate the top-performing method per metric.
$t(2)$ $\mathcal{N}(\cdot,1)$ $Logistic(\cdot,5)$
Raw SMOTE Bias Corr Raw SMOTE Bias Corr Raw SMOTE Bias Corr
Recall 0.5942 0.7240 0.7317 0.6404 0.7916 0.7962 0.5449 0.5497 0.5553
Precision 0.0059 0.0071 0.0072 0.0065 0.0080 0.0080 0.0055 0.0056 0.0056
$F_{\beta}$ Score 0.5927 0.7221 0.7298 0.6388 0.7896 0.7942 0.5435 0.5484 0.5539
Based on an analysis of 100 simulations across varying data distributions in Table 1, the application of the SMOTE synthetic data generator produced a marked performance improvement over using raw data alone, as measured by recall, precision, and $F_{\beta}$-score. Moreover, incorporating the bias correction technique on top of SMOTE leads to an additional improvement, indicating that the combined approach enhances model robustness and generalization. These results demonstrate that while SMOTE remains an effective foundation for handling data imbalance, integrating bias correction further refines the model’s predictive performance, yielding consistent gains across multiple evaluation metrics.
To further demonstrate the effectiveness of the bias correction method on synthetic data, we conduct simulations using some other synthetic generators that can introduce substantial bias. Although directly using the synthetic data from these generators already improves performance compared with methods relying solely on raw data, applying the bias correction procedure leads to a further and substantial performance gain. The results are summarized in Table 2.
Table 2: Performance metrics (Recall, Precision, F1-score, Jaccard index) evaluated on raw and synthetic data (with/without bias correction) across three synthetic methods (Gaussian mixture, perturbed sampling and biased SMOTE) based on the mean shift model. The results are based on 100 simulations and bold values indicate the top-performing method per metric.
Gaussian-Mixture Perturbed-Sampling Biased-SMOTE
Raw Synthetic Bias Corr Raw Synthetic Bias Corr Raw Synthetic Bias Corr
Recall 0.0968 0.3161 0.3247 0.0980 0.2845 0.3083 0.0940 0.8672 0.9074
Precision 0.4951 0.4982 0.4990 0.5058 0.4969 0.4948 0.4989 0.5031 0.5030
F1 0.1616 0.3782 0.3841 0.1639 0.3501 0.3701 0.1576 0.6363 0.6469
Jaccard 0.0881 0.2367 0.2415 0.0895 0.2158 0.2304 0.0859 0.4670 0.4784
As presented in Table 2, the generation of synthetic data itself provided a significant performance improvement in recall, F1-score, and Jaccard index over models trained solely on raw data, with precision remaining similar. However, in contrast to high-quality generators, these poorer methods introduced substantial bias, which is evidenced by a further marked improvement in performance after the application of a bias correction technique. This demonstrates that for such suboptimal synthetic data, the bias correction procedure is a critically important step that successfully mitigates inherent biases and leads to the best overall model performance.
4.2 Non-linear Classification
To further evaluate the effectiveness and robustness of the proposed bias correction technique under diverse conditions, we conducted a series of controlled simulations using four non-linear classification settings. Each setting represents a distinct geometric relationship between the majority and minority classes, designed to capture a range of challenging data imbalance scenarios. The datasets were constructed with varying degrees of class overlap, non-convexity, and variance, providing a comprehensive test bed for assessing model robustness. For each configuration, synthetic samples were generated using the SMOTE algorithm followed by a bias correction step. The classification performance was evaluated in terms of the $F_{\beta}$ score, averaged over 100 independent runs to ensure statistical reliability.
Figure 2: Distribution and $F_{\beta}$ score for four non-linear classification settings.
Figure 2 illustrates the data distributions and corresponding $F_{\beta}$ scores for four non-linear classification settings. Across all configurations, applying the bias correction technique consistently improves the performance of synthetic data augmentation, indicating its effectiveness in refining the representativeness of SMOTE-generated samples. Notably, in the right two cases, the support of the majority class distribution is clearly non-convex. Since SMOTE generates new samples through convex combinations of existing data points, this result is particularly striking—it demonstrates that the bias correction term can effectively enhance model performance even when the synthetic generator fails to capture the true underlying distribution. Furthermore, in most settings, the minority distributions exhibit large variance, a condition that typically poses a challenge in imbalanced classification. The improved performance under these high-variance conditions further highlights the robustness and adaptability of the bias correction approach in complementing synthetic oversampling methods.
4.3 Sigmoid Bernoulli Model
In this section, a simulation study on the sigmoid Bernoulli model is conducted. We mainly study the behavior for four exponential family distributions (Gaussian distribution, Gumbel distribution, location-scale $t$ distribution and HSD) with a relatively high-quality synthetic generator (SMOTE) and a suboptimal synthetic generator with random sampling with noise (perturbed sampling). Two distributions outside of the exponential family (Laplace distribution and logistic distribution) are also tested on the SMOTE-based synthetic data and give good results, demonstrating the robustness of the proposed method to various distributions. The sample size is set to n=1,000n=1{,}000, and the dimension to $d=10$. The mean squared errors are reported based on 100 simulation runs.
Table 3: Estimation error of parameter $\bm{\beta}$ evaluated on raw and synthetic data (with/without bias correction) across varying distributions based on sigmoid Bernoulli model. The results are based on 100 simulations and bold values indicate the top-performing method.
Gaussian Gumbel Loc-Scale t HSD Laplace Logistic
Raw 2.441 2.426 2.431 2.496 2.401 2.384
SMOTE 2.310 2.311 2.347 2.380 2.330 2.362
Bias Corr 2.299 2.308 2.332 2.369 2.318 2.350
The evaluation of parameter estimation error for $\beta$ under the sigmoid Bernoulli model, shown in Table 3, reveals that employing SMOTE-generated synthetic data consistently enhances estimation accuracy across all tested data distributions when compared with models trained solely on raw data. Moreover, incorporating the bias correction technique yields an additional and notable reduction in estimation error, indicating that the correction effectively compensates for residual bias in the synthetic samples. This improvement highlights that, although SMOTE alone serves as a strong baseline for generating high-quality synthetic data, the bias correction step further refines the fidelity of parameter estimation, leading to more accurate recovery of the true underlying model. Overall, these results demonstrate that the proposed bias correction approach provides a meaningful and reliable performance gain beyond standard synthetic augmentation.
Table 4: Estimation error of parameter $\bm{\beta}$ evaluated on raw and synthetic data (with/without bias correction) with two synthetic methods (Perturbed Sampling and Gaussian Mixture) across varying distributions based on sigmoid Bernoulli model. The results are based on 100 simulations and bold values indicate the top-performing method.
Gaussian Mixture Perturbed Sampling
Raw Synthetic Bias Corr Raw Synthetic Bias Corr
Gumbel 2.459 2.418 2.325 2.458 2.438 2.219
HSD 2.511 2.472 2.440 2.512 2.471 2.297
Logistic 2.349 2.507 2.311 2.347 2.658 2.205
The parameter estimation error for $\beta$ in Table 4 demonstrates that the two relatively low-quality synthetic data generators—perturbed sampling and Gaussian mixture—can, in fact, degrade estimation accuracy, yielding higher errors than those obtained using the raw data alone. This deterioration occurs because these generators introduce systematic bias and distort the underlying data distribution, leading to unreliable parameter estimates. However, applying the bias correction technique effectively eliminates this detrimental effect and not only restores performance to the baseline level but also surpasses the accuracy achieved with raw data. This outcome underscores the robustness and corrective strength of the proposed method: even when the synthetic data generator fails to model the true distribution faithfully, bias correction can compensate for these deficiencies and produce more accurate and stable parameter estimates across diverse data settings.
4.4 Average Treatment Effect Estimation
In this section, we conduct a simulation study to evaluate the performance of three methods for estimating the ATE. We compare the standard estimator applied to the raw data with estimators applied to data augmented by SMOTE with and without bias correction. The covariates $\bm{X}$ are generated from four distributions: t(6), t(4), Logistic, and Laplace. The results are summarized in Figure 3.
Figure 3: Square Root MSE for ATE Estimation of Three Methods across Four Distributions of Covariates $\bm{X}$.
The simulation results, presented in Figure 3, lead to two main conclusions. First, incorporating synthetic data, either through SMOTE or through our proposed bias correction method, improves ATE estimation relative to using raw data alone. Second, the bias correction procedure plays a crucial role, as it substantially reduces estimation error and consistently delivers superior performance across all examined distributional settings. These findings also indicate that the SMOTE generator introduces a non-negligible bias in this specific task, suggesting that it may not be universally effective across all applications.
5 Data Analysis for MNIST Dataset
To evaluate the practical efficacy of our proposed framework, we apply it to the MNIST dataset (LeCun, 1998). We use Perturbed Sampling to generate synthetic data for a binary classification task where digit 1 or 4 is treated as the minority class in a five-digit subset. The results are detailed in Figure 4.
Figure 4: Seven metrics (Recall, Precision, Accuracy, Fowlkes-Mallows score (FM), F1-score, Matthews Correlation Coefficient (MCC), and Jaccard index) for the three methods applied to the MNIST dataset (digits 0-4). Results are shown for binary classification of digit 1 (top panel) and digit 4 (bottom panel), treated as minority classes.
Based on an evaluation in Figure 4, the perturbed sampling method demonstrates a clear performance hierarchy for classifying digits 1 and 4 as minority classes. While maintaining a similar precision score, the generation of synthetic data provides a foundational improvement, yielding superior results across all other metrics compared to the model trained exclusively on the raw, imbalanced data. The significant performance leap observed after applying bias correction reveals the substantial inherent bias introduced by the perturbed sampling technique. This bias correction step is not merely beneficial but critical, as it consistently produces the most accurate and reliable classifications. The results underscore that bias correction is an important step for mitigating distortion and achieving better model performance.
6 Discussions
We propose a novel bias correction algorithm to improve the performance of imbalanced classification when using synthetic data augmentation. Treating synthetic data equally to the true data often introduces a systematic bias because synthetic generators rarely perfectly recover the true data distribution. For many popular techniques like SMOTE, the irreducible bias term between the unobserved true data and synthetic data can hinder model generalization. Our methodology focuses on estimating the bias within the minority group. This is achieved by partitioning the majority group into disjoint generation and correction sets. We first generate majority-class synthetic samples from the generation set and then quantify the bias by comparing the difference between the majority synthetic samples and the samples from the correction set. Theoretical results confirm the soundness of the corrected bias and the effectiveness of the resulting predictor. A key advantage of our approach is that its theoretical guarantees do not directly depend on the discrepancy between the true and synthetic distributions. Consequently, the bias correction approach demonstrates robust performance when using suboptimal synthetic generators. This framework can be extended to other domains, including multi-task learning and causal inference. Simulation studies and an empirical application to MNIST handwritten digit image dataset validate the performance of the bias correction algorithm.
Despite its strengths, the bias correction approach has several limitations: (i) The current theoretical bias term is defined with respect to a one-dimensional loss function. For high-dimensional data, this approach risks losing information about the complex discrepancy between the true and synthetic distributions when compressing the distributional error into a scalar loss bias. (ii) When synthetic generators produce synthetic samples of high validity, the performance gains achieved by the bias correction approach are often marginal compared to simply using the synthetic-augmented data directly. Employing the bias correction in such scenarios leads to an unnecessary cost of time and computational resources. (iii) Although the application to MNIST multi-class problems yields satisfactory results, a formal theoretical derivation supporting the algorithm’s extension to the multi-class setting is not provided in this paper.
Future work will focus on three main areas: First, developing a more comprehensive metric to characterize the discrepancy between the true and synthetic distributions beyond a simple loss function bias and integrating this into the correction framework. Second, establishing a computationally scalable criterion to determine whether the bias correction approach is likely to yield substantial performance gains, thus helping to optimize resources. Finally, providing a rigorous theoretical extension of the bias correction framework to formally support its application to the multi-class classification problem.
Acknowledgements
P. L. and A. R. Z. were partially supported by NIH Grant R01HL169347; Z. M. and A. R. Z. were partially supported by NIH Grant R01HL168940; A. R. Z. was also partially supported by NSF Grant CAREER-2203741. L.Z. was partially supported by NSF Grant CAREER 2340241 and Renaissance Philanthropy “AI for Math” Fund.
Appendix A Proof of main results
A.1 Proof of Proposition 1
Proof.
By the triangle inequality, we have
$|\hat{\Delta}_{1}-\hat{\Delta}_{0}|\leq|\Delta_{1}-\Delta_{0}|+|\hat{\Delta}_{1}-\Delta_{1}|+|\hat{\Delta}_{0}-\Delta_{0}|,$ (A.1)
where
$\Delta_{1}=\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1),\quad\Delta_{0}=\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\tilde{\bm{X}},0).$
Since $\mathcal{P}_{1}=(\mathcal{P}_{0})_{\#T}$ for transformation $T$ by Assumption 1 (A1), $\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)=\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(T(\bm{X}),1)$. Similarly, we also have $\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(T(\tilde{\bm{X}}),1)=\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)$. We can thus rewrite $\Delta_{1}$ as
$\Delta_{1}=\big\{\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(T(\bm{X}),1)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(T(\tilde{\bm{X}}),1)\big\}+\big\{\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\big\}.$
Denote $h(\bm{x}):=\ell_{f}(T(\bm{x}),1)-\ell_{f}(\bm{x},0)$, and we have
$\Delta_{1}-\Delta_{0}=$ $\big\{\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})\big\}+\big\{\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\big\}.$ (A.2)
Under Assumption 1 (A2), $\ell_{f}(\bm{x},1)$ has Lipschitz constant $L_{\ell}>0$. By the Kantorovich-Rubinstein duality and Assumption 1 (A1),
$\big|\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\big|\leq L_{\ell}\cdot\mathcal{W}_{1}((\tilde{\mathcal{P}}_{0})_{\#T},\tilde{\mathcal{P}}_{1})\leq L_{\ell}\cdot\varepsilon_{T}.$ (A.3)
Assumption 1 (A3) guarantees that $h$ is bounded by $\varepsilon_{h}$ in expectation with respect to $\mathcal{P}_{0}$ and $\tilde{\mathcal{P}}_{0}$, then
$\big|\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})\big|\leq 2\varepsilon_{h}.$ (A.4)
Therefore, combining (A.2), (A.3) and (A.4), we have
$|\Delta_{1}-\Delta_{0}|\leq 2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}.$ (A.5)
Since $\ell_{f}\in[0,1]$, and the samples in $\hat{\Delta}_{1}$ and $\hat{\Delta}_{0}$ are independent, by Hoeffding’s inequality, for any $t>0$,
$\mathbb{P}\left(\left|\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)\right|>t\right)\leq$ $2\exp\{-2n_{1}^{*}t^{2}\},$
$\mathbb{P}\left(\left|\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\right|>t\right)\leq$ $2\exp\{-2\tilde{n}_{1}t^{2}\},$
$\mathbb{P}\left(\left|\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},0)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)\right|>t\right)\leq$ $2\exp\{-2n_{0c}t^{2}\},$ (A.6)
$\mathbb{P}\left(\left|\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},0)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\tilde{\bm{X}},0)\right|>t\right)\leq$ $2\exp\{-2\tilde{n}_{0}t^{2}\}.$ (A.7)
Let
$t_{1}=\sqrt{\frac{\log(8/\alpha)}{2n_{1}^{*}}},\quad t_{2}=\sqrt{\frac{\log(8/\alpha)}{2\tilde{n}_{1}}},\quad t_{3}=\sqrt{\frac{\log(8/\alpha)}{2n_{0c}}},\quad t_{4}=\sqrt{\frac{\log(8/\alpha)}{2\tilde{n}_{0}}}.$
A union bound over the four two-sided events yields that with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}-\Delta_{1}|\leq t_{1}+t_{2},\quad|\hat{\Delta}_{0}-\Delta_{0}|\leq t_{3}+t_{4}.$ (A.8)
Therefore, combining (A.1), (A.5) and (A.8), we obtain the result: with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}-\hat{\Delta}_{0}|\leq 2\varepsilon_{h}+L_{\ell}\cdot\varepsilon_{T}+\sqrt{\frac{\log(8/\alpha)}{2}}\left(\frac{1}{\sqrt{n_{1}^{*}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right).$
This completes the proof of Proposition 1. ∎
A.2 Proof of Proposition 2
Proof.
First, we show that
$\mathcal{O}\left((K/n_{1})^{1/d}\right)\leq\mathcal{W}_{2}(\mathcal{P}_{1},\tilde{\mathcal{P}}_{1})\leq\mathcal{O}((K/n_{1})^{1/(3d)}),$ (A.9)
where $\mathcal{W}_{2}(\cdot,\cdot)$ denotes the Wasserstein distance defined as
$\mathcal{W}_{2}(\mathcal{P}_{1},\tilde{\mathcal{P}}_{1})=\inf_{\nu\in\Pi(\mathcal{P}_{1},\tilde{\mathcal{P}}_{1})}\left\{\mathbb{E}_{(\bm{X},\tilde{\bm{X}})\sim\nu}(\|\bm{X}-\tilde{\bm{X}}\|_{2}^{2})\right\}^{1/2}.$
Denote $\tilde{\bm{X}}_{j}$ as the generated synthetic sample with center $X_{i}$ for $i=1,\ldots,n_{1}$ and $j=1,\ldots,\tilde{n}_{1}$. Specifically, denote $\bm{X}_{i(1)},\ldots,\bm{X}_{i(K)}$ as the $K$ nearest neighbors of $\bm{X}_{i}$, then $\tilde{\bm{X}}_{j}$ is generated by
$\tilde{\bm{X}}_{j}=\bm{X}_{i}+U_{i}(\bm{X}_{i(k)}-\bm{X}_{i}),$
where $U_{i}$ is sampled uniformly from $(0,1)$ and $k$ is sampled uniformly from $\{1,2,\ldots,K\}$. Recall that $R$, defined in Assumption 2, represents the upper bound for the covariates, i.e., $\mathbb{P}_{\mathcal{P}_{1}}(\|\bm{X}\|\leq R)=1.$ By Sakho et al. [2024], for any $\gamma\in(0,1/d)$, we have
$\mathbb{P}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2}\geq 12R(K/n_{1})^{\gamma})\leq(K/n_{1})^{2/d-2\gamma}.$
Thus we have
$\mathbb{E}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2})=$ $\mathbb{E}\{\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|\cdot I(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2}<12R(K/n_{1})^{\gamma})\}$
$+\mathbb{E}\{\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|\cdot I(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2}\geq 12R(K/n_{1})^{\gamma})\}$
$\leq$ $12R(K/n_{1})^{\gamma}+(2R)^{2}\cdot\mathbb{P}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2}\geq 12R(K/n_{1})^{\gamma})$
$\leq$ $12R(K/n_{1})^{\gamma}+4R^{2}(K/n_{1})^{2/d-2\gamma}$
$=$ $\begin{cases}\mathcal{O}\left((K/n_{1})^{\gamma}\right),\quad&\gamma\in(0,2/(3d)],\\ \mathcal{O}\left((K/n_{1})^{2/d-2\gamma}\right),\quad&\gamma\in(2/(3d),1/d).\end{cases}$
The upper bound achieves its minimum when $\gamma=2/(3d)$. In this case, we have $\mathbb{E}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2})\leq\mathcal{O}\left((K/n_{1})^{2/(3d)}\right)$. Thus we obtain the upper bound for the Wasserstein distance between $\mathcal{P}_{1}$ and $\tilde{\mathcal{P}}_{1}$ as
$\mathcal{W}_{2}(\mathcal{P}_{1},\tilde{\mathcal{P}}_{1})=$ $\inf_{\nu\in\Pi(\mathcal{P}_{1},\tilde{\mathcal{P}}_{1})}\left\{\mathbb{E}_{(\bm{X},\tilde{\bm{X}})\sim\nu}(\|\bm{X}-\tilde{\bm{X}}\|_{2}^{2})\right\}^{1/2}.$
$\leq$ $\mathcal{O}\left(\left(\frac{K}{n_{1}}\right)^{\frac{1}{3d}}\right).$
Hence, the upper bound in (A.9) holds.
To validate the lower bound for (A.9), we need to show the following sufficient condition: there exists a constant $C>0$ such that
$\mathbb{E}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|^{2}\mid\bm{X}_{i})\geq C\left(\frac{K}{n_{1}}\right)^{2/d}.$ (A.10)
Denote $V_{d}(r)$ as the volume of a $d$-dimensional ball with radius $r$, then
$V_{d}(r)=\frac{\pi^{d/2}}{\Gamma(d/2+1)}r^{d}\Rightarrow\log V_{d}(r)\propto d\log r.$
Recall that $C_{1}\leq f_{1}(\bm{x})\leq C_{2}$ for $\bm{x}\in{\rm supp}(\mathcal{P}_{1})$. Denote $N_{\bm{x}}(r)$ as the number of samples in $B(\bm{x},r)$ for $n_{1}$ i.i.d. samples from $\mathcal{P}_{1}$, then for any $\bm{x}$ satisfying $B(\bm{x},r)\subseteq{\rm supp}(\mathcal{P}_{1})$, we have
$n_{1}C_{1}V_{d}(r)\leq\mathbb{E}\{N_{\bm{x}}(r)\}\leq n_{1}C_{2}V_{d}(r).$ (A.11)
Consequently, for any $i=1,\ldots,n_{1}$ we have $N_{\bm{X}_{i}}(r)=\mathcal{O}_{p}(n_{1}r^{d})$. On the other hand, since $\bm{X}_{i(k)}$ is the $k$th nearest neighbor of $X_{i}$, we have $N_{\bm{X}_{i}}(\|\bm{X}_{i}-\bm{X}_{i(k)}\|)=k$ for any $k=1,\ldots,K$. Replace $r$ by $\|\bm{X}_{i}-\bm{X}_{i(k)}\|$ in (A.11), we have
$n_{1}C_{1}V_{d}(\|\bm{X}_{i}-\bm{X}_{i(k)}\|)\leq k\leq n_{1}C_{2}V_{d}(\|\bm{X}_{i}-\bm{X}_{i(k)}\|),$
which yields
$\frac{\Gamma(d/2+1)}{C_{2}\pi^{d/2}}(k/n_{1})^{2/d}\leq\|\bm{X}_{i}-\bm{X}_{i(k)}\|^{2}\leq\frac{\Gamma(d/2+1)}{C_{1}\pi^{d/2}}(k/n_{1})^{2/d}.$
Taking conditional expectation given $X_{i}$, we have
$\mathbb{E}(\|\bm{X}_{i(k)}-\bm{X}_{i}\|^{2}\mid\bm{X}_{i})=\mathcal{O}((k/n_{1})^{2/d}).$
Recall that $\tilde{\bm{X}}_{j}=\bm{X}_{i}+U_{i}(\bm{X}_{i(k)}-\bm{X}_{i})$, where $U\sim{\rm Unif}(0,1)$ and $k\sim{\rm Unif}(1,\ldots,K)$. We have
$\mathbb{E}[\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|_{2}^{2}\mid\bm{X}_{i}]$ $=\frac{1}{3}\cdot\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}[\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}^{2}\mid\bm{X}_{i}].$
Thus (A.10) holds and
$\mathbb{E}[\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|_{2}^{2}]=\mathbb{E}\left[\mathbb{E}[\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|_{2}^{2}\mid\bm{X}_{i}]\right]\geq C\left(\frac{K}{n_{1}}\right)^{2/d}.$
This implies that the distribution $\mathcal{P}_{1}$ and $\tilde{\mathcal{P}}_{1}$ are separated by a certain amount, providing a lower bound on the cost of any optimal transport plan between them. Therefore, we have
$\mathcal{W}_{2}(\tilde{\mathcal{P}}_{1},\mathcal{P}_{1})\geq\mathcal{O}\left(\left(\frac{K}{n_{1}}\right)^{1/d}\right).$
Thus (A.9) holds.
Denote $\mu_{1}=\mathbb{E}_{\bm{X}\sim\mathcal{P}_{1}}\{\ell_{f}(\bm{X},1)\}$ and $\tilde{\mu}_{1}=\mathbb{E}_{\bm{X}\sim\tilde{\mathcal{P}}_{1}}\{\ell_{f}(\bm{X},1)\}$. Since $\bm{X}_{1}^{*},\ldots,\bm{X}_{n_{1}^{*}}^{*}$ are iid generated from $\mathcal{P}_{1}$, with probability at least $1-\alpha$,
$\left|\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\mu_{1}\right|\leq\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}.$
Similarly, we also have
$\left|\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i},\tilde{Y}_{i})-\tilde{\mu}_{1}\right|\leq\left(\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}\right).$
By (A.9), we have
$\mathcal{O}_{p}\left((K/n_{1})^{1/d}\right)\leq\mathcal{W}_{2}(\mathcal{P}_{1},\tilde{\mathcal{P}}_{1})\leq\mathcal{O}_{p}\left((K/n_{1})^{1/(3d)}\right).$
By Assumption 2 (A3), we have
$|\mu_{1}-\tilde{\mu}_{1}|=$ $|\mathbb{E}\{\ell_{f}(\bm{X},1)\}-\mathbb{E}\{\ell_{f}(\tilde{\bm{X}},1)\}|$
$=$ $|\mathbb{E}\{\ell_{f}(\bm{X}_{i},1)-\ell_{f}(\bm{X}_{i}+U_{i}(\bm{X}_{i(k)}-\bm{X}_{i}),1)\}|$
$\geq$ $\mathbb{E}\{C_{3}U_{i}\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}\}$
$\geq$ $C_{3}\cdot\mathbb{E}[U_{i}]\cdot\mathbb{E}\{\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}\}$
$\geq$ $c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}$
for some constant $c_{1}>0$. Then with probability at least $1-\alpha$,
$|\hat{\Delta}_{1}|\geq$ $\left|\mu_{1}-\tilde{\mu}_{1}\right|-\left|\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\mu_{1}\right|-\left|\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i},\tilde{Y}_{i})-\tilde{\mu}_{1}\right|$
$\geq$ $c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}-\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}-\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}.$
This completes the proof of Proposition 2. ∎
A.3 Proof of Theorem 3.1
Proof.
Note that
$L^{*}(\hat{f}^{\rm bc})-L^{*}(f^{*})=$ $\left(L^{*}(\hat{f}^{\rm bc})-L^{\rm bc}(\hat{f}^{\rm bc})\right)+\left(L^{\rm bc}(\hat{f}^{\rm bc})-L^{\rm bc}(f^{*})\right)+\left(L^{\rm bc}(f^{*})-L^{*}(f^{*})\right)$
and
$L^{\rm bc}(\hat{f}^{\rm bc})-L^{\rm bc}(f^{*})\leq 0.$
We just need to show that for any prediction function $f\in\mathcal{F}$ and any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$\begin{split}&|L^{\rm bc}(f)-L^{*}(f)|\\ \leq&\frac{\pi_{1}}{2\pi_{0}}\varepsilon_{\rm BT}+\frac{1}{2}\sqrt{\frac{\log(10/\alpha)}{2}}\left\{\frac{\pi_{1}/\pi_{0}}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\left(\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}\right)\right\}.\end{split}$ (A.12)
Recall that
$L^{\rm bc}(f)=$ $\frac{1}{n+\tilde{n}_{1}}\left[\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\tilde{n}_{1}\left\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)+\hat{\Delta}_{0}\right\}\right],$
$L^{*}(f)=$ $\frac{1}{2}\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]+\frac{1}{2}\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)].$
First, rewrite $L^{\rm bc}(f)-L^{*}(f)$ as
$L^{\rm bc}(f)-L^{*}(f)=$ $\underbrace{\frac{n_{1}}{n+\tilde{n}_{1}}\cdot\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\ell_{f}(\bm{X}_{i},1)-\left(\frac{1}{2}-\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\right)\cdot\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]}_{(\rm I)}$
$+\underbrace{\frac{n_{0}}{n+\tilde{n}_{1}}\cdot\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\ell_{f}(\bm{X}_{i},0)-\frac{1}{2}\cdot\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)]}_{(\rm II)}$
$+\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\left\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)+\Delta_{0}-\underbrace{(\hat{\Delta}_{0}-\Delta_{0})}_{(\rm III)}-\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]\right\}.$
Noting that $\mathcal{P}_{1}$ is the pushforward of $\mathcal{P}_{0}$ under $T$, i.e., $\mathcal{P}_{1}=(\mathcal{P}_{0})_{\#T}$, and $h(\bm{x})=\ell_{f}(T(\bm{x}),1)-\ell_{f}(\bm{x},0)$, we have
$\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)+\Delta_{0}-\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]$
$=$ $\underbrace{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]}_{(\rm IV)}$
$+\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]+\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)]-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}[\ell_{f}(\tilde{\bm{X}},0)]-\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(T(\bm{X}),1)]$
$=$ $({\rm IV})+\underbrace{\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]-\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}[\ell_{f}(\tilde{\bm{X}},1)]}_{(\rm V)}$
$+\big(\mathbb{E}_{\tilde{\mathcal{P}}_{0}}[\ell_{f}(T(\tilde{\bm{X}}),1)]-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}[\ell_{f}(\tilde{\bm{X}},0)]\big)-\big(\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(T(\bm{X}),1)]-\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)]\big)$
$=$ $({\rm IV})+({\rm V})+\underbrace{\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})-\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})}_{(\rm VI)}.$
By the Kantorovich-Rubinstein duality and Assumption 1, we have
$|({\rm V})|=$ $\bigg|\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(T(\tilde{\bm{X}}),0)]-\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}[\ell_{f}(T(\tilde{\bm{X}}),0)]\bigg|\leq L_{T}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{1},(\tilde{\mathcal{P}}_{0})_{\#T})\leq L_{T}\varepsilon_{T}.$
By Assumption 1 (A3),
$|({\rm VI})|=\bigg|\mathbb{E}_{\tilde{\mathcal{P}}_{1}}h(\tilde{\bm{X}})-\mathbb{E}_{\mathcal{P}_{1}}h(\bm{X})\bigg|\leq 2\varepsilon_{h}.$
Under the assumption that $n_{0}/(n+\tilde{n}_{1})\to 1/2$, we have
$\frac{n_{1}}{n+\tilde{n}_{1}}\to\frac{\pi_{1}}{2\pi_{0}},\quad\frac{1}{2}-\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\to\frac{\pi_{1}}{2\pi_{0}},\quad\text{ and}\quad\frac{n_{0}}{n+\tilde{n}_{1}}\to\frac{1}{2}.$
Then by Hoeffding’s inequality, for any $t>0$,
$\mathbb{P}\left(|({\rm I})|>\frac{\pi_{1}}{2\pi_{0}}t\right)\leq$ $2\exp\{-2n_{1}t^{2}\},$
$\mathbb{P}\left(|({\rm II})|>\frac{1}{2}t\right)\leq$ $2\exp\{-2n_{0}t^{2}\}.$
$\mathbb{P}\left(|(\rm IV)|>t\right)\leq$ $2\exp\{-2\tilde{n}_{1}t^{2}\}.$
Noting that
$\hat{\Delta}_{0}-\Delta_{0}=\left(\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},0)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)\right)+\left(\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},0)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\tilde{\bm{X}},0)\right),$
the concentration probabilities of the above two terms are given in (A.6) and (A.7). Let
$\begin{split}t_{1}=&\sqrt{\frac{\log(10/\alpha)}{2n_{1}}},\quad t_{2}=\sqrt{\frac{\log(10/\alpha)}{2n_{0}}},\\ t_{31}=&\sqrt{\frac{\log(10/\alpha)}{2n_{0c}}},\quad t_{32}=\sqrt{\frac{\log(10/\alpha)}{2\tilde{n}_{0}}},\quad t_{4}=\sqrt{\frac{\log(10/\alpha)}{2\tilde{n}_{1}}}.\end{split}$ (A.13)
A union bound over the five two-sided events yields that with probability at least $1-\alpha$,
$|(\rm I)|\leq\frac{\pi_{1}}{2\pi_{0}}t_{1},\quad|(\rm II)|\leq\frac{1}{2}t_{2},\quad|(\rm III)|\leq t_{31}+t_{32},\quad|(\rm IV)|\leq t_{4}.$
Since
$L^{\rm bc}(f)-L^{*}(f)=$ $({\rm I})+({\rm II})+\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\left(({\rm III})+({\rm IV})+({\rm V})+({\rm VI})\right),$
we conclude that for any $f\in\mathcal{F}$, with probability at least $1-\alpha$,
$|L^{\rm bc}(f)-L^{*}(f)|\leq\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{\rm BT}+\frac{\pi_{1}}{2\pi_{0}}t_{1}+\frac{1}{2}t_{2}+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}(t_{31}+t_{32}+t_{4}).$
Finally, we have
$L^{*}(\hat{f}^{\rm bc})-L^{*}(f^{*})=$ $\left(L^{*}(\hat{f}^{\rm bc})-L^{\rm bc}(\hat{f}^{\rm bc})\right)+\left(L^{\rm bc}(\hat{f}^{\rm bc})-L^{\rm bc}(f^{*})\right)+\left(L^{\rm bc}(f^{*})-L^{*}(f^{*})\right)$
$\leq$ $\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\varepsilon_{\rm BT}+\frac{\pi_{1}}{\pi_{0}}t_{1}+t_{2}+\frac{\pi_{0}-\pi_{1}}{\pi_{0}}(t_{31}+t_{32}+t_{4}).$
with probability at least $1-\alpha$, where the inequality holds since $L^{\rm bc}(\hat{f}^{\rm bc})-L^{\rm bc}(f^{*})\leq 0$. This completes the proof of Theorem 3.1. ∎
A.4 Proof of Theorem 3.2
Proof.
First, we derive the error bounds for $\|\hat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}$. For the prediction function $f(\bm{x};\bm{\beta})=\sigma(\bm{x}^{\top}\bm{\beta})$ with logistic function $\sigma(t)=1/(1+\exp(-t))$, denote the loss function $\ell(\bm{x},y;\bm{\beta})=\ell_{f}(\bm{x},y)$ for convenience. For example, consider the cross entropy loss function $\ell(\bm{x},y;\bm{\beta})=-y\log(\sigma(\bm{x}^{\top}\bm{\beta}))-(1-y)\log(1-\sigma(\bm{x}^{\top}\bm{\beta}))$.
Considering each source $k=1,\dots,K$, denote
$\psi_{k}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{X}_{ki}\left(Y_{ki}-\sigma(\bm{X}_{ki}^{\top}\bm{\beta})\right).$
Then $\hat{\bm{\beta}}_{k}$ satisfies $\psi_{k}(\hat{\bm{\beta}}_{k})=\bm{0}_{d\times 1}$. Denote the population Hessian matrix as
$\bm{H}_{k}=-\mathbb{E}[\nabla_{\bm{\beta}_{k}}^{2}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta}_{k})]=\mathbb{E}[\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta}_{k})\bm{X}_{ki}\bm{X}_{ki}^{\top}].$
By Assumption 3 (A2), $\lambda_{\min}(\bm{H}_{k})\geq\underline{\kappa}_{k}$, where $\lambda_{\min}(\bm{H}_{k})$ denotes the smallest eigenvalue of $\bm{H}_{k}$. Taking the first-order Taylor expansion of $\psi_{k}(\hat{\bm{\beta}}_{k})$, we have
$\bm{0}_{d\times 1}=\psi_{k}(\bm{\beta}_{k})+\bar{\bm{J}_{k}}(\hat{\bm{\beta}}_{k}-\bm{\beta}_{k}),$
where $\bar{\bm{J}_{k}}=\int_{0}^{1}\bm{J}_{k}(\bm{\beta}_{k}+(\hat{\bm{\beta}}_{k}-\bm{\beta}_{k})t){\rm d}t$ for the Jacobian matrix $\bm{J}_{k}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\left\{\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta})\bm{X}_{ki}\bm{X}_{ki}^{\top}\right\}.$ Rearranging the above equation, we have
$\hat{\bm{\beta}}_{k}-\bm{\beta}_{k}=-(\bar{\bm{J}_{k}})^{-1}\psi_{k}(\bm{\beta}_{k}).$
Next, in order to derive the error bound for $\hat{\bm{\beta}}_{k}-\bm{\beta}_{k}$, we consider the order of $\|(\bar{\bm{J}_{k}})^{-1}\|_{\rm op}$ and $\|\psi_{k}(\bm{\beta}_{k})\|_{2}$. Noting that the Hessian stability gives that $\|\bar{\bm{J}_{k}}-\bm{H}_{k}\|_{\rm op}=o_{p}(1)$, we have
$\|(\bar{\bm{J}_{k}})^{-1}\|_{\rm op}=(1+o_{p}(1))\|\bm{H}_{k}^{-1}\|_{\rm op}\leq(1+o_{p}(1))/\lambda_{\min}(\bm{H}_{k}).$
By the construction, we have $\mathbb{E}[\psi(\bm{\beta}_{k})]=\bm{0}_{d\times 1}$ and $\mathbb{E}\|\psi_{k}(\bm{\beta}_{k})\|_{2}^{2}=\frac{1}{n_{k}}{\rm tr}\left(\bm{H}_{k}\right)$. By Jensen’s inequality,
$\|\psi_{k}(\bm{\beta}_{k})\|_{2}=O_{p}\left(\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}\right).$
Therefore, we have
$\|\hat{\bm{\beta}}_{k}-\bm{\beta}_{k}\|_{2}\leq$ $\|(\bm{J}_{k}(\bar{\bm{\beta}}_{k}))^{-1}\|_{\rm op}\|\psi_{k}(\bm{\beta}_{k})\|_{2}\leq O_{p}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\min}(\bm{H}_{k})}\right).$
$\|\hat{\bm{\beta}}_{k}-\bm{\beta}_{k}\|_{2}\geq$ $O_{p}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\max}(\bm{H}_{k})}\right).$
Secondly, we derive the error bounds for $\|\hat{\bm{\beta}}_{k}^{\rm syn}-\hat{\bm{\beta}}_{k}\|_{2}$. Let $w_{k}=\tilde{n}_{k}/(n_{k}+\tilde{n}_{k1})$. The mixed score is
$\psi_{k}^{\rm mix}(\bm{\beta})=(1-w_{k})\psi_{k}^{\rm raw}(\bm{\beta})+w_{k}\psi_{k}^{\rm syn}(\bm{\beta}),$
with $\psi_{k}^{\rm raw}(\bm{\beta})=n_{k}^{-1}\sum_{i=1}^{n_{k}}\bm{X}_{i}(Y_{i}-\sigma(\bm{X}_{i}^{\top}\bm{\beta}))$ and $\psi_{k}^{\rm syn}(\bm{\beta})=\tilde{n}_{k1}^{-1}\sum_{i=1}^{\tilde{n}_{k1}}\tilde{\bm{X}}_{i}(1-\sigma(\tilde{\bm{X}}_{i}^{\top}\bm{\beta}))$. At $\bm{\beta}_{k}$, $\mathbb{E}[\psi_{k}^{\rm raw}(\bm{\beta}_{k})]=0$ and $\mathbb{E}[\psi_{k}^{\rm syn}(\bm{\beta}_{k})]=\bm{\delta}_{k}$. A mean value expansion gives
$\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}=\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})^{-1}\big\{(1-w_{k})\psi_{k}^{\rm raw}(\bm{\beta}_{k})+w_{k}\psi_{k}^{\rm syn}(\bm{\beta}_{k})\big\},$
with $\bm{J}_{k}^{\rm mix}(\bm{\beta})=(1-w_{k})\frac{1}{n}\sum_{i=1}^{n_{k}}\sigma^{\prime}(\bm{X}_{i}^{\top}\bm{\beta})\bm{X}_{i}\bm{X}_{i}^{\top}+w_{k}\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\sigma^{\prime}(\tilde{\bm{X}}_{j}^{\top}\bm{\beta})\tilde{\bm{X}}_{i}\tilde{\bm{X}}_{i}^{\top}$. By the law of large numbers, $\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})\to\bm{H}_{k}^{\rm mix}:=(1-w_{k})\bm{H}_{k}+w_{k}\tilde{\bm{H}}_{k}$ in probability and $\|\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})^{-1}\|_{\rm op}=(1+o_{p}(1))/\lambda_{\min}(\bm{H}_{k}^{\rm mix})$ and $\|\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})\|_{\rm op}\leq(1+o_{p}(1))\lambda_{\max}(\bm{H}_{k}^{\rm mix})$. Moreover, denote the synthetic score bias
$\bm{\delta}_{k}=\mathbb{E}_{\tilde{\mathcal{P}}_{k}}[\tilde{\bm{X}}(\tilde{Y}-\sigma(\tilde{\bm{X}}^{\top}\bm{\beta}_{k}))]-\mathbb{E}_{\mathcal{P}_{k}}[\bm{X}(Y-\sigma(\bm{X}^{\top}\bm{\beta}_{k}))]=\mathbb{E}_{\tilde{\mathcal{P}}_{k}}[\tilde{\bm{X}}(\tilde{Y}-\sigma(\tilde{\bm{X}}^{\top}\bm{\beta}_{k}))],$
where the last equality holds since $\mathbb{E}_{\mathcal{P}_{k}}[\bm{X}(Y-\sigma(\bm{X}^{\top}\bm{\beta}_{k}))]=0$. Since $d((\bm{x}_{1},y_{1}),(\bm{x}_{2},y_{2}))\leq\|\bm{x}_{1}-\bm{x}_{2}\|+c|y_{1}-y_{2}|$, $\nabla_{\bm{\beta}}\ell(\bm{X},Y,\bm{\beta})$ is Lipschitz continuous with constant $L_{k}=1+(B_{x}B_{\beta})/4+B_{x}/c$. Thus we have $\|\bm{\delta}_{k}\|\leq L_{k}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})$. Note that
$\|\psi_{k}^{\rm raw}(\bm{\beta}_{k})\|_{2}=O_{p}\left(\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}\right),\qquad\|\psi_{k}^{\rm syn}(\bm{\beta}_{k})-\bm{\delta}_{k}\|_{2}=O_{p}\left(\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}\right).$
Therefore
$\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\leq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})+O_{p}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{\rm mix})}\right),$
$\|\hat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\geq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})-O_{p}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{\rm mix})}\right).$
Thirdly, we derive the error bounds for $\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}$. For any $\bm{\beta}\in\mathbb{R}^{d}$, let $\Delta L_{k}^{\rm bc}(\bm{\beta})=L_{k}^{\rm bc}(\bm{\beta})-L_{k}^{*}(\bm{\beta})$. By the optimality, we have
$\nabla_{\bm{\beta}}L_{k}^{\rm bc}(\hat{\bm{\beta}}_{k}^{\rm bc})=\bm{0},\quad\nabla_{\bm{\beta}}L_{k}^{*}(\bm{\beta}_{k}^{*})=\bm{0}.$
Taking a mean value expansion, we have
$\bm{H}_{k}(\bar{\bm{\beta}}_{k})(\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*})=-\nabla_{\bm{\beta}}\left(\Delta L_{k}^{\rm bc}(\hat{\bm{\beta}}_{k}^{\rm bc})\right)\quad\text{ with }\quad\lambda_{\min}(\bm{H}_{k}(\bar{\bm{\beta}}_{k}))\geq\underline{\kappa}_{k},$
for $\bar{\bm{\beta}}_{k}$ on the segment between $\hat{\bm{\beta}}_{k}^{\rm bc}$ and $\bm{\beta}_{k}^{*}$. Thus we have
$\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*}=-\left(\bm{H}_{k}(\bar{\bm{\beta}}_{k})\right)^{-1}\cdot\nabla_{\bm{\beta}}\left(\Delta L_{k}^{\rm bc}(\hat{\bm{\beta}}_{k}^{\rm bc})\right)$
and
$\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}\left\|\nabla_{\bm{\beta}}\left(\Delta L_{k}^{\rm bc}(\hat{\bm{\beta}}_{k}^{\rm bc})\right)\right\|_{2}.$
Taking the derivative of $L_{k}^{\rm bc}$ and $L_{k}^{*}$ with respect to $\bm{\beta}$, we have
$\nabla_{\bm{\beta}}L_{k}^{\rm bc}(\bm{\beta})=\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\tilde{n}_{k1}\cdot\left\{\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})+\nabla_{\bm{\beta}}\hat{\Delta}_{k0}\right\}\right],$
where
$\nabla_{\bm{\beta}}\hat{\Delta}_{k0}=\frac{1}{n_{k0,c}}\sum_{i\in\mathcal{S}_{k0,c}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},0;\bm{\beta})+\frac{1}{\tilde{n}_{k0}}\sum_{i=1}^{\tilde{n}_{k0}}\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}}_{ki}^{(0)},0;\bm{\beta}).$
Denote
$\nabla_{\bm{\beta}}L_{k}^{*}(\bm{\beta})=\frac{1}{2}\nabla_{\bm{\beta}}\mathbb{E}_{\tilde{\mathcal{P}}_{k1}}\ell(\bm{X},1;\bm{\beta})+\frac{1}{2}\nabla_{\bm{\beta}}\mathbb{E}_{\tilde{\mathcal{P}}_{k0}}\ell(\bm{X},0;\bm{\beta}).$
Using the same calculation in Section A.1, we have
$\nabla_{\bm{\beta}}\left\{L_{k}^{\rm bc}(\bm{\beta})-L_{k}^{*}(\bm{\beta})\right\}=({\rm i})+({\rm ii})+\frac{\tilde{n}_{k1}}{n_{k}+\tilde{n}_{k1}}(({\rm iii})+({\rm iv})+({\rm v})+({\rm vi})),$
where
$({\rm i})=$ $\frac{n_{k1}}{n_{k}+\tilde{n}_{k1}}\cdot\frac{1}{n_{k1}}\sum_{i=1}^{n_{k1}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},1;\bm{\beta})-\left(\frac{1}{2}-\frac{\tilde{n}_{k1}}{n_{k}+\tilde{n}_{k1}}\right)\cdot\mathbb{E}_{\mathcal{P}_{k1}}[\nabla_{\bm{\beta}}\ell(\bm{X},1;\bm{\beta})],$
$({\rm ii})=$ $\frac{n_{k0}}{n_{k}+\tilde{n}_{k0}}\cdot\frac{1}{n_{k1}}\sum_{i=n_{k1}+1}^{n_{k}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},0;\bm{\beta})-\frac{1}{2}\cdot\mathbb{E}_{\mathcal{P}_{k0}}[\nabla_{\bm{\beta}}\ell(\bm{X},0;\bm{\beta})],$
$({\rm iii})=$ $\nabla_{\bm{\beta}}\left(\hat{\Delta}_{k0}-\Delta_{k0}\right),$
$({\rm iv})=$ $\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})-\mathbb{E}_{\tilde{\mathcal{P}}_{k1}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})],$
$({\rm v})=$ $\mathbb{E}_{\tilde{\mathcal{P}}_{k1}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})]-\mathbb{E}_{(\tilde{\mathcal{P}}_{k0})_{\#T_{k}}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})],$
$({\rm vi})=$ $\mathbb{E}_{\tilde{\mathcal{P}}_{k0}}[\nabla_{\bm{\beta}}h_{k}(\bm{X})]-\mathbb{E}_{\mathcal{P}_{k0}}[\nabla_{\bm{\beta}}h_{k}(\bm{X})].$
By Hoeffding’s inequality and a union bound across $d$ coordinates, there exists a constant $C_{0k}>0$ such that, with probability at least $1-\alpha$,
$\|(\rm i)\|_{2}\leq$ $\frac{\pi_{1}}{2\pi_{0}}C_{0k}R\sqrt{\frac{\log(10d/\alpha)}{n_{k1}}},$
$\|(\rm ii)\|_{2}\leq$ $\frac{1}{2}C_{0k}R\sqrt{\frac{\log(10d/\alpha)}{n_{k0}}},$
$\|(\rm iii)\|_{2}\leq$ $C_{0k}R\sqrt{\frac{\log(10d/\alpha)}{n_{k0,c}}}+C_{0k}R\sqrt{\frac{\log(10d/\alpha)}{\tilde{n}_{k0}}},$
$\|(\rm iv)\|_{2}\leq$ $C_{0k}R\sqrt{\frac{\log(10d/\alpha)}{\tilde{n}_{k1}}}.$
By the Kantorovich-Rubinstein duality, we have
$\|({\rm v})\|_{2}\leq L_{g}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k1},(\tilde{\mathcal{P}}_{k0})_{\#T_{k}})\leq L_{g}\varepsilon_{T}.$
Finally, by Assumption 1 (A3),
$\|(\rm vi)\|_{2}\leq 2\varepsilon_{h}.$
Therefore, with probability at least $1-\alpha$, we have
$\left\|\nabla_{\bm{\beta}}(L_{k}^{\rm bc}(\hat{\bm{\beta}}_{k}^{\rm bc})-L_{k}^{*}(\hat{\bm{\beta}}_{k}^{\rm bc}))\right\|_{2}\leq\underbrace{\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\left(2\varepsilon_{h}+L_{g}\cdot\varepsilon_{T}\right)}_{\varepsilon_{{\rm BT},k}}$
$\quad+\underbrace{C_{0k}R\sqrt{\log(10d/\alpha)}\left\{\frac{\pi_{1}/(2\pi_{0})}{\sqrt{n_{k1}}}+\frac{1}{2\sqrt{n_{k0}}}+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\left(\frac{1}{\sqrt{n_{k0,c}}}+\frac{1}{\sqrt{\tilde{n}_{k0}}}+\frac{1}{\sqrt{\tilde{n}_{k1}}}\right)\right\}}_{\varepsilon_{{\rm sampling},k}}.$
Consequently, with probability at least $1-\alpha$,
$\|\hat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}(\varepsilon_{{\rm BT},k}+\varepsilon_{{\rm sampling},k}).$
This completes the proof of Theorem 3.2.
A.5 Proof of Theorem 3.3
Proof.
Denote
$\varphi(W)=\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})-\tau.$
Thus it suffices to show that $\mathbb{E}[\varphi(W)]=0.$ Decompose the error of the AIPW estimator as
$\hat{\tau}^{\rm AIPW}-\tau=$ $\mathbb{P}_{n}[\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})]-\mathbb{P}[\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})]$
$=$ $(\mathbb{P}_{n}-\mathbb{P})[\varphi(W)]+\mathbb{P}\big(\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)$
$+(\mathbb{P}_{n}-\mathbb{P})[\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})].$
We first focus on the influence function fluctuation term $(\mathbb{P}_{n}-\mathbb{P})[\varphi(W)]$. By Assumption 4,
$|\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})-\tau|\leq|\mu_{1}^{*}(\bm{X})-\mu_{0}^{*}(\bm{X})|+\frac{|Y-\mu_{1}^{*}(\bm{X})|}{\eta}+\frac{|Y-\mu_{0}^{*}(\bm{X})|}{\eta}\leq C(\eta,M).$
By Hoeffding’s inequality, with probability at least $1-\alpha/2$,
$|(\mathbb{P}_{n}-\mathbb{P})[\varphi(W)]|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}.$ (A.14)
Next, we derive the upper bound for the population bias term $\mathbb{P}\big(\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)$. For simplicity, let $\delta_{e}=\hat{e}-e^{*}$ and $\delta_{a}=\hat{\mu}_{a}-\mu_{a}^{*}$ for $a\in\{0,1\}$. Note that
$\left\{\hat{\mu}_{1}(\bm{X})-\frac{Z(Y-\hat{\mu}_{1}(\bm{X}))}{\hat{e}(\bm{X})}\right\}-\left\{\mu_{1}^{*}(\bm{X})-\frac{Z(Y-\mu_{1}^{*}(\bm{X}))}{e^{*}(\bm{X})}\right\}$
$=$ $\delta_{1}(\bm{X})\left(1-\frac{Z}{\hat{e}(\bm{X})}\right)+Z\left(\frac{e^{*}(\bm{X})}{\hat{e}(\bm{X})}-1\right)\cdot\frac{Y-\mu_{1}^{*}(\bm{X})}{e^{*}(\bm{X})}.$
Taking the expectation conditional on $\bm{X}$ on the right-hand side, by Assumption 4 (C1), we have
$\mathbb{E}\left[\left\{\hat{\mu}_{1}(\bm{X})-\frac{Z(Y-\hat{\mu}_{1}(\bm{X}))}{\hat{e}(\bm{X})}\right\}-\left\{\mu_{1}^{*}(\bm{X})-\frac{Z(Y-\mu_{1}^{*}(\bm{X}))}{e^{*}(\bm{X})}\right\}\bigg|\bm{X}\right]=\delta_{1}(\bm{X})\left(1-\frac{e^{*}(\bm{X})}{\hat{e}(\bm{X})}\right).$
We can use a similar way to derive that
$\mathbb{E}\left[\left\{\hat{\mu}_{0}(\bm{X})-\frac{(1-Z)(Y-\hat{\mu}_{0}(\bm{X}))}{1-\hat{e}(\bm{X})}\right\}-\left\{\mu_{0}^{*}(\bm{X})-\frac{(1-Z)(Y-\mu_{0}^{*}(\bm{X}))}{1-e^{*}(\bm{X})}\right\}\bigg|\bm{X}\right]$
$=$ $\delta_{0}(\bm{X})\left(1-\frac{1-e^{*}(\bm{X})}{1-\hat{e}(\bm{X})}\right).$
Expanding the population bias term, we have
$\mathbb{P}\big(\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)$
$=$ $\mathbb{E}\left\{\delta_{1}(\bm{X})\left(1-\frac{e^{*}(\bm{X})}{\hat{e}(\bm{X})}\right)\right\}-\mathbb{E}\left\{\delta_{0}(\bm{X})\left(1-\frac{1-e^{*}(\bm{X})}{1-\hat{e}(\bm{X})}\right)\right\}.$
Using the overlap condition that $\hat{e},e^{*}\in[\eta,1-\eta]$ and the Cauchy-Schwarz inequality, we have
$|\mathbb{P}\big(\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)|\leq$ $\frac{1}{\eta}\left(\mathbb{E}[|\delta_{1}(\bm{X})\delta_{e}(\bm{X})|]+\mathbb{E}[|\delta_{0}(\bm{X})\delta_{e}(\bm{X})|]\right)$
$\leq$ $\frac{C_{1}}{\eta}(r_{1}+r_{0})r_{e}.$ (A.15)
Finally, we consider the second-order empirical remainder $(\mathbb{P}_{n}-\mathbb{P})[\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})]$. Denote $\Delta\psi(\bm{W})=\psi(W;\hat{\mu}_{1},\hat{\mu}_{0},\hat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})$. By the Cauchy-Schwarz inequality, we have
$\|\Delta\psi\|_{L_{2}(\mathcal{P}_{X})}\leq C(\eta,M)[(r_{1}+r_{0})r_{e}+r_{1}r_{0}].$
Hoeffding’s inequality gives that with probability at least $1-\alpha/2$,
$|(\mathbb{P}_{n}-\mathbb{P})\Delta\psi(W)|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}[(r_{1}+r_{0})r_{e}+r_{1}r_{0}].$ (A.16)
Consequently, combining (A.14), (A.15) and (A.16), the proof of Theorem 3.3 is completed.
Appendix B Synthetic Generators
We briefly review some synthetic generating methods in this section.
Reweighting and Bootstrap.
Reweighting is an intuitive oversampling technique used to address imbalanced data in machine learning. This approach works by assigning a higher weight to samples from the minority class so that the training process focuses more on learning from the underrepresented group. For instance, consider a dataset with $n_{1}$ minority samples and $n_{0}$ majority samples with $n_{1}\ll n_{0}$. A common reweighting approach assigns a weight of $w_{1}=\lfloor n_{0}/n_{1}\rfloor$ to each minority sample and a weight of $w_{0}=1$ to each majority sample [Breiman et al., 2017]. This approach is equivalent to oversampling the minority class by replicating each minority sample $\lfloor n_{0}/n_{1}\rfloor-1$ times and training on the resulting augmented dataset with equal weight.
In contrast, bootstrap methods [Efron and Tibshirani, 1994] for imbalanced classification generate synthetic samples by randomly drawing with replacement from the minority samples. Bootstrap can be regarded as a generalization of the fixed-weight reweighting approach as it effectively assigns random weights to the minority samples in each resampling step. While both reweighting and bootstrap are intuitive and straightforward to implement, they are sensitive to outliers in the minority class. By heavily emphasizing or replicating the outliers, these approaches can potentially lead to overfitting to the noise present in the minority group.
Gaussian Mixture Model (GMM).
Gaussian mixture model [McLachlan and Peel, 2000] is an oversampling technique that assumes the minority samples follow a mixture of multivariate Gaussian distributions with unknown means and covariance matrices. This technique typically fits a single Gaussian component to the minority class. Given the minority samples $\bm{X}_{1},\dots,\bm{X}_{n_{1}}$, this approach first estimates the distributional parameters, including the empirical mean $\hat{\bm{\mu}}_{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\bm{X}_{i}$ and the sample covariance matrix $\hat{\bm{\Sigma}}_{1}=\frac{1}{n_{1}-1}\sum_{i=1}^{n_{1}}(\bm{X}_{i}-\hat{\bm{\mu}}_{1})(\bm{X}_{i}-\hat{\bm{\mu}}_{1})^{\top}$. Next, the synthetic samples are generated by randomly drawing from the estimated Gaussian distribution with mean $\hat{\bm{\mu}}_{1}$ and covariance matrix $\hat{\bm{\Sigma}}_{1}$. This approach effectively captures the first two moments of the minority distribution. However, the strong underlying Gaussian distribution assumption imposes significant constraints. It might generate poorly representative synthetic samples when the true minority distribution is not unimodal, particularly when the distribution is heavy tailed. For example, it involves a non-convex support or is highly skewed. In such cases, the synthetic data fails to accurately reflect the manifold of the minority class. Consequently, the introduction of the synthetic noise can potentially degrade the performance and robustness of the following training step.
Synthetic Minority Oversampling TEchnique (SMOTE).
SMOTE, introduced by Chawla et al. [2002], is a widely used oversampling method that generates synthetic minority samples in imbalanced datasets. SMOTE generates new synthetic samples by linearly interpolating between pairs of minority samples. It works as follows: for a randomly selected minority class sample, first find its $K$ nearest neighbors in the minority group. Then randomly select one of these $K$ nearest neighbors and create a new point along the line segment between the original point and the chosen neighbor. This procedure is repeated until the desired number of synthetic samples is reached. SMOTE requires a hyperparameter $K$, the number of nearest neighbors considered for each minority sample. Algorithm 2 provides a step-wise description on how SMOTE generates $\tilde{n}_{1}$ synthetic samples based on input data $\bm{X}_{1},\ldots,\bm{X}_{n_{1}}$.
Algorithm 2 Synthetic Minority Oversampling TEchnique (SMOTE)
0: Samples $(\bm{X}_{i})_{i=1}^{n_{1}}$, the number of nearest neighbors $K$, synthetic sample size $\tilde{n}_{1}$.
1:for each $i$ in $1:n_{1}$ do
2:  Find the $K$ nearest neighbors of $\bm{X}_{i}$, denoted as $\bm{X}_{i(1)},\ldots,\bm{X}_{i(K)}$.
3:end for
4:for each $i$ in $1:\tilde{n}_{1}$ do
5:  Sample index $t$ uniformly from $\{1,2,\dots,n_{1}\}$.
6:  Sample $U_{i}$ from $\mathcal{U}(0,1)$, i.e., from the uniform distribution on the interval [0, 1].
7:  Sample $k$ uniformly from $\{1,\ldots,K\}$.
8:  Generate the SMOTE sample $\tilde{\bm{X}}_{i}^{(1)}\leftarrow\bm{X}_{t}+U_{i}(\bm{X}_{t(k)}-\bm{X}_{t})$.
9:end for
9: Synthetic samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$.
Diffusion Model.
Diffusion models [Ho et al., 2020, Song et al., 2020] form one of the most popular classes of generative models for data synthesis. A diffusion model learns the distribution of observed samples by simulating and statistically revising a Markovian diffusion process that maps the data to standard Gaussian noise and then reconstructs data from noise. The framework consists of two phases: a fixed forward process, which maps a data example to Gaussian noise, and a learned backward process, which iteratively maps random noise back to a realistic data sample.
The forward process is a fixed Markov chain that progressively corrupts a sample with Gaussian noise over $T$ time steps, parameterized by a schedule of variance terms $\beta_{t}\in(0,1)$ for $t=1,\dots,T$. Starting with an original data sample $\bm{x}\in\mathbb{R}^{d}$ and letting $\bm{z}_{0}=\bm{x}$, a series of intermediate latent variables $\bm{z}_{1},\dots,\bm{z}_{T}\in\mathbb{R}^{d}$ are generated according to the following iterative equation,
$\bm{z}_{t}=$ $\>\sqrt{1-\beta_{t}}\cdot\bm{z}_{t-1}+\sqrt{\beta_{t}}\cdot\bm{\epsilon}_{t},\quad t=1,\dots,T,$
where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ is noise added at time $t$. Denoting $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{t})$ for $t=1,\dots,T$, this process allows for a direct-sampling property, which makes it possible to obtain $\bm{z}_{t}$ from $\bm{x}$ in one step:
$\bm{z}_{t}=\sqrt{\alpha_{t}}\cdot\bm{x}+\sqrt{1-\alpha_{t}}\cdot\bm{\epsilon},$
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$. Since $\beta_{t}<1$ is chosen such that $\alpha_{T}\approx 0$ for large $T$, the final latent variable $\bm{z}_{T}$ is guaranteed to be close to the standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}_{d})$.
The backward process defines the generative model. It defines a learned Markov chain that attempts to reverse the diffusion process, starting from pure noise $\bm{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ and iteratively denoising it back to a data sample $\bm{z}_{0}$:
$\bm{z}_{t-1}\mid(\bm{z}_{t},\phi_{t})\sim\mathcal{N}(\bm{f}_{t}(\bm{z}_{t},\phi_{t}),\sigma_{t}^{2}\bm{I}),\quad t=T,T-1,\dots,1.$
The function $\bm{f}_{t}(\bm{z}_{t},\phi_{t})$ is a neural network that is trained to estimate the mean of the approximate Gaussian distribution for the mapping from $\bm{z}_{t}$ to $\bm{z}_{t-1}$, and $\sigma_{t}$ is predetermined by the variance parameter $\beta_{t}$. By chaining these steps, diffusion models can synthesize high-fidelity data by gradually transforming Gaussian noise to structured samples.
Flow matching.
Flow matching [Lipman et al., 2022] aims to learn a smooth and invertible map from a simple base distribution, say, the standard Gaussian distribution, to the target data distribution. For observation $\bm{x}\in\mathbb{R}^{d}$, consider the probability density path $p:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\int p_{t}(\bm{x}){\rm d}\bm{x}=1$ for any $t\in[0,1]$. Let $p_{0}$ be the simple base distribution, and $p_{1}$ be the target data distribution. Define the flow $\bm{\phi}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ as a time-dependent differomorphic map satisfying that if $\bm{x}\sim p_{0}$, then $\bm{\phi}_{t}(\bm{x})\sim p_{t}$. Without loss of generality, let $\bm{\phi}_{0}(\bm{x})=\bm{x}$. The flow can be generated by a continuous normalizing flow vector field $\bm{v}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that satisfies $\frac{\rm d}{{\rm d}t}\bm{\phi}_{t}(\bm{x})=\bm{v}_{t}(\bm{\phi}_{t}(\bm{x}))$ [Chen et al., 2018]. Flow matching simplifies the learning problem by utilizing conditional flows that define a straight path between the noise and a data point. For a given data point $\bm{x}_{1}\sim p_{1}$ and a noise sample $\bm{x}_{0}\sim p_{0}$, the optimal conditional vector field is the straight line path $\bm{u}_{t}(\bm{x})=\bm{x}_{1}-\bm{x}_{0}$.
The goal of flow matching is to train a neural network field $\bm{v}_{t}(\bm{x};\theta)$ parameterized by $\theta$ to match the ideal conditional vector field $\bm{u}_{t}$ in expectation. Suppose the target probability density path $p_{t}$ is generated by the vector field $\bm{u}_{t}$, flow matching aims to minimize the objective function
$\mathcal{L}_{\rm FM}(\theta)=\mathbb{E}_{t,p_{0},p_{1}}\left[\|v_{t}(\bm{x}_{0}+t(\bm{x}_{1}-\bm{x}_{0});\theta)-(\bm{x}_{1}-\bm{x}_{0})\|^{2}\right],$
where $t\sim U(0,1)$, $\bm{x}_{0}\sim p_{0}$ and $\bm{x}_{1}\sim p_{1}$. With the learned vector field $\bm{v}_{t}(\bm{x};\theta)$ and a random noise sample $\bm{z}\sim p_{0}$, synthetic samples are generated by
$\tilde{\bm{x}}=\bm{\phi}_{1}(\bm{z}),\quad\text{where }\frac{\rm d}{{\rm d}t}\bm{\phi}_{t}(\bm{z})=\bm{v}_{t}(\bm{\phi}_{t}(\bm{z});\theta).$
There are many other synthetic generators, such as generative adversarial networks (GANs) [Goodfellow et al., 2014, 2020], normalizing flows [Rezende and Mohamed, 2015], and variational autoencoders (VAE) [Kingma and Welling, 2013]. Please see Figueira and Vaz [2022], Lu et al. [2023] for a comprehensive survey.