### v2

# Bias-Corrected Data Synthesis for Imbalanced Learning

Abstract
Imbalanced data, where the positive samples represent only a small proportion compared to the negative samples, makes it challenging for classification problems to balance the false positive and false negative rates. A common approach to addressing the challenge involves generating synthetic data for the minority group and then training classification models with both observed and synthetic data. However, since the synthetic data depends on the observed data and fails to replicate the original data distribution accurately, prediction accuracy is reduced when the synthetic data is naïvely treated as the true data. In this paper, we address the bias introduced by synthetic data and provide consistent estimators for this bias by borrowing information from the majority group. We propose a bias correction procedure to mitigate the adverse effects of synthetic data, enhancing prediction accuracy while avoiding overfitting. This procedure is extended to broader scenarios with imbalanced data, such as imbalanced multi-task learning and causal inference. Theoretical properties, including bounds on bias estimation errors and improvements in prediction accuracy, are provided. Simulation results and data analysis on handwritten digit datasets demonstrate the effectiveness of our method.
Keywords: Bias correction; Empirical risk minimization; Imbalanced classification; Oversampling; Synthetic data.
1 Introduction
1.1 Background
Imbalanced classification is a fundamental challenge in modern machine learning, arising when the number of observations in one class significantly exceeds that in another class. This issue is prevalent in diverse applications, including detecting rare diseases in medical diagnosis (Rajkomar et al., 2019; Faviez et al., 2020), fraud detection (Subudhi and Panigrahi, 2018), anomaly detection in industrial systems (Kong et al., 2020), and cybersecurity (Sarker, 2019). Traditional classification algorithms often perform poorly under such an imbalance, as they tend to be biased towards the majority class, leading to suboptimal sensitivity and an increasing risk of overlooking critical minority instances.
A common strategy for addressing such challenges in imbalanced classification is data augmentation, which aims to rebalance the samples in different classes by artificially modifying or expanding the training dataset. Resampling-based approaches include undersampling (removing samples from the majority class) and oversampling (expanding the minority class). Undersampling techniques, such as Tomek’s links (Tomek, 1976) and cluster centroid (Lemaître et al., 2017), often suffer from information loss due to discarding potentially informative majority samples. In contrast, oversampling is typically preferred, and various methods have been proposed to enrich the minority class. The reweighting procedure, which assigns higher weights to the minority samples, is equivalent to oversampling by replicating the minority samples. While bootstrap (Efron and Tibshirani, 1994) is a widely used resampling method in statistics, its naïve application in oversampling may be sensitive to outliers and may introduce variance inflation. The Gaussian mixture model (McLachlan and Peel, 2000) is a classical tool for modeling heterogeneous populations through parametric mixture distributions. Perturbed sampling techniques (He and Garcia, 2009) have been widely used to address class imbalance by modifying the sampling mechanism across classes. Both approaches provide convenient modeling abstractions, though they rely on simplified assumptions and may not fully capture the complexity of real-world data.
Among oversampling methods, the Synthetic Minority Oversampling TEchnique (SMOTE, Chawla et al. (2002)) has been especially influential. SMOTE generates synthetic samples by interpolating between minority samples and has inspired numerous variants, such as Borderline-SMOTE (Han et al., 2005), ADASYN (He et al., 2008), and safe-level-SMOTE (Bunkhumpornpat et al., 2009), which aim to better capture the geometry of the data and concentrate synthetic sample generation near the decision boundary where classification is challenging. For a comprehensive review of resampling techniques in imbalanced settings, see Mohammed et al. (2020).
Beyond empirical success, recent theoretical studies have examined the statistical properties of synthetic procedures and their impact on classification risk. For example, Elreedy et al. (2024) and Sakho et al. (2024) separately derive the probability distribution of SMOTE-generated synthetic samples, with the latter further proving that the synthetic density function vanishes near the boundary of the minority support. Another widely used augmentation method is Mixup (Zhang et al., 2017), which generates new samples by convex combinations of covariates and their labels. In Xia and Klusowski (2026), imbalanced classification is viewed through a transfer-learning lens under label shift, and regimes are identified in which oversampling can be detrimental. From a learning-theoretic perspective, uniform concentration bounds are established in Ahmad et al. (2025) that link the empirical risk computed on synthetic samples to the true minority risk, yielding excess-risk guarantees for kernel-based classifiers. Theoretical results for Mixup include robustness against adversarial attacks and improved generalization by reducing overfitting (Zhang et al., 2020), as well as conditions under which Mixup helps reduce calibration errors (Naeini et al., 2015; Zhang et al., 2022).
Some recent work suggests that resampling/upsampling does not always improve performance under class imbalance and can even be counterproductive in certain prediction settings. Large empirical comparisons find that the benefits of common sampling strategies are highly dataset- and model-dependent, with no uniformly best choice across imbalance regimes (Newaz et al., 2022). In clinical risk prediction with logistic regression, it is shown that oversampling and SMOTE can substantially degrade probability calibration without improving discrimination, and that comparable operating points can often be achieved via threshold adjustment (Van den Goorbergh et al., 2022). Assunção et al. (2024) argue, under a calibration-centric perspective, that adjusting the decision cutoff may recover much of the benefit typically attributed to augmentation. Motivated by the dependence introduced when synthetic samples are generated from the training data, Tian and Shen (2025) propose a sample-splitting framework in which one subset is used to generate synthetic samples and an independent subset is used for training. Nevertheless, a fundamental question remains: under what conditions do synthetic procedures improve classification, and how can potential adverse effects be avoided? In this work, we aim to address this question by developing a principled framework to analyze and correct the distributional discrepancy introduced by synthetic data.
1.2 Our Contributions
We summarize our contributions as follows:
Bias-corrected synthetic data augmentation for imbalanced classification. We develop a bias correction methodology that effectively estimates and adjusts for the discrepancy between the synthetic distribution and the true distribution. By borrowing information from the majority class, our procedure builds a bridge between the observed data and the otherwise unobservable bias in the minority class. Since the minority bias induced by synthetic data is non-negligible, our procedure effectively reduces the bias by an explicit correction term. Theoretically, this bias correction procedure results in improved performance for suboptimal synthetic generators, as confirmed by both simulations and data analysis.
Theoretical guarantees and error bounds. We provide non-asymptotic error bounds for estimators based on raw data, synthetic augmentation, and bias correction methodology. These results identify the regimes where bias correction yields substantial improvement and clarify the trade-offs between variance reduction and bias inflation under different levels of imbalance. Our theoretical results answer the questions of when synthetic augmentation alone suffices and when bias correction is indispensable.
Unified framework with practical validation. We design a general framework that integrates bias correction with diverse synthetic generators, including Gaussian mixture, perturbed sampling, and SMOTE. Through extensive simulations and real-world data analysis, we demonstrate that the proposed method consistently enhances both predictive accuracy and parameter estimation, offering robustness across different imbalance ratios and model architectures.
2 Methodology
We begin by introducing the setting for binary classification with imbalanced data. Suppose that the training data consist of $n$ independent and identically distributed (i.i.d.) samples $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, where $\bm{X}_{i}\in\mathcal{X}\subseteq\mathbb{R}^{d}$ is a $d$-dimensional covariate vector and $Y_{i}\in\{0,1\}$ represents the class label. Assume that $Y_{i}$’s follow a Bernoulli distribution with $\pi_{1}=\mathbb{P}(Y_{i}=1)$ and $\pi_{0}=\mathbb{P}(Y_{i}=0)$. In the imbalanced setting, we assume $0<\pi_{1}<1/2$, so that the class $Y=1$ is underrepresented. For convenience, we refer to $Y=1$ as the minority class and $Y=0$ as the majority class. Let $n_{1}=\sum_{i=1}^{n}Y_{i}$ and $n_{0}=n-n_{1}$ denote the respective sample sizes, with $n_{1}\ll n_{0}$ with high probability. Without loss of generality, we assume that the samples are ordered such that $Y_{1}=\dots=Y_{n_{1}}=1$ and $Y_{n_{1}+1}=\dots=Y_{n}=0$. We also assume that $\bm{X}\mid(Y=1)\sim\mathcal{P}_{1}$ and $\bm{X}\mid(Y=0)\sim\mathcal{P}_{0}$, where $\mathcal{P}_{1},\mathcal{P}_{0}$ represent the class-conditional distributions.
2.1 Bias Correction with Synthetic Data
Our goal is to learn a prediction function $f:\mathcal{X}\to(0,1)$ that provides accurate probability estimates for $Y_{n+1}$ given a new covariate vector $\bm{X}_{n+1}$. Our framework focuses on risk minimization with respect to a general loss function and does not impose a specific classification or decision rule. In particular, the learned function $f$ can be combined with any downstream classification rule depending on the application, such as thresholding or cost-sensitive decisions. One widely used example is to obtain the prediction via $\widehat{Y}_{n+1}=I(f(\bm{X}_{n+1})\geq 1/2)$ where $I(\cdot)$ denotes the indicator function taking values in $\{0,1\}$. Considering the family of prediction functions $\mathcal{F}=\{f:\mathbb{R}^{d}\to(0,1)\}$, denote the corresponding loss function as a binary cross-entropy loss, for example, $\ell_{f}(\bm{X},Y)=-Y\log f(\bm{X})-(1-Y)\log(1-f(\bm{X}))$. Throughout this paper, the proposed correction is formulated at the loss and risk level and is agnostic to the specific classification architecture. Our framework applies to classifiers $f\in\mathcal{F}$ trained via empirical risk minimization, including (but not limited to) logistic regression and neural networks with cross-entropy loss, as well as margin-based methods such as calibrated support vector machines and boosting.
With the raw data $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, we can simply train the prediction function by minimizing the empirical loss function $L^{\rm raw}$:
$L^{\rm raw}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i}).$ (2.1)
In the case that $n_{1}\ll n_{0}$, a trivial guess that all samples are from the majority group will result in accuracy as high as $n_{0}/n$, which is close to $1$, but it does not provide information from the data. To deal with this problem, an intuitive way is to make the data balanced by adding synthetic data samples to the minority group. Assume that we have $\tilde{n}_{1}$ synthetic samples for the minority group: $(\tilde{\bm{X}}^{(1)}_{i},\tilde{Y}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$, where $\tilde{Y}_{i}^{(1)}=1$ for $i=1,\ldots,\tilde{n}_{1}$ with $\tilde{n}_{1}\approx n_{0}-n_{1}$. By equally treating the synthetic and raw samples, we can run the algorithm by minimizing the synthetic-augmented loss function $L^{\rm syn}$:
$L^{\rm syn}(f)=\frac{1}{n+\tilde{n}_{1}}\bigg(\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})\bigg).$ (2.2)
See Figure 1 for the illustration of the imbalanced learning based on raw data and synthetic augmentation. By introducing $\tilde{n}_{1}$ synthetic samples from the minority group, $L^{\rm syn}$ is a loss function from a “balanced” dataset, especially compared with $L^{\rm raw}$. While $L^{\rm syn}$ helps improve the prediction accuracy, a concern arises when the synthetic data fails to exactly recover the distribution of the minority group $\mathcal{P}_{1}$. This will cause a bias between the loss functions of data from the true distribution $\mathcal{P}_{1}$ to the synthetic distribution $\tilde{\mathcal{P}}_{1}$ as follows,
$\Delta_{1}=\mathbb{E}_{\bm{X}\sim\mathcal{P}_{1}}\{\ell_{f}(\bm{X},1)\}-\mathbb{E}_{\tilde{\bm{X}}\sim\tilde{\mathcal{P}}_{1}}\{\ell_{f}(\tilde{\bm{X}},1)\}.$ (2.3)
Figure 1: Imbalanced learning based on raw data, synthetic augmentation and bias correction.
Since the sample size in the minority class is limited, and the synthetic samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$ are generated conditional on the observed samples $(\bm{X}_{i})_{i=1}^{n_{1}}$, a direct comparison between the empirical losses $n_{1}^{-1}\sum_{i=1}^{n_{1}}\ell_{f}(\bm{X}_{i},1)$ and $\tilde{n}_{1}^{-1}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}},1)$ leads to variance inflation and induces a complicated dependence structure in the resulting difference. To overcome the two challenges, consider $n_{1}^{*}$ satisfying $n_{1}^{*}\approx n_{0}-n_{1}$. Imagine that we have $n_{1}^{*}$ i.i.d. unobserved samples from the minority group $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ independent of the observed data, where $Y_{i}^{*}=1$ and $\bm{X}_{i}^{*}\mid(Y_{i}^{*}=1)\sim\mathcal{P}_{1}$. By introducing the unobserved minority samples, the total dataset $(\bm{X}_{i},Y_{i})_{i=1}^{n}$ and $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ is roughly balanced. Denote the sample bias caused by the synthetic data as
$\widehat{\Delta}_{1}=\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)}).$ (2.4)
Suppose we have the balanced data including observed dataset $(\bm{X}_{i},Y_{i})_{i=1}^{n}$ and unobserved dataset $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$. The empirical loss function for the balanced dataset is
$L^{\rm bal}(f)=$ $\frac{1}{n+n_{1}^{*}}\left\{\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})\right\}$
$=$ $\frac{1}{n+n_{1}^{*}}\left\{\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+n_{1}^{*}\cdot\left(\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\widehat{\Delta}_{1}\right)\right\}.$ (2.5)
Since $(\bm{X}_{i}^{*},Y_{i}^{*})_{i=1}^{n_{1}^{*}}$ are not observable, it is impossible to calculate $\widehat{\Delta}_{1}$ with the raw data. As formalized through explicit assumptions introduced later, we can estimate the bias from the available data in the majority group as follows. First, randomly partition the majority indices into a generation subgroup $\mathcal{S}_{0g}$ and a correction subgroup $\mathcal{S}_{0c}$ with corresponding sizes $n_{0g}$ and $n_{0c}$, respectively. In practice, we recommand chossing $n_{0g}\approx n_{1}$ and therefore $n_{0c}\approx n_{0}-n_{1}$. Next, using samples from the generation subgroup $(\bm{X}_{i},Y_{i})_{i\in\mathcal{S}_{0g}}$, generate $\tilde{n}_{0}$ synthetic samples $(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ by the same synthetic generator, where $\tilde{n}_{0}\approx n_{0}-n_{1}$ and $\tilde{Y}_{i}^{(0)}=0$ for $i=1,\ldots,\tilde{n}_{0}$. Then consider the population bias of the loss function from the true majority distribution $\mathcal{P}_{0}$ to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ by
$\Delta_{0}=\mathbb{E}_{\bm{X}\sim\mathcal{P}_{0}}\{\ell_{f}(\bm{X},0)\}-\mathbb{E}_{\tilde{\bm{X}}\sim\tilde{\mathcal{P}}_{0}}\{\ell_{f}(\tilde{\bm{X}},0)\}.$ (2.6)
Finally, obtain the sample loss bias for the majority group using the majority synthetic samples and the correction subsamples by
$\widehat{\Delta}_{0}=\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},Y_{i})-\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},\tilde{Y}_{i}^{(0)}).$ (2.7)
Note that all elements for calculating $\widehat{\Delta}_{0}$ are available, and $\widehat{\Delta}_{0}$ is constructed the same way as $\widehat{\Delta}_{1}$. Suppose there exists a measurable mapping $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ such that the minority distribution $\mathcal{P}_{1}$ is well approximated by the pushforward measure $(\mathcal{P}_{0})_{\#T}$, up to a controlled distributional error. Furthermore, assume that this transformation structure is approximately preserved under the corresponding synthetic distributions. This assumption is formally elaborated in Section 3.1. Given the above properties, it is possible to estimate the unobservable $\widehat{\Delta}_{1}$ using the data-driven estimator $\widehat{\Delta}_{0}$. To obtain a computable surrogate of the balanced loss, we further assume that the number of unobserved minority samples $n_{1}^{*}$ is matched by the number of generated synthetic minority samples, i.e., $n_{1}^{*}\approx\tilde{n}_{1}$. Thus, by modifying the balanced loss $L^{\rm bal}(f)$ in (2.5), we propose the following bias-correction loss function $L^{\rm bc}(f)$:
$\begin{split}L^{\rm bc}(f)=\frac{1}{n+\tilde{n}_{1}}\bigg[\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\tilde{n}_{1}\cdot\bigg\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\widehat{\Delta}_{0}\bigg\}\bigg].\end{split}$ (2.8)
The majority bias correction term $\widehat{\Delta}_{0}$ captures the loss function bias induced by the discrepancy between the true and synthetic distributions from the majority group. Under mild assumptions, $\widehat{\Delta}_{0}$ is a good representation of $\widehat{\Delta}_{1}$, the loss function bias from the minority group, up to a fixed bias-transfer error and sampling fluctuations, as illustrated in Section 3.1. With this property, the term $\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},\tilde{Y}_{i}^{(1)})+\widehat{\Delta}_{0}$ is regarded as the average loss from the unobserved minority samples after correcting the synthetic bias. Consequently, the bias-corrected loss $L^{\rm bc}$ represents a valid average loss from a roughly balanced dataset. Finally, we can find the prediction function by minimizing the bias-corrected loss:
$\widehat{f}^{\rm bc}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{\rm bc}(f).$ (2.9)
The process of bias correction for imbalanced classification is summarized in Algorithm 1. See Figure 1 for an illustration.
Algorithm 1 Bias Correction for Imbalanced Classification
0: Imbalanced data $(\bm{X}_{i},Y_{i})_{i=1}^{n}$, prediction function class $\mathcal{F}$, loss function $\ell_{f}$, synthetic generator $\mathcal{G}$, minority synthetic size $\tilde{n}_{1}$, majority synthetic size $\tilde{n}_{0}$ and generation size $n_{0g}$.
1: Minority augmentation: Generate $\tilde{n}_{1}$ synthetic minority samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$ via generator $\mathcal{G}$ trained on the minority group $(\bm{X}_{i})_{Y_{i}=1}$.
2: Partition the majority index set $\mathcal{S}_{0}=\{i:Y_{i}=0\}$ into a generation set $\mathcal{S}_{0g}$ and a correction set $\mathcal{S}_{0c}$ with corresponding sizes $n_{0g}$ and $n_{0c}=n_{0}-n_{0g}$.
3: Generate $\tilde{n}_{0}$ synthetic majority samples $(\tilde{\bm{X}}_{i}^{(0)})_{i=1}^{\tilde{n}_{0}}$ via generator $\mathcal{G}$ trained on the generation set $(\bm{X}_{i})_{i\in\mathcal{S}_{0g}}$.
4: Compute the empirical majority bias $\widehat{\Delta}_{0}$ according to Equation (2.7).
5: Form the bias-corrected loss $L^{\rm bc}(f)$ as defined in Equation (2.8).
6: Obtain the predictor $\widehat{f}^{\rm bc}$ by (2.9).
6: The prediction function $\widehat{f}^{\rm bc}:\mathbb{R}^{d}\to(0,1)$.
2.2 Multi-Task Imbalanced Learning
In this subsection, we focus on applying bias correction techniques to datasets involving multiple related tasks, a scenario commonly addressed by multi-task learning (MTL). MTL is a machine learning paradigm where multiple related tasks are learned simultaneously, enabling the model to leverage shared information and learn a common robust representation (Caruana, 1997; Zhang and Yang, 2021). For example, in genomic studies, researchers analyze gene expression data from different regional populations to identify genetic markers for specific diseases such as Alzheimer’s disease (Zhang and Shen, 2011). In this case, each regional population is regarded as a separate learning task. While the goal of identifying Alzheimer’s disease is shared, the genetic and environmental differences between populations lead to unique data distributions. This makes it necessary to utilize a multi-task learning framework to leverage the common structure. However, the number of individuals with the disease is typically smaller than the number of healthy individuals, creating a within-task imbalance problem (Wu et al., 2018; Guo et al., 2025). We apply the bias correction procedure to such imbalanced MTL problems to improve the predictive performance by leveraging information from all tasks.
Consider datasets from $K$ learning tasks and for each task $k=1,\ldots,K$, there are $n_{k}$ samples of covariates $\bm{X}_{ki}\in\mathbb{R}^{d}$ and class labels $Y_{ki}\in\{0,1\}$ for $i=1,\ldots,n_{k}$. Under the imbalanced setting, the class labels are imbalanced within each task such that the marginal probability $\pi_{k1}=\mathbb{P}(Y_{ki}=1)<1/2$. The dependence structure of the class labels on the covariates is captured by the following Bernoulli model:
$\mathbb{P}(Y_{ki}=1\mid\bm{X}_{ki}=\bm{x})=\sigma(\bm{x}^{\top}\bm{B}\bm{\alpha}_{k})\quad\text{ for }i=1,\ldots,n_{k},$ (2.10)
where $\sigma(t)=1/(1+\exp(-t))$ denotes the logistic function, $\bm{B}\in\mathbb{R}^{d\times r}$ denotes the shared coefficient matrix across $K$ tasks and $\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K}\in\mathbb{R}^{r}$ are task-specific latent loading vectors. Denote $\bm{M}=\bm{B}(\bm{\alpha}_{1},\ldots,\bm{\alpha}_{K})\in\mathbb{R}^{d\times K}$ as the unknown coefficient matrix with ${\rm rank}(\bm{M})=r$. Consider the task-specific coefficient vector $\bm{\beta}_{k}=\bm{B}\bm{\alpha}_{k}\in\mathbb{R}^{d}$ as the $k$th column of $\bm{M}$ for $k=1,\ldots,K$. Our goal is to learn the left singular vector space of the shared matrix $\bm{B}$.
For any $\bm{\beta}\in\mathbb{R}^{d}$, the prediction function is provided by $f(\bm{x})=\sigma(\bm{x}^{\top}\bm{\beta})$ and denote the loss function as
$\ell_{f}(\bm{x},y)=\ell(\bm{x},y;\bm{\beta})=-y\log(\sigma(\bm{x}^{\top}\bm{\beta}))-(1-y)\log(1-\sigma(\bm{x}^{\top}\bm{\beta})).$
For each task, we obtain the estimation $\widehat{\bm{\beta}}_{k}^{\rm raw}$ from $(\bm{X}_{ki},Y_{ki})_{i=1}^{n_{k}}$ by minimizing the loss function from the raw data $L^{\rm raw}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta}).$ Consider the minority synthetic samples $(\tilde{\bm{X}}_{ki}^{(1)})_{i=1}^{\tilde{n}_{k1}}$ and majority synthetic samples $(\tilde{\bm{X}}_{ki}^{(0)})_{i=1}^{\tilde{n}_{k0}}$. We can also obtain synthetic and bias-corrected loss functions by
$L^{\rm syn}(\bm{\beta})=$ $\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\sum_{i=1}^{\tilde{n}_{k1}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})\right],$
$L^{\rm bc}(\bm{\beta})=$ $\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\tilde{n}_{k1}\cdot\left\{\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})+\widehat{\Delta}_{k0}\right\}\right],$
where $\widehat{\Delta}_{k0}$ denotes the bias correction term from the majority samples in the correction set $\mathcal{S}_{kc}$ and majority synthetic samples:
$\widehat{\Delta}_{k0}=$ $\frac{1}{|\mathcal{S}_{kc}|}\sum_{i\in\mathcal{S}_{kc}}\ell(\bm{X}_{ki},0;\bm{\beta})-\frac{1}{\tilde{n}_{k0}}\sum_{i=1}^{\tilde{n}_{k0}}\ell(\tilde{\bm{X}}_{ki}^{(0)},0;\bm{\beta}).$
Denote $\widehat{\bm{\beta}}_{k}^{\rm syn}$ and $\widehat{\bm{\beta}}_{k}^{\rm bc}$ as the minimizers of the synthetic loss $L^{\rm syn}$ and bias-corrected loss $L^{\rm bc}$, respectively.
Next, we consider the estimation of the left singular matrix $\bm{U}$ from any coefficient estimators $(\widehat{\bm{\beta}}_{k})_{k=1}^{K}$. First, collect all estimators into $\widehat{\bm{M}}=(\widehat{\bm{\beta}}_{1},\dots,\widehat{\bm{\beta}}_{K})\in\mathbb{R}^{d\times K}$. Next, conduct eigendecomposition of $\widehat{\bm{M}}\widehat{\bm{M}}^{T}$ such that $\widehat{\bm{M}}\widehat{\bm{M}}^{T}=\widehat{\bm{U}}^{\prime}\widehat{\bm{\Lambda}}\widehat{\bm{U}}^{\prime\top}$, where $\widehat{\bm{U}}^{\prime}\in\mathbb{R}^{d\times d}$ is an orthonormal eigenvector matrix satisfying $(\widehat{\bm{U}^{\prime}})^{\top}\widehat{\bm{U}}^{\prime}=\bm{I}_{d}$ and $\widehat{\bm{\Lambda}}={\rm diag}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{d})$ is a diagonal eigenvalue matrix with decreasing eigenvalues $\widehat{\lambda}_{1}\geq\cdots\geq\widehat{\lambda}_{d}$. The rank of the latent embedding matrix is estimated using an eigenvalue ratio criterion that has been widely adopted for determining the number of latent factors in high-dimensional spectral and factor models (Ahn and Horenstein, 2013; Fan et al., 2019; Wang et al., 2026). Specifically, the rank is selected by maximizing the eigenvalue ratio such that $\widehat{r}=\arg\max_{1\leq r\leq d_{-}}\widehat{\lambda}_{r}/\widehat{\lambda}_{r+1}$, where $d_{-}<d$ is a constant to avoid numerical instability caused by extremely small eigenvalues. Finally, take the first $\widehat{r}$ columns of $\widehat{\bm{U}}^{\prime}$ to obtain the estimated shared embedding matrix $\widehat{\bm{U}}=\widehat{\bm{U}}^{\prime}_{1:\widehat{r}}$. Applying the same construction to different coefficient estimators yields the corresponding latent embedding matrix estimators. In particular, using the raw estimator $\widehat{\bm{\beta}}_{k}^{\rm raw}$, the synthetic estimator $\widehat{\bm{\beta}}_{k}^{\rm syn}$, and the bias-corrected estimator $\widehat{\bm{\beta}}_{k}^{\rm bc}$ in place of $\widehat{\bm{\beta}}_{k}$ results in $\widehat{\bm{U}}^{\rm raw}$, $\widehat{\bm{U}}^{\rm syn}$, and $\widehat{\bm{U}}^{\rm bc}$, respectively.
Suppose we are then provided with samples from a new task $(\bm{X}_{K+1,i},Y_{K+1,i})_{i=1}^{n_{K+1}}$ from the following Bernoulli model with the same shared structure:
$\mathbb{P}(Y_{K+1,i}=1\mid\bm{X}_{K+1,i}=\bm{x})=\sigma(\bm{x}^{\top}\bm{B}\bm{\alpha}_{K+1}).$
The estimated shared embedding matrix $\widehat{\bm{U}}$ helps us to estimate the coefficient in a lower dimension $\widehat{r}$ rather than $d$. To obtain the estimation, we can first project the covariates into a lower-dimensional embedding subspace by $\widehat{\bm{Z}}_{K+1,i}=\widehat{\bm{U}}^{\top}\bm{X}_{K+1,i}$ for each $i=1,\dots,n_{K+1}$. Next, obtain $\widehat{\bm{\theta}}_{K+1}$ which minimizes the loss function, for example, $L^{\rm raw}(\bm{\theta})$ on the dataset $(\widehat{\bm{Z}}_{K+1,i},Y_{K+1,i})_{i=1}^{n_{K+1}}$. Note that the empirical loss function $L^{\rm raw}$ can be replaced by $L^{\rm syn}$ and $L^{\rm bc}$ depending on the imbalance of task $K+1$. Then project $\widehat{\bm{\theta}}_{K+1}$ back to the coefficient space by $\widehat{\bm{\beta}}_{K+1}=\widehat{\bm{U}}\widehat{\bm{\theta}}_{K+1}$ and obtain the prediction function $\widehat{f}_{K+1}(\bm{x})=\sigma(\bm{x}^{\top}\widehat{\bm{\beta}}_{K+1})$. By substituting $\widehat{\bm{U}}$ with the above $\widehat{\bm{U}}^{\rm raw}$, $\widehat{\bm{U}}^{\rm syn}$ and $\widehat{\bm{U}}^{\rm bc}$, we can obtain the corresponding coefficient estimators and prediction function. In Section 3.2, we provide the theoretical results of the coefficient and embedding matrix estimations of the three methods and show the conditions under which the bias correction procedure outperforms the synthetic procedure.
2.3 Average Treatment Effect Estimation
Our proposed methodology has an application to average treatment effect (ATE) estimation, one fundamental problem in causal inference (Rubin, 1974; Lunceford and Davidian, 2004). In this context, imbalanced data often arises when the number of individuals receiving the treatment is significantly smaller than the number of individuals receiving the control. This imbalance makes the estimation of the ATE, more specifically, the expected outcome for the minority group, less reliable due to the limited sample size. This directly degrades the credibility and robustness of the final ATE estimate. Consequently, addressing this imbalanced data is essential for accurate causal inference.
Suppose $Y(1)$ and $Y(0)$ are the potential responses under treatment $Z=1$ and control $Z=0$. The observed response is a function of the potential responses and the treatment indicator: Y=ZY(1)+(1Z)Y(0)Y=ZY(1)+(1-Z)Y(0). Then the ATE is defined as $\tau=\mathbb{E}\{Y(1)\}-\mathbb{E}\{Y(0)\}$.
Consider i.i.d. observations $(\bm{X}_{i},Y_{i},Z_{i})_{i=1}^{n}$, where
$Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0),\qquad i=1,\ldots,n,$
$\bm{X}_{i}\in\mathbb{R}^{d}$ is a vector of covariates, and $Z_{i}\in\{0,1\}$ is the treatment indicator. Suppose there are $n_{1}$ and $n_{0}$ samples from the treatment group and control group, respectively, and define the treated and control covariate indices as $\mathcal{S}_{1}=\{i:Z_{i}=1\}$ and $\mathcal{S}_{0}=\{i:Z_{i}=0\}$. In the imbalanced case where $n_{1}<n_{0}$, we aim to estimate the ATE augmented with synthetic data by the bias correction approach.
Consider the propensity score (Rosenbaum and Rubin, 1983), which is defined as the conditional probability of a sample receiving treatment given the corresponding covariate $\bm{X}_{i}=\bm{x}$: $e^{*}(\bm{x})=\mathbb{P}(Z_{i}=1\mid\bm{X}_{i}=\bm{x}).$ For the propensity score, suppose we have an estimating model denoted as $e(\bm{x})$. Similarly, for the conditional means of the responses given the covariates $\mu_{1}^{*}(\bm{x})=\mathbb{E}\{Y_{i}(1)\mid\bm{X}_{i}=\bm{x}\}$ under treatment and $\mu_{0}^{*}(\bm{x})=\mathbb{E}\{Y_{i}(0)\mid\bm{X}_{i}=\bm{x}\}$ under control, suppose we have the estimating models $\mu_{1}(\bm{x})$ and $\mu_{0}(\bm{x})$, respectively. We consider the augmented inverse propensity weighting (AIPW) estimators (Rubin, 1978; Glynn and Quinn, 2010):
$\widehat{\mu}_{1}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{Z_{i}\{Y_{i}-\mu_{1}(\bm{X}_{i})\}}{e(\bm{X}_{i})}+\mu_{1}(\bm{X}_{i})\right],$ (2.11)
$\widehat{\mu}_{0}^{\rm AIPW}=$ $\frac{1}{n}\sum_{i=1}^{n}\left[\frac{(1-Z_{i})\{Y_{i}-\mu_{0}(\bm{X}_{i})\}}{1-e(\bm{X}_{i})}+\mu_{0}(\bm{X}_{i})\right],$ (2.12)
$\widehat{\tau}^{\rm AIPW}=$ $\widehat{\mu}_{1}^{\rm AIPW}-\widehat{\mu}_{0}^{\rm AIPW}.$ (2.13)
With the observations $(\bm{X}_{i},Y_{i},Z_{i})_{i=1}^{n}$, we can first fit separate regression models of the responses on the covariates and obtain the coefficient estimators $\widehat{\bm{\beta}}_{1}$ and $\widehat{\bm{\beta}}_{0}$ under treatment and control, respectively. Let $\widehat{\mu}_{1}(\bm{x})=\bm{x}^{\top}\widehat{\bm{\beta}}_{1}$ and $\widehat{\mu}_{0}(\bm{x})=\bm{x}^{\top}\widehat{\bm{\beta}}_{0}$ be the estimated responses under treatment and control, respectively.
Next, we consider the propensity score estimation. Suppose we are interested in a loss function $\ell_{f}$, e.g., logistic loss, for a prediction function $f:\mathbb{R}^{d}\to(0,1)$. With the raw data, we can obtain the propensity score estimation $\widehat{f}^{\rm raw}$ by minimizing the empirical loss function $L^{\rm raw}$ in (2.1). Assume that there are $\tilde{n}_{1}$ synthetic minority covariate samples $\tilde{\bm{X}}_{i}^{(1)}$, then we can obtain the synthetic-augmented propensity score estimation $\widehat{f}^{\rm syn}$ by minimizing the $L^{\rm syn}$ in (2.2). Partition the control index set into a generation set $\mathcal{S}_{0g}$ and a correction set $\mathcal{S}_{0c}$. Applying the same synthetic generator to obtain $\tilde{n}_{0}$ control synthetic covariates $\tilde{\bm{X}}_{i}^{(0)}$ from the generation set, we can obtain the bias-corrected propensity score estimation $\widehat{f}^{\rm bc}$ by minimizing $L^{\rm bc}$ in (2.8).
Plugging the treatment and control models $\widehat{\mu}_{1}(\bm{x})$ and $\widehat{\mu}_{0}(\bm{x})$ as well as the propensity score estimation $\widehat{e}(\bm{x})\in\{\widehat{f}^{\rm raw}(\bm{x}),\widehat{f}^{\rm syn}(\bm{x}),\widehat{f}^{\rm bc}(\bm{x})\}$ into (2.13), the AIPW estimators $\widehat{\tau}^{\rm raw},\widehat{\tau}^{\rm syn}$ and $\widehat{\tau}^{\rm bc}$ can be derived corresponding to the propensity score estimators.
3 Theoretical Properties
3.1 Bias Correction for Risk Functions
In this subsection, we first propose an upper bound for the difference between the minority and majority bias correction terms, $|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|$. This result provides a theoretical guarantee for the construction of the bias-corrected loss function $L^{\rm bc}$ in (2.8). We then derive a lower bound for the minority bias correction term $\widehat{\Delta}_{1}$ of SMOTE, thereby illustrating that treating synthetic samples as real data may introduce bias. Finally, we present the properties of the bias-corrected predictor with respect to the balanced population risk function, demonstrating that the proposed procedure effectively leverages synthetic samples while ensuring strong performance on imbalanced data.
To establish finite-sample guarantees for the proposed bias-corrected loss and predictor, we impose a set of standard regularity conditions. These assumptions are introduced to control the behavior of the loss function, the complexity of the classifier class, and the dependence structure of synthetic samples. They are commonly adopted in empirical risk minimization and learning with synthetic or augmented data, and do not impose algorithm-specific restrictions.
We first impose a boundedness condition on the prediction function and the associated loss. This assumption ensures that the empirical risk and its bias correction terms are well-defined and uniformly controlled, which is essential for deriving concentration inequalities and uniform deviation bounds. Such boundedness conditions are standard in the analysis of classification methods with probabilistic outputs.
Assumption 3.1 (Bounded prediction and loss).
There exists a constant $\varepsilon\in(0,1/2)$ such that for all $f\in\mathcal{F}$ and all $\bm{x}\in\mathcal{X}$, $f(\bm{x})\in[\varepsilon,1-\varepsilon]$ and there exists a constant $M_{\varepsilon}>0$ such that $0\leq\ell_{f}(\bm{x},y)\leq M_{\varepsilon}$ for all $(\bm{x},y)$. For instance, for the binary cross-entropy loss $\ell_{f}(\bm{x},y)=-y\log f(\bm{x})-(1-y)\log(1-f(\bm{x}))$, we have $M_{\varepsilon}=-\log\varepsilon$.
Next, we control the richness of the classifier class through its induced loss class. Bounding the Rademacher complexity of the normalized loss functions allows us to quantify the uniform generalization error of empirical risk minimization over $\mathcal{F}$. This assumption provides a standard measure of function class complexity and plays a central role in our finite-sample analysis.
Assumption 3.2 (Function class complexity).
Let $\mathcal{L}=\{(\bm{x},y)\mapsto\ell_{f}(\bm{x},y)/M_{\varepsilon}:f\in\mathcal{F}\}$. The Rademacher complexity of a function class $\mathcal{L}$ with sample size $m\geq 1$ is defined as
$\mathfrak{R}_{m}(\mathcal{L}):=\mathbb{E}_{\bm{Z}_{1:m},\,\bm{\sigma}}\left[\sup_{g\in\mathcal{L}}\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\,g(\bm{Z}_{i})\right],$
where $\bm{Z}_{1},\ldots,\bm{Z}_{m}$ are i.i.d. observations with $\bm{Z}_{i}=(\bm{X}_{i},Y_{i})$, $\sigma_{1},\ldots,\sigma_{m}$ are i.i.d. Rademacher random variables satisfying $\mathbb{P}(\sigma_{i}=1)=\mathbb{P}(\sigma_{i}=-1)=1/2$, and $(\sigma_{i})_{i=1}^{m}$ are independent of $(\bm{Z}_{i})_{i=1}^{m}$. There exists a constant $C_{\mathcal{F}}>0$ such that for all $m\geq 1$, the Rademacher complexity satisfies $\mathfrak{R}_{m}(\mathcal{L})\leq C_{\mathcal{F}}/\sqrt{m}$.
We now formalize the stochastic structure of the synthetic samples. Rather than assuming unconditional independence, we introduce a conditional independence framework in which synthetic data are generated from a random generator $\mathcal{G}$ and are independent of the evaluation samples given $\mathcal{G}$. This formulation accommodates many synthetic data generation mechanisms, including resampling and model-based procedures, while preserving tractable theoretical analysis.
Assumption 3.3 (Synthetic conditional independence).
There exists a random object $\mathcal{G}$ (the synthetic data generator) such that conditional on $\mathcal{G}$, the synthetic samples $\{\tilde{\bm{X}}_{i}^{(1)}\}_{i=1}^{\tilde{n}_{1}}$ are i.i.d. from $\tilde{\mathcal{P}}_{1}$ and $\{\tilde{\bm{X}}_{i}^{(0)}\}_{i=1}^{\tilde{n}_{0}}$ are i.i.d. from $\tilde{\mathcal{P}}_{0}$. Moreover, conditional on $\mathcal{G}$, the synthetic samples are independent of the observed samples used to evaluate $L^{\rm bc}$ and $\widehat{\Delta}_{0}$.
Finally, to characterize the structural relationship between the minority and majority populations, we introduce an alignment assumption based on distributional transformations. This assumption is imposed at an abstract level to quantify the degree of distributional similarity between populations, rather than to require an explicit or computable transformation in practice. It serves as an analytical device for controlling the discrepancy between bias correction terms across populations and for formalizing the effect of synthetic samples under distributional shift.
Assumption 3.4.
Consider the support of covariates $\mathcal{X}\subseteq\mathbb{R}^{d}$. Suppose the following conditions are satisfied uniformly for all $f\in\mathcal{F}$:
  1. (A1)
    Distribution transformation. There exists a measurable function $T:\mathcal{X}\to\mathcal{X}$ and a constant $\varepsilon_{T}$ such that
    $\mathcal{W}_{1}(\mathcal{P}_{1},(\mathcal{P}_{0})_{\#T})\leq\varepsilon_{T},\quad\text{and }\quad\mathcal{W}_{1}(\tilde{\mathcal{P}}_{1},(\tilde{\mathcal{P}}_{0})_{\#T})\leq\varepsilon_{T}\quad\text{a.s. conditional on }\mathcal{G},$
    where $(\cdot)_{\#T}$ denotes the pushforward distribution induced by transformation $T$, and $\mathcal{W}_{1}(\cdot,\cdot)$ represents the Wasserstein–1 distance.
  2. (A2)
    Lipschitz smoothness of the transformed loss. There exists constant $L_{\ell}>0$ such that for all $\bm{x}_{1},\bm{x}_{2}\in\mathcal{X}$ and $y\in\{0,1\}$,
    $|\ell_{f}(T(\bm{x}_{1}),y)-\ell_{f}(T(\bm{x}_{2}),y)|\leq$ $L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|_{2},$
    $|\ell_{f}(\bm{x}_{1},y)-\ell_{f}(\bm{x}_{2},y)|\leq$ $L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|_{2}.$
  3. (A3)
    Transformation bound for the loss function. There exists $\varepsilon_{h}$ such that
    $|\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)|\leq\varepsilon_{h},\quad\text{ and }\quad|\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\bm{X},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\bm{X},0)|\leq\varepsilon_{h}.$
By Assumption 3.4 (A1), the minority distribution $\mathcal{P}_{1}$ is approximately characterized as a transformation of the majority distribution through a mapping $T$, in the sense that the pushforward measure $(\mathcal{P}_{0})_{\#T}$ is close to $\mathcal{P}_{1}$ in $\mathcal{W}_{1}$ distance up to $\varepsilon_{T}$. Assumption 3.4 (A1) is imposed as a structural alignment condition rather than a requirement for an explicitly known or computable transformation. The transformation $T$ is introduced solely for theoretical analysis to characterize the degree of distributional similarity between majority and minority populations. In practice, our method does not require estimating or specifying $T$, and the validity of the results depends only on the existence of such a transformation with small Wasserstein discrepancy. For the synthetic distributions, the same transformation structure is approximately preserved: $\tilde{\mathcal{P}}_{1}$ remains close in $\mathcal{W}_{1}$ distance to $(\tilde{\mathcal{P}}_{0})_{\#T}$, with the discrepancy controlled by the same tolerance $\varepsilon_{T}$. This ensures that the synthetic generator retains the transformation structure between the majority and minority distributions. (A2) ensures the Lipschitz continuity of the loss and the transformed loss. Small covariate perturbations change the loss after transformation $T$ by at most $L_{\ell}$ times the perturbation size. The stability guarantees that small input fluctuations do not lead to extremely large changes in loss. (A3) further bounds the discrepancy between the minority loss and the majority loss. This condition requires that the loss function $\ell_{f}$ is roughly symmetric in expectation under the majority distribution $\mathcal{P}_{0}$ and the minority distribution $\mathcal{P}_{1}$ up to an error bound $\varepsilon_{h}$. This property also extends to the synthetic majority distribution $\tilde{\mathcal{P}}_{0}$ and its transformation $(\tilde{\mathcal{P}}_{0})_{\#T}$. It ensures that the transformation does not introduce excessive bias into loss evaluation across the two distributions. Under these assumptions, we now establish a high-probability upper bound for $|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|$. Denote $\varepsilon_{\rm BT}=2L_{\ell}\varepsilon_{T}+2\varepsilon_{h}$ as the bias-transfer error.
Proposition 3.1.
Suppose Assumptions 3.1-3.4 hold and there exists $0<c_{1}<c_{2}<1/2$ such that $c_{1}\leq n_{1}/n\leq c_{2}$. Then for any $f\in\mathcal{F}$ and any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|\leq\varepsilon_{\rm BT}+M_{\varepsilon}\sqrt{\frac{\log(8/\alpha)}{2}}\left(\frac{1}{\sqrt{n_{1}^{*}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right).$
Specifically, when $n_{1}^{*}=O(n_{0}-n_{1})$, $\tilde{n}_{1}=n_{1}^{*}=O(n_{0}-n_{1})$, $n_{0c}=n_{1}^{*}=O(n_{0}-n_{1})$ and $\tilde{n}_{0}=n_{1}^{*}=O(n_{0}-n_{1})$, the result can be written as follows: For any $\alpha>0$ and some constant $C<\infty$, with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|\leq\varepsilon_{\rm BT}+C\sqrt{\frac{\log(8/\alpha)}{n_{0}-n_{1}}}.$
Proposition 3.1 establishes an upper bound on the discrepancy between the majority and minority groups. Notably, This upper bound does not directly depend on how accurately the synthetic distribution $\tilde{\mathcal{P}}_{0}$ recovers the true distribution $\mathcal{P}_{0}$. Instead, noting that $\varepsilon_{\rm BT}=2\varepsilon_{h}+2L_{\ell}\varepsilon_{T}$, the discrepancy is governed by three components: the expected loss gap $\varepsilon_{h}$, the transformation approximation error $\varepsilon_{T}$, and the perturbation fluctuation of order $O((n_{0}-n_{1})^{-1/2})$. Since the class probabilities $\pi_{1}=\mathbb{P}(Y=1)$ and $\pi_{0}=\mathbb{P}(Y=0)$ are fixed constants, the effective sample size governing this stochastic fluctuation is of the same order as the total sample size $n$. This result implies that even when the synthetic generator yields a poor approximation to the true minority distribution, the additional error introduced by bias correction remains within the same well-defined bound. In practice, this means that the procedure remains stable and effective under suboptimal synthetic generators, ensuring the reliability of the bias-corrected estimator.
A key challenge with synthetic oversampling methods such as SMOTE is that the synthetic distribution $\tilde{\mathcal{P}}_{1}$ of the minority class does not perfectly match the true distribution $\mathcal{P}_{1}$. This discrepancy inevitably introduces a bias in the empirical risk. Here we show two complementary results: (i) SMOTE introduces a non-negligible bias in the minority class, for which we establish a population and empirical lower bound; and (ii) by applying a bias-correction procedure, we can upper bound the corrected error in terms of the distribution discrepancy between $\mathcal{P}_{1},\mathcal{P}_{0}$ and their synthetic counterparts.
Assumption 3.5.
  1. (A1)
    The loss $\ell_{f}(\bm{x},y)$ is Lipschitz in $\bm{x}$ with constant $L_{\ell}$ uniformly over $y\in\{0,1\}$, i.e., $|\ell_{f}(\bm{x}_{1},y)-\ell_{f}(\bm{x}_{2},y)|\leq L_{\ell}\|\bm{x}_{1}-\bm{x}_{2}\|_{2}$ for any $\bm{x}_{1},\bm{x}_{2}\in\mathbb{R}^{d}$ and $y\in\{0,1\}$.
  2. (A2)
    $\mathcal{P}_{1}$ is supported on a bounded set $B(0,R)\subseteq\mathbb{R}^{d}$ and has a density $f_{1}$ satisfying $0<C_{1}\leq f_{1}(\bm{x})\leq C_{2}<\infty$ for all $\bm{x}\in{\rm supp}(\mathcal{P}_{1})$.
  3. (A3)
    There exists a constant $C_{3}>0$ such that the following holds for fixed $K$. Let $(\bm{X}_{i})_{i=1}^{n_{1}}$ be an i.i.d. sample drawn from $\mathcal{P}_{1}$. Then for any sample $\bm{x}_{1}$, any $\bm{x}_{2}$ among its $K$ nearest neighbors within the samples, and any $u\in[0,1]$,
    $\ell_{f}(\bm{x}_{1}+u(\bm{x}_{2}-\bm{x}_{1}),1)-\ell_{f}(\bm{x}_{1},1)\geq C_{3}u\|\bm{x}_{2}-\bm{x}_{1}\|_{2}.$
Assumption 3.5 (A1) guarantees that the loss function $\ell_{f}$ is Lipschitz continuous with respect to $\bm{x}$ uniformly over $y\in\{0,1\}$. (A2) ensures that the minority distribution has bounded support and a bounded density function on its support. (A3) holds, for example, if $\ell_{f}(\cdot,1)$ is differentiable on ${\rm supp}(\mathcal{P}_{1})$ and its gradient satisfies $(\bm{x}_{2}-\bm{x}_{1})^{\top}\nabla\ell_{f}(\bm{z},1)\geq C_{3}\|\bm{x}_{2}-\bm{x}_{1}\|_{2}$ for any $\bm{z}$ on the segment between $\bm{x}_{1}$ and $\bm{x}_{2}$.
The following proposition shows that when the minority synthetic samples are generated by SMOTE, the induced bias cannot vanish too quickly. Specifically, there is a lower bound that scales with $(K/n_{1})^{1/d}$, reflecting the discrepancy between $\mathcal{P}_{1}$ and $\tilde{\mathcal{P}}_{1}$.
Proposition 3.2.
Suppose Assumptions 3.1-3.5 hold, and the synthetic minority samples are i.i.d. generated from $\tilde{\mathcal{P}}_{1}$ via SMOTE with parameter $K$. Then there exists a constant $c_{1}>0$, depending on $(d,C_{1},C_{2},C_{3},R)$, such that for any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}|\geq c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}-M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}-M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}.$
Proposition 3.2 demonstrates that the bias introduced into the loss function by the synthetic samples is statistically non-negligible. This bias prevents the synthetic augmented loss function (2.2) from serving as a close substitution for the balanced loss function (2.5). This discrepancy between these loss functions leads to a noticeable difference between their corresponding minimizers, which will potentially reduce the performance of the trained classifier.
Beyond quantifying the bias induced by synthetic sampling, an important question is how such bias affects the learning procedure itself. The natural target is the population balanced risk, which is defined as
$L^{*}(f)=\frac{1}{2}\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]+\frac{1}{2}\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)].$ (3.1)
The population balanced risk function $L^{*}$ represents the optimal case where the two classes are balanced with equal probability, which eliminates the effects of overfitting from imbalanced data. Denote the corresponding population balanced risk minimizer as
$f^{*}=\underset{f\in\mathcal{F}}{\arg\min}\>L^{*}(f),$ (3.2)
where $\mathcal{F}=\{f:\mathcal{X}\to(0,1)\}$ represents the prediction function class. In practice, however, one does not have access to the population balanced loss $L^{*}$, but instead minimizes an empirical loss. With synthetic samples generated and the bias correction procedure, we have the empirical minimizer $\widehat{f}^{\rm bc}$ of the bias-corrected loss function (2.9). Next, we investigate the following question: how close is the empirical bias-corrected minimizer $\widehat{f}^{\rm bc}$ to the population balanced minimizer $f^{*}$?
To answer this question, we consider a uniform assumption, which requires that the gap between the bias terms for the minority and majority groups remains controlled. Suppose that Assumption 3.4 is satisfied for all prediction functions $f\in\mathcal{F}$. The basic idea indicates that the bias observed in the minority group can be transferred to the majority group up to a controlled error $\varepsilon_{\rm BT}$, where $\varepsilon_{\rm BT}=2\varepsilon_{h}+2L_{\ell}\varepsilon_{T}$. Intuitively, if $\varepsilon_{\rm BT}$ is small, then correcting for the minority bias using information from the majority group is reliable.
The following theorem then provides an upper bound on the excess population risk of the bias-corrected minimizer.
Theorem 3.1.
Suppose Assumptions 3.1-3.4 hold and $\tilde{n}_{1}/(n_{0}-n_{1})\to 1$. Then for any $\alpha\in(0,1)$, there exists constant $C$ depends only on $M_{\varepsilon},\pi_{0},\pi_{1}$, such that with probability at least $1-\alpha$,
$L^{*}(\widehat{f}^{\rm bc})-L^{*}(f^{*})\leq$ $\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\varepsilon_{\rm BT}+C(2C_{\mathcal{F}}+\sqrt{\log(10/\alpha)})\left\{\frac{1}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right\}.$
The bound in Theorem 3.1 shows that the excess population risk of the bias-corrected estimator is controlled by two types of terms:
  1. (i)
    A bias transfer term, $\varepsilon_{\rm BT}$, which measures the worst-case mismatch of the bias correction from the majority to the minority group.
  2. (ii)
    Complexity terms with the order of inverse square root of sample sizes, which quantify the statistical fluctuations from randomness of finite samples in each component dataset.
Thus, bias correction ensures that even though synthetic oversampling introduces a nontrivial distributional bias, the bias-corrected empirical risk minimizer achieves population risk close to the optimal $f^{*}$, up to statistical and transfer errors.
3.2 Bias Correction for Multi-Task Learning
In transfer learning with multiple sources, samples in each source $k=1,\ldots,K$ are independently drawn from the corresponding distribution and regression parameter $\bm{\beta}_{k}$. We study how synthetic augmentation and bias correction affect the accuracy of parameter estimation across tasks, and how these errors accumulate to the shared low-rank structure.
For each source $k$, we observe $n_{k}$ i.i.d. samples $(\bm{X}_{ki},Y_{ki})_{i=1}^{n_{k}}$ with $Y_{ki}\in\{0,1\}$ following the logistic model:
$\mathbb{P}(Y_{ki}=1\mid\bm{X}_{ki}=\bm{x})=\sigma(\bm{x}^{\top}\bm{\beta}_{k}),\quad\text{ where }\sigma(t)=1/(1+\exp(-t)).$
We impose the following assumptions.
Assumption 3.6.
For each source index $k\in\{1,\ldots,K\}$:
  1. (A1)
    There exists $R>0$ such that $\|\bm{X}_{ki}\|_{2}\leq R$ almost surely.
  2. (A2)
    The Fisher information at $\bm{\beta}_{k}$ is uniformly positive definite with
    $\underline{\kappa}_{k}\bm{I}_{d}\preceq\bm{H}_{k}:=\mathbb{E}[\sigma^{\prime}(\bm{X}_{k1}^{\top}\bm{\beta}_{k})\bm{X}_{k1}\bm{X}_{k1}^{\top}]\preceq\bar{\kappa}_{k}\bm{I}_{d}\quad\text{ for some constants }0<\underline{\kappa}_{k}\leq\bar{\kappa}_{k}<\infty.$
  3. (A3)
    The synthetic samples $\tilde{\bm{X}}^{(1)}_{ki}$ and $\tilde{\bm{X}}_{ki}^{(0)}$ are generated by a fixed mechanism i.i.d. conditional on the training data.
  4. (A4)
    For $y\in\{0,1\}$, the gradient of the loss function $g_{k,y}(\bm{x})=\nabla_{\bm{\beta}}\ell(\bm{x},y;\bm{\beta}_{k})$ is $L_{k}$-Lipschitz on ${\rm supp}(\mathcal{P}_{k,t})\cup{\rm supp}(\tilde{\mathcal{P}}_{k,t})$, where $\ell$ represents the logistic loss function.
The following theorem establishes nonasymptotic bounds for three types of estimations: the raw MLE $\widehat{\bm{\beta}}_{k}$, the synthetic augmented estimator $\tilde{\bm{\beta}}_{k}$, and the bias-corrected estimator $\tilde{\bm{\beta}}_{k}^{\rm bc}$. It also quantifies how these parameter errors propagate to the estimation of the shared low-rank structure.
Theorem 3.2.
Under Assumptions 3.1-3.4 and 3.6, for any $\alpha\in(0,1)$, there exist constants $C_{1},C_{2},C_{3}>0$ such that, with probability at least $1-\alpha$, the following properties hold simultaneously for each source $k$:
${\rm(i)}\quad$ $C_{1}\left(\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}\right)\leq\|\widehat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}\leq C_{1}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}}{\lambda_{\min}(\bm{H}_{k})}\right),$
${\rm(ii)}\quad$ $\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\leq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})+C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
$\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\geq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})-C_{2}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\right),$
${\rm(iii)}\quad$ $\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}\left(\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{{\rm BT}}+\varepsilon_{{\rm sampling},k}\right),$
where $w_{k}=\tilde{n}_{k}/(n_{k}+\tilde{n}_{k1})$ and
$\varepsilon_{{\rm sampling},k}=$ $C_{3}(2C_{\mathcal{F}}+\log(10d/\alpha))\left\{\frac{1}{\sqrt{n_{k1}}}+\frac{1}{\sqrt{n_{k0}}}+\frac{1}{\sqrt{n_{k0,c}}}+\frac{1}{\sqrt{\tilde{n}_{k0}}}+\frac{1}{\sqrt{\tilde{n}_{k1}}}\right\}.$
Furthermore, consider the $d\times K$ true and estimated matrices as $\bm{M}=[\bm{\beta}_{1},\cdots,\bm{\beta}_{K}]$, $\widehat{\bm{M}}^{\rm raw}=[\widehat{\bm{\beta}}_{1}^{\rm raw},\cdots,\widehat{\bm{\beta}}_{K}^{\rm raw}]$, $\widehat{\bm{M}}^{\rm syn}=[\widehat{\bm{\beta}}_{1}^{\rm syn},\cdots,\widehat{\bm{\beta}}_{K}^{\rm syn}]$ and $\widehat{\bm{M}}^{\rm bc}=[\widehat{\bm{\beta}}_{1}^{(\rm bc)},\cdots,\widehat{\bm{\beta}}_{K}^{(\rm bc)}]$. Let $\bm{U}$ be the matrix of leading $r$ left singular vectors of $\bm{M}$ and $D=\sigma_{r}(\bm{M})-\sigma_{r+1}(\bm{M})$ the spectral gap. Then, with the same probability,
${\rm(iv)}\quad$ $\|\sin\Theta(\widehat{\bm{U}}^{\rm raw},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(v)}\quad$ $\|\sin\Theta(\widehat{\bm{U}}^{\rm syn},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2},$
${\rm(vi)}\quad$ $\|\sin\Theta(\widehat{\bm{U}}^{\rm bc},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2}.$
The bounds illustrate the trade-offs among three estimators:
  1. (i)
    The raw MLE $\widehat{\bm{\beta}}_{k}^{\rm raw}$ converges at the standard parametric rate $n_{k}^{-1/2}$.
  2. (ii)
    The synthetic estimator $\widehat{\bm{\beta}}^{\rm syn}$ inherits an additional bias, which reflects the distributional mismatch between $\mathcal{P}_{k,0}$ and $\tilde{\mathcal{P}}_{k,0}$.
  3. (iii)
    The bias-corrected estimator $\widehat{\bm{\beta}}_{k}^{\rm bc}$ removes the bias at the cost of extra sampling fluctuations and the residual class-difference bias $\varepsilon_{\rm BT}$.
Finally, the subspace error bounds (iv)-(vi) show how these parameter errors accumulate in estimating the shared low-rank structure, with stability governed by the spectral gap $D$.
Remark 3.1.
When
$\varepsilon_{\rm BT}\leq\frac{2\pi_{0}}{\pi_{0}-\pi_{1}}\cdot\frac{\underline{\kappa}_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0}),$ (3.3)
the bias correction parameter estimator $\widehat{\bm{\beta}}_{k}^{\rm bc}$ has smaller errors than the synthetic augmented parameter estimator $\widehat{\bm{\beta}}^{\rm syn}$. In contrast, when (3.3) does not hold, the advantage of the bias correction procedure is not guaranteed. In addition, when $\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})=O_{p}(n_{k}^{-1/2})$, the lower bound for the synthetic-augmented estimator error $\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}$ in Theorem 3.2 becomes negative and thus can be replaced by zero. Theorem 3.2 shows that the bias correction performs well, especially for “bad” synthetic generators.
3.3 Average Treatment Effect Estimation
Let $W=(\bm{X},Y,Z)$ be the observed dataset with covariates $\bm{X}\in\mathbb{R}^{d}$, the treatment/control indicator $Z\in\{0,1\}$ and observed response Y=ZY(1)+(1Z)Y(0)Y=ZY(1)+(1-Z)Y(0). The average treatment effect is $\tau=\mathbb{E}[Y(1)]-\mathbb{E}[Y(0)]$. Denote the conditional treatment and control responses as $\mu_{1}^{*}(\bm{x})=\mathbb{E}[Y(1)\mid\bm{X}=\bm{x}]$ and $\mu_{0}^{*}(\bm{x})=\mathbb{E}[Y(0)\mid\bm{X}=\bm{x}]$, respectively. Let $e^{*}(\bm{x})=\mathbb{P}(Z=1\mid\bm{X}=\bm{x})$ be the propensity score. For any arbitrary functions $\mu_{1}(\cdot)$, $\mu_{0}(\cdot)$ and $e(\cdot)$, define
$\psi(W;\mu_{1},\mu_{0},e)=\left\{\mu_{1}(\bm{X})+\frac{Z(Y-\mu_{1}(\bm{X}))}{e(\bm{X})}\right\}-\left\{\mu_{0}(\bm{X})+\frac{(1-Z)(Y-\mu_{0}(\bm{X}))}{1-e(\bm{X})}\right\}.$
Then the augmented inverse propensity weighting estimator of ATE given $\widehat{\mu}_{1}$, $\widehat{\mu}_{0}$ and $\widehat{e}$ is provided by
[…]
Assumption 3.7.
  1. (A1)
    Identifiability. $(Y(1),Y(0))$ is independent of $Z$ conditional on $\bm{X}$ and $\eta\leq e^{*}(\bm{X})\leq 1-\eta$ almost surely for some $0<\eta<1$.
  2. (A2)
    Bounded moments. $|Y|\leq M$ and $\|\bm{X}\|_{2}\leq R$ for some constants $M,R>0$.
Theorem 3.3.
Under Assumption 3.7, denote $r_{a}=(\mathbb{E}|\widehat{\mu}_{a}(\bm{X})-\mu_{a}^{*}(\bm{X})|^{2})^{1/2}$ for $a\in\{0,1\}$ and $r_{e}=(\mathbb{E}|\widehat{e}(\bm{X})-e^{*}(\bm{X})|^{2})^{1/2}$. There exist constants $C_{0},C_{1}>0$ depending only on $\eta$ and $M$, such that with probability at least $1-\alpha$,
$|\widehat{\tau}^{\rm AIPW}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}}{\eta}(r_{1}+r_{0})r_{e}+C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{e}+r_{1}r_{0}\big).$
Corollary 3.1.
Under Assumptions 3.6 and 3.7, suppose the propensity score estimation is obtained from the bias-corrected coefficient estimator $\widehat{e}^{\rm bc}(\bm{x})=\sigma(\bm{x}^{\top}\widehat{\bm{\beta}}^{\rm bc})$, then the corresponding ATE estimator satisfies
$|\widehat{\tau}^{\rm AIPW,bc}-\tau|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}+\frac{C_{1}R}{\eta}(r_{1}+r_{0})r_{\beta}+C_{0}R\sqrt{\frac{\log(4/\alpha)}{2n}}\big((r_{1}+r_{0})r_{\beta}+r_{1}r_{0}\big),$
where $r_{\beta}$ represents the error bound for $\widehat{\bm{\beta}}^{\rm bc}$:
$r_{\beta}=\frac{1}{\underline{\kappa}_{k}}\left(\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{{\rm BT}}+\varepsilon_{{\rm sampling}}\right).$
Suppose that the treatment and control effect estimations $\widehat{\mu}_{1}(\cdot)$ and $\widehat{\mu}_{0}(\cdot)$ are obtained from the raw data in the treatment and control groups, respectively. For example, when the model is correctly specified for linear regression, the corresponding errors scale with the sample sizes, i.e., $r_{a}=O(n_{a}^{-1/2})$ for $a\in\{0,1\}$. Then according to Theorem 3.3, how close $\widehat{\tau}^{\rm AIPW}$ is to the true ATE $\tau$ depends on the propensity score estimation error $r_{e}$. Since the propensity score estimation is obtained from the coefficient estimation $\widehat{\bm{\beta}}$, we can have the following result: when (3.3) holds, the ATE estimation from the bias correction procedure $\widehat{e}^{\rm bc}(\bm{X})=\sigma(\bm{X}^{\top}\widehat{\bm{\beta}}^{\rm bc})$ has a smaller error than the synthetic augmented estimation from $\widehat{e}^{\rm syn}(\bm{X})=\sigma(\bm{X}^{\top}\widehat{\bm{\beta}}^{\rm syn})$. In contrast, when (3.3) does not hold, the advantage of the bias correction procedure is not guaranteed.
4 Data Analysis for MNIST Dataset
To evaluate the practical efficacy of our proposed framework, we apply it to the MNIST dataset (LeCun, 1998). We use Perturbed Sampling to generate synthetic data for a binary classification task where digit 1 or 4 is treated as the minority class in a five-digit subset. The results are detailed in Figure 2.
Figure 2: Seven metrics (Recall, Precision, Accuracy, Fowlkes-Mallows score (FM), F1-score, Matthews Correlation Coefficient (MCC), and Jaccard index) for the three methods applied to the MNIST dataset (digits 0-4). Results are shown for binary classification of digit 1 (top panel) and digit 4 (bottom panel), treated as minority classes.
Based on an evaluation in Figure 2, the perturbed sampling method demonstrates a clear performance hierarchy for classifying digits 1 and 4 as minority classes. While maintaining a similar precision score, the generation of synthetic data provides a foundational improvement, yielding superior results across all other metrics compared to the model trained exclusively on the raw, imbalanced data. The significant performance leap observed after applying bias correction reveals the substantial inherent bias introduced by the perturbed sampling technique. This bias correction step is not merely beneficial but critical, as it consistently produces the most accurate and reliable classifications. The results underscore that bias correction is an important step for mitigating distortion and achieving better model performance.
5 Discussions
We propose a novel bias correction framework for imbalanced classification with synthetic data augmentation, motivated by the observation that treating synthetic data as real data often introduces systematic bias due to imperfect data generation. The proposed method estimates and corrects the bias by leveraging majority-class data, where the majority samples are partitioned into generation and correction subsets, enabling bias estimation and correction through discrepancies between synthetic and real majority samples. Theoretical results establish the validity of the corrected bias and demonstrate that its guarantees do not explicitly rely on the discrepancy between the true and synthetic distributions. This property makes the approach robust even when suboptimal synthetic generators are used. Empirical studies, including simulations and experiments on the MNIST dataset, confirm the effectiveness of the proposed method.
This approach has several limitations: (i) The current theory characterizes bias through a scalar loss function, which may inadequately capture complex discrepancies in high-dimensional settings. (ii) When synthetic generators already produce highly accurate samples, the additional gains from bias correction are limited and may not justify the computational cost. (iii) While empirical results suggest that the method extends naturally to multi-class problems, a formal theoretical justification is currently lacking. Future work will focus on developing richer discrepancy measures beyond loss-based bias, establishing principled criteria to identify scenarios where bias correction is beneficial, and providing a rigorous theoretical extension to multi-class classification.
Acknowledgements
P. L. and A. R. Z. were partially supported by NIH Grant R01HL169347; Z. M. and A. R. Z. were partially supported by NIH Grant R01HL168940; A. R. Z. was also partially supported by NSF Grant CAREER 2203741. L.Z. was partially supported by NSF Grant CAREER 2340241 and Renaissance Philanthropy “AI for Math” Fund.
Disclosure statement
The authors declare no conflict of interest.
Appendix A Proof of main results
A.1 Proof of Proposition 3.1
Proof.
By the triangle inequality, we have
$|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|\leq|\Delta_{1}-\Delta_{0}|+|\widehat{\Delta}_{1}-\Delta_{1}|+|\widehat{\Delta}_{0}-\Delta_{0}|,$ (A.1)
where
$\Delta_{1}=\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1),\quad\Delta_{0}=\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\tilde{\bm{X}},0).$
Considering the transformation map $T$ by Assumption 3.4 (A1), $\mathbb{E}_{\mathcal{(}P_{0})_{\#T}}\ell_{f}(\bm{X},1)=\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(T(\bm{X}),1)$. Similarly, we also have $\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(T(\tilde{\bm{X}}),1)=\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)$. We can thus rewrite $\Delta_{1}$ as
$\Delta_{1}=$ $\big\{\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}\ell_{f}(\bm{X},1)\big\}+\big\{\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(T(\bm{X}),1)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(T(\tilde{\bm{X}}),1)\big\}$
$+\big\{\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\big\}.$
Denote $h(\bm{x}):=\ell_{f}(T(\bm{x}),1)-\ell_{f}(\bm{x},0)$, and we have
$\Delta_{1}-\Delta_{0}=$ $\big\{\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})\big\}+\big\{\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}\ell_{f}(\bm{X},1)\big\}$
$+\big\{\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\big\}.$ (A.2)
Under Assumption 3.4 (A2), $\ell_{f}(\bm{x},1)$ has Lipschitz constant $L_{\ell}>0$. By the Kantorovich-Rubinstein duality and Assumption 3.4 (A1),
$|\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)-\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}\ell_{f}(\bm{X},1)|\leq L_{\ell}\cdot\mathcal{W}_{1}(\mathcal{P}_{1},(\mathcal{P}_{0})_{\#T})\leq L_{\ell}\cdot\varepsilon_{T},$ (A.3)
$\big|\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}\ell_{f}(\tilde{\bm{X}},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\big|\leq L_{\ell}\cdot\mathcal{W}_{1}((\tilde{\mathcal{P}}_{0})_{\#T},\tilde{\mathcal{P}}_{1})\leq L_{\ell}\cdot\varepsilon_{T}.$ (A.4)
Assumption 3.4 (A3) guarantees that $h$ is bounded by $\varepsilon_{h}$ in expectation with respect to $\mathcal{P}_{0}$ and $\tilde{\mathcal{P}}_{0}$, then
$\big|\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})\big|\leq 2\varepsilon_{h}.$ (A.5)
Therefore, combining (A.1), (A.3), (A.4) and (A.5), we have
$|\Delta_{1}-\Delta_{0}|\leq 2\varepsilon_{h}+2L_{\ell}\cdot\varepsilon_{T}.$ (A.6)
Since $\ell_{f}\in[0,M_{\varepsilon}]$ by Assumption 3.1, and the samples in $\widehat{\Delta}_{1}$ and $\widehat{\Delta}_{0}$ are independent, by Hoeffding’s inequality, for any $t>0$,
$\mathbb{P}\left(\left|\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)\right|>t\right)\leq$ $2\exp\{-2n_{1}^{*}t^{2}/M_{\varepsilon}^{2}\},$
$\mathbb{P}\left(\left|\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}\ell_{f}(\tilde{\bm{X}},1)\right|>t\right)\leq$ $2\exp\{-2\tilde{n}_{1}t^{2}/M_{\varepsilon}^{2}\},$
$\mathbb{P}\left(\left|\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},0)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)\right|>t\right)\leq$ $2\exp\{-2n_{0c}t^{2}/M_{\varepsilon}^{2}\},$ (A.7)
$\mathbb{P}\left(\left|\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},0)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\tilde{\bm{X}},0)\right|>t\right)\leq$ $2\exp\{-2\tilde{n}_{0}t^{2}/M_{\varepsilon}^{2}\}.$ (A.8)
Let
$t_{1}=M_{\varepsilon}\sqrt{\frac{\log(8/\alpha)}{2n_{1}^{*}}},\quad t_{2}=M_{\varepsilon}\sqrt{\frac{\log(8/\alpha)}{2\tilde{n}_{1}}},\quad t_{3}=M_{\varepsilon}\sqrt{\frac{\log(8/\alpha)}{2n_{0c}}},\quad t_{4}=M_{\varepsilon}\sqrt{\frac{\log(8/\alpha)}{2\tilde{n}_{0}}}.$
A union bound over the four two-sided events yields that with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}-\Delta_{1}|\leq t_{1}+t_{2},\quad|\widehat{\Delta}_{0}-\Delta_{0}|\leq t_{3}+t_{4}.$ (A.9)
Therefore, combining (A.1), (A.6) and (A.9), we obtain the result: with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}-\widehat{\Delta}_{0}|\leq 2\varepsilon_{h}+2L_{\ell}\cdot\varepsilon_{T}+\sqrt{\frac{\log(8/\alpha)}{2}}\left(\frac{1}{\sqrt{n_{1}^{*}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right).$
This completes the proof of Proposition 3.1. ∎
A.2 Proof of Proposition 3.2
Proof.
We first show that there exists a constant $C>0$ such that
$\mathbb{E}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|_{2}\mid\bm{X}_{i})\geq C\left(\frac{K}{n_{1}}\right)^{1/d}.$ (A.10)
Denote $V_{d}(r)$ as the volume of a $d$-dimensional ball with radius $r$, then
$V_{d}(r)=\frac{\pi^{d/2}}{\Gamma(d/2+1)}r^{d}.$
For simplicity, let $v_{d}=V_{d}(1)=\pi^{d/2}/\Gamma(d/2+1)$ and thus $V_{d}(r)=v_{d}r^{d}$. Let $(\bm{X}_{i})_{i=1}^{n_{1}}$ be i.i.d. samples from $\mathcal{P}_{1}$ with density $f_{1}$ satisfying $C_{1}\leq f_{1}(\bm{x})\leq C_{2}$ for any $\bm{x}\in{\rm supp}(\mathcal{P}_{1})$. Fix $\bm{x}$ such that there exists $r_{0}>0$ with $B(\bm{x},r_{0})\subseteq{\rm supp}(\mathcal{P}_{1})$. For $r\leq r_{0}$, denote $N_{i}(r)=\sum_{j\neq i}I(\|\bm{X}_{j}-\bm{X}_{i}\|_{2}\leq r)$ as the neighbor count. Let $R_{i,k}=\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}$ be the $k$th nearest-neighbor distance of $\bm{X}_{i}$. Since $\mathcal{P}_{1}$ admits a continuous density, ties occur with probability zero and $N_{i}(R_{i,k})=k$ holds almost surely.
Conditioned on $\bm{X}_{i}$, $N_{i}(r)\sim{\rm Bin}(n_{1}-1,p_{i}(r))$, where $p_{i}(r)=\mathbb{P}(\|\bm{X}-\bm{X}_{i}\|_{2}\leq r\mid\bm{X}_{i})=\int_{B(\bm{X}_{i},r)}f_{1}(\bm{z}){\rm d}\bm{z}$. For $r\leq r_{0}$, the density bounds imply
$C_{1}V_{d}(r)\leq p_{i}(r)\leq C_{2}V_{d}(r),$
and hence the conditional mean $\mu_{i}(r)=\mathbb{E}(N_{i}(r)\mid\bm{X}_{i})=(n_{1}-1)p_{i}(r)$ satisfies
$(n_{1}-1)C_{1}V_{d}(r)\leq\mu_{i}(r)\leq(n_{1}-1)C_{2}V_{d}(r).$
For a fixed $k\in\{1,\ldots,K\}$, define
$r_{+}(k)=\left(\frac{2k}{(n_{1}-1)C_{1}v_{d}}\right)^{1/d},\quad r_{-}(k)=\left(\frac{k}{2(n_{1}-1)C_{2}v_{d}}\right)^{1/d}.$
Assume $n_{1}$ is large enough such that $r_{+}(k)\leq r_{0}$ and $r_{-}(k)\leq r_{0}$ for all $k\leq K$. Then
$\mu_{i}(r_{+}(k))\geq(n_{1}-1)C_{1}v_{d}r_{+}(k)^{d}=2k,$
$\mu_{i}(r_{-}(k))\leq(n_{1}-1)C_{2}v_{d}r_{-}(k)^{d}=k/2.$
The event $R_{i,k}>r_{+}(k)$ implies that $N_{i}(r_{+}(k))<k$. Then Chernoff’s lower-tail bounds for binomials give
$\mathbb{P}(R_{i,k}>r_{+}(k)\mid\bm{X}_{i})\leq\mathbb{P}(N_{i}(r_{+}(k))<k\mid\bm{X}_{i})\leq\exp(-c_{+}k)$
for some constant $c_{+}>0$. Similarly, we have
$\mathbb{P}(R_{i,k}<r_{-}(k)\mid\bm{X}_{i})\leq\mathbb{P}(N_{i}(r_{-}(k))\geq k\mid\bm{X}_{i})\leq\exp(-c_{-}k)$
for some constant $c_{-}>0$. Therefore, for some constant $c_{0}>0$, we have
$\mathbb{P}(r_{-}(k)\leq R_{i,k}\leq r_{+}(k)\mid\bm{X}_{i})\geq 1-2\exp(-c_{0}k).$
Moreover, since ${\rm supp}(\mathcal{P}_{1})\subseteq B(0,R)$, we have $R_{i,k}\leq 2R$ almost surely. Hence
$\mathbb{E}(R_{i,k}\mid\bm{X}_{i})$ $=\mathbb{E}\left\{R_{i,k}I(R_{i,k}\leq r_{+}(k))\mid\bm{X}_{i}\right\}+\mathbb{E}\left\{R_{i,k}I(R_{i,k}>r_{+}(k))\mid\bm{X}_{i}\right\}$
$\leq r_{+}(k)+2R\cdot\mathbb{P}(R_{i,k}>r_{+}(k)\mid\bm{X}_{i})$
$\leq r_{+}(k)+2R\exp(-c_{+}k).$
Similarly, we also obtain that
$\mathbb{E}(R_{i,k}\mid\bm{X}_{i})$ $=\mathbb{E}\left\{R_{i,k}I(R_{i,k}<r_{-}(k)\mid\bm{X}_{i}\right\}+\mathbb{E}\left\{R_{i,k}I(R_{i,k}\geq r_{-}(k)\mid\bm{X}_{i}\right\}$
$\geq r_{-}(k)\cdot\mathbb{P}(R_{i,k}\geq r_{-}(k)\mid\bm{X}_{i})$
$\geq r_{-}(k)(1-\exp(-c_{-}k)).$
Combining the two inequalities above and the definitions of $r_{+}(k),r_{-}(k)$ yields
$C_{-}(k/n_{1})^{1/d}\leq\mathbb{E}(R_{i,k}\mid\bm{X}_{i})\leq C_{+}(k/n_{1})^{1/d}+2R\exp(-c_{+}K),$
where the constants depend only on $d,C_{1},C_{2},R$ and $r_{0}$.
Let $\tilde{\bm{X}}_{j}=\bm{X}_{i}+U(\bm{X}_{i(k)}-\bm{X}_{i})$, where $U\sim{\rm Unif}(0,1)$ and $k\sim{\rm Unif}\{1,\ldots,K\}$, independent of the samples. Then
$\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|_{2}=U\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}=UR_{i,k},$
so
$\mathbb{E}(\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|_{2}\mid\bm{X}_{i})=$ $\mathbb{E}(U)\cdot\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}(R_{i,k}\mid\bm{X}_{i})$
$\geq$ $\frac{1}{2K}\sum_{k=1}^{K}C_{-}(k/n_{1})^{1/d}$
$\geq$ $C(K/n_{1})^{1/d},$
where the last equation holds since $K^{-1}\sum_{k=1}^{K}k^{1/d}\asymp K^{1/d}$. Taking expectation over $\bm{X}_{i}$ gives
$\mathbb{E}\|\tilde{\bm{X}}_{j}-\bm{X}_{i}\|\geq C(K/n_{1})^{1/d}.$ (A.11)
Denote $\mu_{1}=\mathbb{E}_{\bm{X}\sim\mathcal{P}_{1}}\{\ell_{f}(\bm{X},1)\}$ and $\tilde{\mu}_{1}=\mathbb{E}_{\bm{X}\sim\tilde{\mathcal{P}}_{1}}\{\ell_{f}(\bm{X},1)\}$. Since $\bm{X}_{1}^{*},\ldots,\bm{X}_{n_{1}^{*}}^{*}$ are iid generated from $\mathcal{P}_{1}$, with probability at least $1-\alpha$,
$\left|\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\mu_{1}\right|\leq M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}.$
Similarly, we also have
$\left|\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i},\tilde{Y}_{i})-\tilde{\mu}_{1}\right|\leq M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}.$
We have
$|\mu_{1}-\tilde{\mu}_{1}|=$ $|\mathbb{E}\{\ell_{f}(\bm{X},1)\}-\mathbb{E}\{\ell_{f}(\tilde{\bm{X}},1)\}|$
$\geq$ $|\mathbb{E}\{\ell_{f}(\bm{X}_{i},1)-\ell_{f}(\bm{X}_{i}+U_{i}(\bm{X}_{i(k)}-\bm{X}_{i}),1)\}|$
$\geq$ $C_{3}\cdot\mathbb{E}[U_{i}\cdot\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}]$ (A.12)
$=$ $C_{3}\cdot\mathbb{E}[U_{i}]\cdot\mathbb{E}\{\|\bm{X}_{i(k)}-\bm{X}_{i}\|_{2}\}$ (A.13)
$\geq$ $c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}$ (A.14)
for some constant $c_{1}>0$. Here, (A.12) is due to Assumption 3.5 (A3); (A.13) is due to the independence of $U_{i}$ and $\bm{X}_{i(k)}-\bm{X}_{i}$; (A.14) holds due to (A.11).
Then, with probability at least $1-\alpha$,
$|\widehat{\Delta}_{1}|\geq$ $\left|\mu_{1}-\tilde{\mu}_{1}\right|-\left|\frac{1}{n_{1}^{*}}\sum_{i=1}^{n_{1}^{*}}\ell_{f}(\bm{X}_{i}^{*},Y_{i}^{*})-\mu_{1}\right|-\left|\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i},\tilde{Y}_{i})-\tilde{\mu}_{1}\right|$
$\geq$ $c_{1}\left(\frac{K}{n_{1}}\right)^{1/d}-M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2n_{1}^{*}}}-M_{\varepsilon}\sqrt{\frac{\log(2/\alpha)}{2\tilde{n}_{1}}}.$
This completes the proof of Proposition 3.2. ∎
A.3 Proof of Theorem 3.1
Proof.
Note that
$L^{*}(\widehat{f}^{\rm bc})-L^{*}(f^{*})=$ $\left(L^{*}(\widehat{f}^{\rm bc})-L^{\rm bc}(\widehat{f}^{\rm bc})\right)+\left(L^{\rm bc}(\widehat{f}^{\rm bc})-L^{\rm bc}(f^{*})\right)+\left(L^{\rm bc}(f^{*})-L^{*}(f^{*})\right),$
and
$L^{\rm bc}(\widehat{f}^{\rm bc})-L^{\rm bc}(f^{*})\leq 0.$
We just need to show that for any prediction function $f\in\mathcal{F}$ and any $\alpha\in(0,1)$, with probability at least $1-\alpha$,
$\begin{split}&\sup_{f\in\mathcal{F}}|L^{\rm bc}(f)-L^{*}(f)|\\ \leq&\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{\rm BT}+\frac{C}{2}\cdot(2C_{\mathcal{F}}+\sqrt{\log(10/\alpha)})\left\{\frac{1}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right\}.\end{split}$ (A.15)
Recall that
$L^{\rm bc}(f)=$ $\frac{1}{n+\tilde{n}_{1}}\left[\sum_{i=1}^{n}\ell_{f}(\bm{X}_{i},Y_{i})+\tilde{n}_{1}\left\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)+\widehat{\Delta}_{0}\right\}\right],$
$L^{*}(f)=$ $\frac{1}{2}\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]+\frac{1}{2}\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)].$
First, rewrite $L^{\rm bc}(f)-L^{*}(f)$ as
$L^{\rm bc}(f)-L^{*}(f)=$ $\underbrace{\frac{n_{1}}{n+\tilde{n}_{1}}\cdot\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\ell_{f}(\bm{X}_{i},1)-\left(\frac{1}{2}-\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\right)\cdot\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]}_{(\rm I)}$
$+\underbrace{\frac{n_{0}}{n+\tilde{n}_{1}}\cdot\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\ell_{f}(\bm{X}_{i},0)-\frac{1}{2}\cdot\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)]}_{(\rm II)}$
$+\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\left\{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)+\Delta_{0}-\underbrace{(\widehat{\Delta}_{0}-\Delta_{0})}_{(\rm III)}-\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]\right\}.$
Note that $\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(T(\bm{X}),1)]=\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}[\ell_{f}(\bm{X},1)]$, and let $h(\bm{x})=\ell_{f}(T(\bm{x}),1)-\ell_{f}(\bm{x},0)$, we have
$\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)+\Delta_{0}-\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]$
$=$ $\underbrace{\frac{1}{\tilde{n}_{1}}\sum_{i=1}^{\tilde{n}_{1}}\ell_{f}(\tilde{\bm{X}}_{i}^{(1)},1)-\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]}_{(\rm IV)}+\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]$
$+\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)]-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}[\ell_{f}(\tilde{\bm{X}},0)]-\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(T(\bm{X}),1)]+\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}[\ell_{f}(\bm{X},1)]-\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]$
$=$ $({\rm IV})+\underbrace{\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]-\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}[\ell_{f}(\tilde{\bm{X}},1)]}_{(\rm V)}+\underbrace{\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}[\ell_{f}(\bm{X},1)]-\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]}_{(\rm VI)}$
$+\big(\mathbb{E}_{\tilde{\mathcal{P}}_{0}}[\ell_{f}(T(\tilde{\bm{X}}),1)]-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}[\ell_{f}(\tilde{\bm{X}},0)]\big)-\big(\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(T(\bm{X}),1)]-\mathbb{E}_{\mathcal{P}_{0}}[\ell_{f}(\bm{X},0)]\big)$
$=$ $({\rm IV})+({\rm V})+({\rm VI})+\underbrace{\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})-\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})}_{(\rm VII)}.$
Thus we have
$L^{\rm bc}(f)-L^{*}(f)=({\rm I})+({\rm II})+\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\left\{({\rm III})+({\rm IV})+({\rm V})+({\rm VI})+({\rm VII})\right\},$
and we aim to bound each term uniformly over $f\in\mathcal{F}$.
Lemma 1 (Uniform concentration for bounded losses).
Under Assumptions 3.1 and 3.2, for any i.i.d. sample $\bm{Z}_{1},\dots,\bm{Z}_{m}$ with $\bm{Z}_{i}=(\bm{X}_{i},Y_{i})$ and any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\sup_{f\in\mathcal{F}}\left|\frac{1}{m}\sum_{i=1}^{m}\ell_{f}(\bm{Z}_{i})-\mathbb{E}\ell_{f}(\bm{Z})\right|\leq M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{m}}+\sqrt{\frac{\log(2/\delta)}{2m}}\right).$ (A.16)
Proof of Lemma 1.
By Assumption 3.1, $0\leq\ell_{f}(\bm{z})\leq M_{\varepsilon}$ for all $\bm{z}$ and all $f\in\mathcal{F}$. Define $g_{f}(\bm{z})=\ell_{f}(\bm{z})/M_{\varepsilon}$ and $\mathcal{L}=\{g_{f}:f\in\mathcal{F}\}$, then $g_{f}(\bm{z})\in[0,1]$.
Define
$\Phi(\bm{Z}_{1:m})=\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}g(\bm{Z}_{i})-\mathbb{E}g(\bm{Z})\right|.$
We first control $\mathbb{E}\Phi(\bm{Z}_{1:m})$ by symmetrization and Rademacher complexity. Let $\bm{Z}_{1}^{\prime},\dots,\bm{Z}_{m}^{\prime}$ be i.i.d. copies independent of $\bm{Z}_{1:m}$. Then
$\mathbb{E}\Phi(\bm{Z}_{1:m})$ $=\mathbb{E}\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}g(\bm{Z}_{i})-\mathbb{E}g(\bm{Z})\right|$
$=\mathbb{E}\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}\left(g(\bm{Z}_{i})-\mathbb{E}g(\bm{Z}_{i}^{\prime})\right)\right|$
$\leq\mathbb{E}\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}\left(g(\bm{Z}_{i})-g(\bm{Z}_{i}^{\prime})\right)\right|.$
Let $\sigma_{1},\dots,\sigma_{m}$ be i.i.d. Rademacher random variables independent of all random variables. Conditioning on $\bm{Z}_{1:m}$ and $\bm{Z}_{1:m}^{\prime}$, the standard symmetrization step gives
$\mathbb{E}\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}\left(g(\bm{Z}_{i})-g(\bm{Z}_{i}^{\prime})\right)\right|=$ $\mathbb{E}\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}\left(g(\bm{Z}_{i})-g(\bm{Z}_{i}^{\prime})\right)\right|$
$\leq$ $2\mathbb{E}\sup_{g\in\mathcal{L}}\left|\frac{1}{m}\sum_{i=1}^{m}\sigma_{i}g(\bm{Z}_{i})\right|=2\mathfrak{R}_{m}(\mathcal{L}).$
Therefore $\mathbb{E}\Phi(\bm{Z}_{1:m})\leq 2\mathfrak{R}_{m}(\mathcal{L})$.
Next we apply McDiarmid’s inequality to $\Phi(\bm{Z}_{1:m})$. Fix $j\in\{1,\dots,m\}$ and replace $\bm{Z}_{j}$ by an independent copy $\bm{Z}_{j}^{\prime}$. Denote by $\bm{Z}_{1:m}^{(j)}$ the sample with only the $j$th coordinate replaced. Since $g(\bm{z})\in[0,1]$, for any $g\in\mathcal{L}$,
$\left|\frac{1}{m}\sum_{i=1}^{m}g(\bm{Z}_{i})-\frac{1}{m}\sum_{i=1}^{m}g\left(\bm{Z}_{i}^{(j)}\right)\right|\leq\frac{1}{m}.$
Taking supremum over $g\in\mathcal{L}$ and using the inequality $|\sup a-\sup b|\leq\sup|a-b|$ yields
$\left|\Phi(\bm{Z}_{1:m})-\Phi\left(\bm{Z}_{1:m}^{(j)}\right)\right|\leq\frac{1}{m}.$
Thus $\Phi$ satisfies the bounded differences condition with constants $c_{j}=1/m$ for all $j$. By McDiarmid’s inequality, for any $\delta\in(0,1)$, with probability at least $1-\delta$,
$\Phi(\bm{Z}_{1:m})\leq\mathbb{E}\Phi(\bm{Z}_{1:m})+\sqrt{\frac{\log(2/\delta)}{2}\sum_{j=1}^{m}c_{j}^{2}}=\mathbb{E}\Phi(\bm{Z}_{1:m})+\sqrt{\frac{\log(2/\delta)}{2m}}.$
Combining this with $\mathbb{E}\Phi(\bm{Z}_{1:m})\leq 2\mathfrak{R}_{m}(\mathcal{L})$ and Assumption 3.2 gives
$\Phi(\bm{Z}_{1:m})\leq\frac{2C_{\mathcal{F}}}{\sqrt{m}}+\sqrt{\frac{\log(2/\delta)}{2m}}$
with probability at least $1-\delta$. Finally, multiplying both sides by $M_{\varepsilon}$ and using $\ell_{f}=M_{\varepsilon}g_{f}$ yields (A.16). ∎
Applying (A.16) to the minority and majority empirical averages, we have with probability at least $1-\alpha/5$ for each,
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\ell_{f}(\bm{X}_{i},1)-\mathbb{E}_{\mathcal{P}_{1}}\ell_{f}(\bm{X},1)\right|$ $\leq M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{1}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{1}}}\right),$
$\sup_{f\in\mathcal{F}}\left|\frac{1}{n_{0}}\sum_{i=n_{1}+1}^{n}\ell_{f}(\bm{X}_{i},0)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)\right|$ $\leq M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{0}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{0}}}\right).$
Therefore,
$\sup_{f\in\mathcal{F}}|({\rm I})|$ $\leq\frac{n_{1}}{n+\tilde{n}_{1}}M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{1}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{1}}}\right)+\left|\frac{n_{1}+\tilde{n}_{1}}{n+\tilde{n}_{1}}-\frac{1}{2}\right|M_{\varepsilon},$
$\sup_{f\in\mathcal{F}}|({\rm II})|$ $\leq\frac{n_{0}}{n+\tilde{n}_{1}}M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{0}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{0}}}\right)+\left|\frac{n_{0}}{n+\tilde{n}_{1}}-\frac{1}{2}\right|M_{\varepsilon}.$
Recall that
$({\rm III})=\left(\frac{1}{n_{0c}}\sum_{i\in\mathcal{S}_{0c}}\ell_{f}(\bm{X}_{i},0)-\mathbb{E}_{\mathcal{P}_{0}}\ell_{f}(\bm{X},0)\right)-\left(\frac{1}{\tilde{n}_{0}}\sum_{i=1}^{\tilde{n}_{0}}\ell_{f}(\tilde{\bm{X}}_{i}^{(0)},0)-\mathbb{E}_{\tilde{\mathcal{P}}_{0}}\ell_{f}(\tilde{\bm{X}},0)\right).$
Applying (A.16) to both terms, with probability at least $1-2\alpha/5$,
$\sup_{f\in\mathcal{F}}|({\rm III})|\leq M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{0c}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{0c}}}\right)+M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{\tilde{n}_{0}}}+\sqrt{\frac{\log(10/\alpha)}{2\tilde{n}_{0}}}\right).$
Applying (A.16) to the synthetic minority samples, with probability at least $1-\alpha/5$,
$\sup_{f\in\mathcal{F}}|({\rm IV})|\leq M_{\varepsilon}\left(\frac{2C_{\mathcal{F}}}{\sqrt{\tilde{n}_{1}}}+\sqrt{\frac{\log(10/\alpha)}{2\tilde{n}_{1}}}\right).$
By the Kantorovich-Rubinstein duality and Assumption 3.4 (A), we have
$\sup_{f\in\mathcal{F}}|({\rm V})|=$ $\sup_{f\in\mathcal{F}}\bigg|\mathbb{E}_{\tilde{\mathcal{P}}_{1}}[\ell_{f}(\tilde{\bm{X}},1)]-\mathbb{E}_{(\tilde{\mathcal{P}}_{0})_{\#T}}[\ell_{f}(\tilde{\bm{X}},1)]\bigg|\leq L_{\ell}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{1},(\tilde{\mathcal{P}}_{0})_{\#T})\leq L_{\ell}\varepsilon_{T},$
$\sup_{f\in\mathcal{F}}|({\rm VI})|=$ $\sup_{f\in\mathcal{F}}\bigg|\mathbb{E}_{\mathcal{P}_{1}}[\ell_{f}(\bm{X},1)]-\mathbb{E}_{(\mathcal{P}_{0})_{\#T}}[\ell_{f}(\bm{X},1)]\bigg|\leq L_{\ell}\mathcal{W}_{1}(\mathcal{P}_{1},(\mathcal{P}_{0})_{\#T})\leq L_{\ell}\varepsilon_{T}.$
By Assumption 3.4 (A3), it holds uniformly over $f\in\mathcal{F}$ that
$\sup_{f\in\mathcal{F}}|({\rm VII})|=\sup_{f\in\mathcal{F}}\bigg|\mathbb{E}_{\tilde{\mathcal{P}}_{0}}h(\tilde{\bm{X}})-\mathbb{E}_{\mathcal{P}_{0}}h(\bm{X})\bigg|\leq 2\varepsilon_{h}.$
Thus $\sup_{f\in\mathcal{F}}|({\rm V})+({\rm VI})+({\rm VII})|\leq 2L_{\ell}\varepsilon_{T}+2\varepsilon_{h}=\varepsilon_{\rm BT}$.
By the triangle inequality,
$\sup_{f\in\mathcal{F}}|L^{\rm bc}(f)-L^{*}(f)|\leq$ $\sup_{f\in\mathcal{F}}|({\rm I})|+\sup_{f\in\mathcal{F}}|({\rm II})|$
$+\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\left(\sup_{f\in\mathcal{F}}|({\rm III})|+\sup_{f\in\mathcal{F}}|({\rm IV})|+\sup_{f\in\mathcal{F}}|({\rm V})+({\rm VI})+({\rm VII})|\right).$
Under the assumption that $\tilde{n}_{1}/(n_{0}-n_{1})\to 1$, we have
$\frac{n_{1}}{n+\tilde{n}_{1}}\to\frac{\pi_{1}}{2\pi_{0}},\quad\frac{1}{2}-\frac{\tilde{n}_{1}}{n+\tilde{n}_{1}}\to\frac{\pi_{1}}{2\pi_{0}},\quad\text{ and}\quad\frac{n_{0}}{n+\tilde{n}_{1}}\to\frac{1}{2}.$
Collecting the uniform bounds for $({\rm I})$ - $({\rm VII})$ and applying a union bound, with probability at least $1-\alpha$, we obtain
$\sup_{f\in\mathcal{F}}|L^{\rm bc}(f)-L^{*}(f)|\leq$ $M_{\varepsilon}\left\{\frac{\pi_{1}}{2\pi_{0}}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{1}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{1}}}\right)+\frac{1}{2}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{0}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{0}}}\right)\right$.
$\left.+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{0c}}}+\sqrt{\frac{\log(10/\alpha)}{2n_{0c}}}+\frac{2C_{\mathcal{F}}}{\sqrt{\tilde{n}_{0}}}+\sqrt{\frac{\log(10/\alpha)}{2\tilde{n}_{0}}}\right)\right$.
$\left.+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\left(\frac{2C_{\mathcal{F}}}{\sqrt{\tilde{n}_{1}}}+\sqrt{\frac{\log(10/\alpha)}{2\tilde{n}_{1}}}\right)\right\}+\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{\rm BT}.$
$\leq$ $\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\varepsilon_{\rm BT}+\frac{C}{2}\cdot(2C_{\mathcal{F}}+\sqrt{\log(10/\alpha)})\left\{\frac{1}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right\},$
where $C<\infty$ is a constant that depends on $M_{\varepsilon},\pi_{0},\pi_{1}$.
Finally, we have
$L^{*}(\widehat{f}^{\rm bc})-L^{*}(f^{*})=$ $\left(L^{*}(\widehat{f}^{\rm bc})-L^{\rm bc}(\widehat{f}^{\rm bc})\right)+\left(L^{\rm bc}(\widehat{f}^{\rm bc})-L^{\rm bc}(f^{*})\right)+\left(L^{\rm bc}(f^{*})-L^{*}(f^{*})\right)$
$\leq$ $\frac{\pi_{0}-\pi_{1}}{\pi_{0}}\varepsilon_{\rm BT}+C(2C_{\mathcal{F}}+\sqrt{\log(10/\alpha)})\left\{\frac{1}{\sqrt{n_{1}}}+\frac{1}{\sqrt{n_{0}}}+\frac{1}{\sqrt{n_{0c}}}+\frac{1}{\sqrt{\tilde{n}_{1}}}+\frac{1}{\sqrt{\tilde{n}_{0}}}\right\}$
with probability at least $1-\alpha$, where the inequality holds since $L^{\rm bc}(\widehat{f}^{\rm bc})-L^{\rm bc}(f^{*})\leq 0$. This completes the proof of Theorem 3.1. ∎
A.4 Proof of Theorem 3.2
Proof.
First, we derive the error bounds for $\|\widehat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}$. For the prediction function $f(\bm{x};\bm{\beta})=\sigma(\bm{x}^{\top}\bm{\beta})$ with logistic function $\sigma(t)=1/(1+\exp(-t))$, denote the loss function $\ell(\bm{x},y;\bm{\beta})=\ell_{f}(\bm{x},y)$ for convenience. For example, consider the cross-entropy loss function $\ell(\bm{x},y;\bm{\beta})=-y\log(\sigma(\bm{x}^{\top}\bm{\beta}))-(1-y)\log(1-\sigma(\bm{x}^{\top}\bm{\beta}))$.
Considering each source $k=1,\dots,K$, denote
$\psi_{k}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\bm{X}_{ki}\left(Y_{ki}-\sigma(\bm{X}_{ki}^{\top}\bm{\beta})\right).$
Then $\widehat{\bm{\beta}}_{k}$ satisfies $\psi_{k}(\widehat{\bm{\beta}}_{k})=\bm{0}_{d\times 1}$. Denote the population Hessian matrix as
$\bm{H}_{k}=-\mathbb{E}[\nabla_{\bm{\beta}_{k}}^{2}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta}_{k})]=\mathbb{E}[\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta}_{k})\bm{X}_{ki}\bm{X}_{ki}^{\top}].$
By Assumption 3.6 (A2), $\lambda_{\min}(\bm{H}_{k})\geq\underline{\kappa}_{k}$, where $\lambda_{\min}(\bm{H}_{k})$ denotes the smallest eigenvalue of $\bm{H}_{k}$. Taking the first-order Taylor expansion of $\psi_{k}(\widehat{\bm{\beta}}_{k})$, we have
$\bm{0}_{d\times 1}=\psi_{k}(\bm{\beta}_{k})+\bar{\bm{J}_{k}}(\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}),$
where $\bar{\bm{J}_{k}}=\int_{0}^{1}\bm{J}_{k}(\bm{\beta}_{k}+(\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k})t){\rm d}t$ for the Jacobian matrix $\bm{J}_{k}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\left\{\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta})\bm{X}_{ki}\bm{X}_{ki}^{\top}\right\}.$ Rearranging the above equation, we have
$\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}=-(\bar{\bm{J}_{k}})^{-1}\psi_{k}(\bm{\beta}_{k}).$ (A.17)
Next, in order to derive the error bound for $\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}$, we control $\|\bar{\bm{J}}_{k}^{-1}\|_{\rm op}$ and $\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}$. Denote $\Delta_{k}=\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}$ and recall that
$\Delta_{k}=-(\bar{\bm{J}}_{k})^{-1}\bm{\psi}_{k}(\bm{\beta}_{k}),\qquad\bar{\bm{J}}_{k}=\int_{0}^{1}\bm{J}_{k}(\bm{\beta}_{k}+t\Delta_{k}){\rm d}t,\qquad\bm{J}_{k}(\bm{\beta})=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta})\bm{X}_{ki}\bm{X}_{ki}^{\top}.$
By Assumption 3.6 (A2), $\lambda_{\min}(\bm{H}_{k})\geq\underline{\kappa}_{k}>0$ and $\lambda_{\max}(\bm{H}_{k})\leq\bar{\kappa}_{k}$. Moreover, by Assumption 3.6 (A1), $\|\bm{X}_{ki}\|_{2}\leq R$ almost surely.
Note that for the logistic link $\sup_{t\in\mathbb{R}}\sigma^{\prime}(t)\leq 1/4$ and $\sup_{t\in\mathbb{R}}|\sigma^{\prime\prime}(t)|<\infty$. By the mean value theorem, for any $\bm{\beta},\bm{\beta}^{\prime}$ and any $i$,
$|\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta})-\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta}^{\prime})|\leq\sup_{t\in\mathbb{R}}|\sigma^{\prime\prime}(t)|\cdot|\bm{X}_{ki}^{\top}(\bm{\beta}-\bm{\beta}^{\prime})|.$
Therefore, using Assumption 3.6 (A1),
$\|\bm{J}_{k}(\bm{\beta})-\bm{J}_{k}(\bm{\beta}^{\prime})\|_{\rm op}\leq\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}|\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta})-\sigma^{\prime}(\bm{X}_{ki}^{\top}\bm{\beta}^{\prime})|\cdot\|\bm{X}_{ki}\bm{X}_{ki}^{\top}\|_{\rm op}\leq C_{J}\|\bm{\beta}-\bm{\beta}^{\prime}\|_{2},$
where $C_{J}=\sup_{t\in\mathbb{R}}|\sigma^{\prime\prime}(t)|\cdot R^{3}$. In addition, each summand in $\bm{J}_{k}(\bm{\beta}_{k})$ has operator norm bounded by $(1/4)R^{2}$ and mean $\bm{H}_{k}$, hence a matrix concentration inequality implies that for any $\alpha\in(0,1)$,
$\|\bm{J}_{k}(\bm{\beta}_{k})-\bm{H}_{k}\|_{\rm op}\leq\delta_{J,k}(\alpha)$
with probability at least $1-\alpha/3$, where $\delta_{J,k}(\alpha)\to 0$ as $n_{k}\to\infty$. Let $r_{k}=\underline{\kappa}_{k}/(4C_{J})$ and consider the event
$E_{1,k}=\left\{\|\bm{J}_{k}(\bm{\beta}_{k})-\bm{H}_{k}\|_{\rm op}\leq\underline{\kappa}_{k}/4\right\}.$
On $E_{1,k}$, for any $\bm{\beta}$ satisfying $\|\bm{\beta}-\bm{\beta}_{k}\|_{2}\leq r_{k}$, we have
$\|\bm{J}_{k}(\bm{\beta})-\bm{H}_{k}\|_{\rm op}\leq\|\bm{J}_{k}(\bm{\beta})-\bm{J}_{k}(\bm{\beta}_{k})\|_{\rm op}+\|\bm{J}_{k}(\bm{\beta}_{k})-\bm{H}_{k}\|_{\rm op}\leq C_{J}r_{k}+\underline{\kappa}_{k}/4=\underline{\kappa}_{k}/2,$
which implies
$\lambda_{\min}(\bm{J}_{k}(\bm{\beta}))\geq\lambda_{\min}(\bm{H}_{k})-\|\bm{J}_{k}(\bm{\beta})-\bm{H}_{k}\|_{\rm op}\geq\underline{\kappa}_{k}/2.$
Define the event
$E_{2,k}=\left\{\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}\leq\underline{\kappa}_{k}r_{k}/4\right\}.$
For any $\bm{\beta}$ with $\|\bm{\beta}-\bm{\beta}_{k}\|_{2}=r_{k}$, by the mean-value expansion there exists $\tilde{\bm{\beta}}$ on the line segment between $\bm{\beta}_{k}$ and $\bm{\beta}$ such that
$\bm{\psi}_{k}(\bm{\beta})=\bm{\psi}_{k}(\bm{\beta}_{k})+\bm{J}_{k}(\tilde{\bm{\beta}})(\bm{\beta}-\bm{\beta}_{k}).$
Since $\|\tilde{\bm{\beta}}-\bm{\beta}_{k}\|_{2}\leq r_{k}$, on $E_{1,k}$ we have $\lambda_{\min}(\bm{J}_{k}(\tilde{\bm{\beta}}))\geq\underline{\kappa}_{k}/2$, and therefore
$\|\bm{\psi}_{k}(\bm{\beta})\|_{2}\geq\|\bm{J}_{k}(\tilde{\bm{\beta}})(\bm{\beta}-\bm{\beta}_{k})\|_{2}-\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}\geq(\underline{\kappa}_{k}/2)r_{k}-\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}.$
Hence on $E_{1,k}\cap E_{2,k}$,
$\|\bm{\psi}_{k}(\bm{\beta})\|_{2}\geq\underline{\kappa}_{k}r_{k}/4>0,$
which implies that the root of $\bm{\psi}_{k}(\bm{\beta})=\bm{0}$ cannot lie on or outside the boundary of the ball, and therefore
$\|\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}\|_{2}=\|\Delta_{k}\|_{2}\leq r_{k}.$
On $E_{1,k}\cap E_{2,k}$, for any $t\in[0,1]$ we have $\|\bm{\beta}_{k}+t\Delta_{k}-\bm{\beta}_{k}\|_{2}\leq\|\Delta_{k}\|_{2}\leq r_{k}$, hence by the uniform bound established above,
$\lambda_{\min}(\bm{J}_{k}(\bm{\beta}_{k}+t\Delta_{k}))\geq\underline{\kappa}_{k}/2.$
Consequently, for any $\bm{v}\in\mathbb{R}^{d}$,
$\bm{v}^{\top}\bar{\bm{J}}_{k}\bm{v}=\int_{0}^{1}\bm{v}^{\top}\bm{J}_{k}(\bm{\beta}_{k}+t\Delta_{k})\bm{v}\,{\rm d}t\geq\int_{0}^{1}(\underline{\kappa}_{k}/2)\|\bm{v}\|_{2}^{2}{\rm d}t=(\underline{\kappa}_{k}/2)\|\bm{v}\|_{2}^{2},$
and thus
$\lambda_{\min}(\bar{\bm{J}}_{k})\geq\underline{\kappa}_{k}/2,\qquad\|\bar{\bm{J}}_{k}^{-1}\|_{\rm op}\leq 2/\underline{\kappa}_{k}.$
Moreover, by $\|\bm{J}_{k}(\bm{\beta})\|_{\rm op}\leq(1/4)R^{2}$ we have
$\|\bar{\bm{J}}_{k}\|_{\rm op}\leq\int_{0}^{1}\|\bm{J}_{k}(\bm{\beta}_{k}+t\Delta_{k})\|_{\rm op}{\rm d}t\leq(1/4)R^{2}.$
By $\Delta_{k}=-(\bar{\bm{J}}_{k})^{-1}\bm{\psi}_{k}(\bm{\beta}_{k})$ and Step 3, on $E_{1,k}\cap E_{2,k}$,
$\|\Delta_{k}\|_{2}\leq\|\bar{\bm{J}}_{k}^{-1}\|_{\rm op}\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}\leq\frac{2}{\underline{\kappa}_{k}}\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}.$
In addition, since $\|\bar{\bm{J}}_{k}^{-1}\bm{u}\|_{2}\geq\|\bm{u}\|_{2}/\|\bar{\bm{J}}_{k}\|_{\rm op}$ for any $\bm{u}$, we obtain
$\|\Delta_{k}\|_{2}=\|\bar{\bm{J}}_{k}^{-1}\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}\geq\frac{1}{\|\bar{\bm{J}}_{k}\|_{\rm op}}\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}\geq\frac{4}{R^{2}}\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}.$
Note that $\mathbb{E}[\bm{\psi}_{k}(\bm{\beta}_{k})]=\bm{0}$ and $\mathbb{E}\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}^{2}=\mathrm{tr}(\bm{H}_{k})/n_{k}$. Indeed, writing $\bm{s}_{ki}=\bm{X}_{ki}\{Y_{ki}-\sigma(\bm{X}_{ki}^{\top}\bm{\beta}_{k})\}$, we have $\bm{\psi}_{k}(\bm{\beta}_{k})=n_{k}^{-1}\sum_{i=1}^{n_{k}}\bm{s}_{ki}$ with $\mathbb{E}[\bm{s}_{ki}]=\bm{0}$ and $\mathbb{E}[\bm{s}_{k1}\bm{s}_{k1}^{\top}]=\bm{H}_{k}$. By Assumption 3.6 (A1), $\|\bm{s}_{ki}\|_{2}\leq R$ almost surely, hence a vector Bernstein inequality implies that for any $\alpha\in(0,1)$,
$c_{\psi}(\alpha)\sqrt{\frac{\mathrm{tr}(\bm{H}_{k})}{n_{k}}}\leq\|\bm{\psi}_{k}(\bm{\beta}_{k})\|_{2}\leq\frac{1}{c_{\psi}(\alpha)}\sqrt{\frac{\mathrm{tr}(\bm{H}_{k})}{n_{k}}}$
with probability at least $1-\alpha/3$, where $c_{\psi}(\alpha)\in(0,1)$ depends only on $\alpha$ and $R$. Moreover, for $n_{k}$ sufficiently large such that
$\frac{1}{c_{\psi}(\alpha)}\sqrt{\frac{\mathrm{tr}(\bm{H}_{k})}{n_{k}}}\leq\underline{\kappa}_{k}r_{k}/4,$
the event $E_{2,k}$ holds with probability at least $1-\alpha/3$.
By a union bound, with probability at least $1-\alpha$, the events $E_{1,k}$ and $E_{2,k}$ hold together with the bounds in Step 5. On this event, combining Step 4 with Step 5 yields
$C_{1}\sqrt{\frac{\mathrm{tr}(\bm{H}_{k})}{n_{k}}}\leq\|\widehat{\bm{\beta}}_{k}^{\rm raw}-\bm{\beta}_{k}\|_{2}\leq\frac{C_{1}}{\lambda_{\min}(\bm{H}_{k})}\sqrt{\frac{\mathrm{tr}(\bm{H}_{k})}{n_{k}}},$
for some constant $C_{1}>0$ depending only on $R,\underline{\kappa}_{k},\bar{\kappa}_{k}$ and $\alpha$. This finishes the proof of the first statement.
Secondly, we derive the error bounds for $\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\widehat{\bm{\beta}}_{k}\|_{2}$. Let $w_{k}=\tilde{n}_{k}/(n_{k}+\tilde{n}_{k1})$. The mixed score is
$\psi_{k}^{\rm mix}(\bm{\beta})=(1-w_{k})\psi_{k}^{\rm raw}(\bm{\beta})+w_{k}\psi_{k}^{\rm syn}(\bm{\beta}),$
with $\psi_{k}^{\rm raw}(\bm{\beta})=n_{k}^{-1}\sum_{i=1}^{n_{k}}\bm{X}_{i}(Y_{i}-\sigma(\bm{X}_{i}^{\top}\bm{\beta}))$ and $\psi_{k}^{\rm syn}(\bm{\beta})=\tilde{n}_{k1}^{-1}\sum_{i=1}^{\tilde{n}_{k1}}\tilde{\bm{X}}_{i}(1-\sigma(\tilde{\bm{X}}_{i}^{\top}\bm{\beta}))$. At $\bm{\beta}_{k}$, $\mathbb{E}[\psi_{k}^{\rm raw}(\bm{\beta}_{k})]=0$ and $\mathbb{E}[\psi_{k}^{\rm syn}(\bm{\beta}_{k})]=\bm{\delta}_{k}$. A mean value expansion gives
$\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}=\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})^{-1}\big\{(1-w_{k})\psi_{k}^{\rm raw}(\bm{\beta}_{k})+w_{k}\psi_{k}^{\rm syn}(\bm{\beta}_{k})\big\},$
with $\bm{J}_{k}^{\rm mix}(\bm{\beta})=(1-w_{k})\frac{1}{n}\sum_{i=1}^{n_{k}}\sigma^{\prime}(\bm{X}_{i}^{\top}\bm{\beta})\bm{X}_{i}\bm{X}_{i}^{\top}+w_{k}\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\sigma^{\prime}(\tilde{\bm{X}}_{j}^{\top}\bm{\beta})\tilde{\bm{X}}_{i}\tilde{\bm{X}}_{i}^{\top}$. By the law of large numbers, $\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})\to\bm{H}_{k}^{\rm mix}:=(1-w_{k})\bm{H}_{k}+w_{k}\tilde{\bm{H}}_{k}$ in probability and $\|\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})^{-1}\|_{\rm op}=(1+o_{p}(1))/\lambda_{\min}(\bm{H}_{k}^{\rm mix})$ and $\|\bm{J}_{k}^{\rm mix}(\bar{\bm{\beta}}_{k})\|_{\rm op}\leq(1+o_{p}(1))\lambda_{\max}(\bm{H}_{k}^{\rm mix})$. Moreover, denote the synthetic score bias
$\bm{\delta}_{k}=\mathbb{E}_{\tilde{\mathcal{P}}_{k}}[\tilde{\bm{X}}(\tilde{Y}-\sigma(\tilde{\bm{X}}^{\top}\bm{\beta}_{k}))]-\mathbb{E}_{\mathcal{P}_{k}}[\bm{X}(Y-\sigma(\bm{X}^{\top}\bm{\beta}_{k}))]=\mathbb{E}_{\tilde{\mathcal{P}}_{k}}[\tilde{\bm{X}}(\tilde{Y}-\sigma(\tilde{\bm{X}}^{\top}\bm{\beta}_{k}))],$
where the last equality holds since $\mathbb{E}_{\mathcal{P}_{k}}[\bm{X}(Y-\sigma(\bm{X}^{\top}\bm{\beta}_{k}))]=0$. Since $d((\bm{x}_{1},y_{1}),(\bm{x}_{2},y_{2}))\leq\|\bm{x}_{1}-\bm{x}_{2}\|_{2}+c|y_{1}-y_{2}|$, $\nabla_{\bm{\beta}}\ell(\bm{X},Y,\bm{\beta})$ is Lipschitz continuous with constant $L_{k}=1+(B_{x}B_{\beta})/4+B_{x}/c$. Thus we have $\|\bm{\delta}_{k}\|_{2}\leq L_{k}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})$. Note that
$\|\psi_{k}^{\rm raw}(\bm{\beta}_{k})\|_{2}=O_{p}\left(\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}\right),\qquad\|\psi_{k}^{\rm syn}(\bm{\beta}_{k})-\bm{\delta}_{k}\|_{2}=O_{p}\left(\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}\right).$
Therefore
$\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\leq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})+O_{p}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{\rm mix})}\right),$
$\|\widehat{\bm{\beta}}_{k}^{\rm syn}-\bm{\beta}_{k}\|_{2}\geq\frac{w_{k}L_{k}}{\lambda_{\min}(\bm{H}_{k}^{\rm mix})}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k0},\mathcal{P}_{k0})-O_{p}\left(\frac{\sqrt{{\rm tr}(\bm{H}_{k})/n_{k}}+\sqrt{{\rm tr}(\tilde{\bm{H}}_{k})/\tilde{n}_{k1}}}{\lambda_{\min}(\bm{H}_{\rm mix})}\right).$
This finishes the proof of the second statement of this theorem.
Thirdly, we derive the error bounds for $\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}\|_{2}$. For any $\bm{\beta}\in\mathbb{R}^{d}$, let $\Delta L_{k}^{\rm bc}(\bm{\beta})=L_{k}^{\rm bc}(\bm{\beta})-L_{k}^{*}(\bm{\beta})$. By the optimality, we have
$\nabla_{\bm{\beta}}L_{k}^{\rm bc}(\widehat{\bm{\beta}}_{k}^{\rm bc})=\bm{0},\quad\nabla_{\bm{\beta}}L_{k}^{*}(\bm{\beta}_{k}^{*})=\bm{0}.$
Taking a mean value expansion, we have
$\bm{H}_{k}(\bar{\bm{\beta}}_{k})(\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*})=-\nabla_{\bm{\beta}}\left(\Delta L_{k}^{\rm bc}(\widehat{\bm{\beta}}_{k}^{\rm bc})\right)\quad\text{ with }\quad\lambda_{\min}(\bm{H}_{k}(\bar{\bm{\beta}}_{k}))\geq\underline{\kappa}_{k},$
for $\bar{\bm{\beta}}_{k}$ on the segment between $\widehat{\bm{\beta}}_{k}^{\rm bc}$ and $\bm{\beta}_{k}^{*}$. Thus we have
$\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*}=-\left(\bm{H}_{k}(\bar{\bm{\beta}}_{k})\right)^{-1}\cdot\nabla_{\bm{\beta}}\left(\Delta L_{k}^{\rm bc}(\widehat{\bm{\beta}}_{k}^{\rm bc})\right)$
and
$\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}\left\|\nabla_{\bm{\beta}}\left(\Delta L_{k}^{\rm bc}(\widehat{\bm{\beta}}_{k}^{\rm bc})\right)\right\|_{2}.$
Taking the derivative of $L_{k}^{\rm bc}$ and $L_{k}^{*}$ with respect to $\bm{\beta}$, we have
$\nabla_{\bm{\beta}}L_{k}^{\rm bc}(\bm{\beta})=\frac{1}{n_{k}+\tilde{n}_{k1}}\left[\sum_{i=1}^{n_{k}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},Y_{ki};\bm{\beta})+\tilde{n}_{k1}\cdot\left\{\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})+\nabla_{\bm{\beta}}\widehat{\Delta}_{k0}\right\}\right],$
where
$\nabla_{\bm{\beta}}\widehat{\Delta}_{k0}=\frac{1}{n_{k0,c}}\sum_{i\in\mathcal{S}_{k0,c}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},0;\bm{\beta})+\frac{1}{\tilde{n}_{k0}}\sum_{i=1}^{\tilde{n}_{k0}}\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}}_{ki}^{(0)},0;\bm{\beta}).$
Denote
$\nabla_{\bm{\beta}}L_{k}^{*}(\bm{\beta})=\frac{1}{2}\nabla_{\bm{\beta}}\mathbb{E}_{\tilde{\mathcal{P}}_{k1}}\ell(\bm{X},1;\bm{\beta})+\frac{1}{2}\nabla_{\bm{\beta}}\mathbb{E}_{\tilde{\mathcal{P}}_{k0}}\ell(\bm{X},0;\bm{\beta}).$
Using the same calculation in Section A.1, we have
$\nabla_{\bm{\beta}}\left\{L_{k}^{\rm bc}(\bm{\beta})-L_{k}^{*}(\bm{\beta})\right\}=({\rm i})+({\rm ii})+\frac{\tilde{n}_{k1}}{n_{k}+\tilde{n}_{k1}}(({\rm iii})+({\rm iv})+({\rm v})+({\rm vi})+(\rm vii)),$
where
$({\rm i})=$ $\frac{n_{k1}}{n_{k}+\tilde{n}_{k1}}\cdot\frac{1}{n_{k1}}\sum_{i=1}^{n_{k1}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},1;\bm{\beta})-\left(\frac{1}{2}-\frac{\tilde{n}_{k1}}{n_{k}+\tilde{n}_{k1}}\right)\cdot\mathbb{E}_{\mathcal{P}_{k1}}[\nabla_{\bm{\beta}}\ell(\bm{X},1;\bm{\beta})],$
$({\rm ii})=$ $\frac{n_{k0}}{n_{k}+\tilde{n}_{k1}}\cdot\frac{1}{n_{k1}}\sum_{i=n_{k1}+1}^{n_{k}}\nabla_{\bm{\beta}}\ell(\bm{X}_{ki},0;\bm{\beta})-\frac{1}{2}\cdot\mathbb{E}_{\mathcal{P}_{k0}}[\nabla_{\bm{\beta}}\ell(\bm{X},0;\bm{\beta})],$
$({\rm iii})=$ $\nabla_{\bm{\beta}}\left(\widehat{\Delta}_{k0}-\Delta_{k0}\right),$
$({\rm iv})=$ $\frac{1}{\tilde{n}_{k1}}\sum_{i=1}^{\tilde{n}_{k1}}\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}}_{ki}^{(1)},1;\bm{\beta})-\mathbb{E}_{\tilde{\mathcal{P}}_{k1}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})],$
$({\rm v})=$ $\mathbb{E}_{\tilde{\mathcal{P}}_{k1}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})]-\mathbb{E}_{(\tilde{\mathcal{P}}_{k0})_{\#T_{k}}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})],$
$({\rm vi})=$ $\mathbb{E}_{(\mathcal{P}_{k0})_{\#T_{k}}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})]-\mathbb{E}_{\mathcal{P}_{k1}}[\nabla_{\bm{\beta}}\ell(\tilde{\bm{X}},1;\bm{\beta})],$
$({\rm vii})=$ $\mathbb{E}_{\tilde{\mathcal{P}}_{k0}}[\nabla_{\bm{\beta}}h_{k}(\bm{X})]-\mathbb{E}_{\mathcal{P}_{k0}}[\nabla_{\bm{\beta}}h_{k}(\bm{X})]$
By applying Lemma 1, there exists a constant $C_{0k}>0$ such that, with probability at least $1-\alpha$,
$\|(\rm i)\|_{2}\leq$ $\frac{\pi_{1}}{2\pi_{0}}C_{0k}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{k1}}}+\sqrt{\frac{\log(10d/\alpha)}{n_{k1}}}\right),$
$\|(\rm ii)\|_{2}\leq$ $\frac{1}{2}C_{0k}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{k0}}}+\sqrt{\frac{\log(10d/\alpha)}{n_{k0}}}\right),$
$\|(\rm iii)\|_{2}\leq$ $C_{0k}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{k0,c}}}+\sqrt{\frac{\log(10d/\alpha)}{n_{k0,c}}}\right)+C_{0k}\left(\frac{2C_{\mathcal{F}}}{\sqrt{n_{k0}}}+\sqrt{\frac{\log(10d/\alpha)}{\tilde{n}_{k0}}}\right),$
$\|(\rm iv)\|_{2}\leq$ $C_{0k}\left(\frac{2C_{\mathcal{F}}}{\sqrt{\tilde{n}_{k1}}}+\sqrt{\frac{\log(10d/\alpha)}{\tilde{n}_{k1}}}\right).$
By the Kantorovich-Rubinstein duality, we have
$\|({\rm v})+({\rm vi})\|_{2}\leq L_{g}\mathcal{W}_{1}(\tilde{\mathcal{P}}_{k1},(\tilde{\mathcal{P}}_{k0})_{\#T_{k}})\leq 2L_{g}\varepsilon_{T}.$
Finally, by Assumption 3.4 (A3),
$\|(\rm vi)\|_{2}\leq 2\varepsilon_{h}.$
Therefore, with probability at least $1-\alpha$, we have
$\left\|\nabla_{\bm{\beta}}(L_{k}^{\rm bc}(\widehat{\bm{\beta}}_{k}^{\rm bc})-L_{k}^{*}(\widehat{\bm{\beta}}_{k}^{\rm bc}))\right\|_{2}\leq\frac{\pi_{0}-\pi_{1}}{2\pi_{0}}\underbrace{\left(2\varepsilon_{h}+2L_{g}\cdot\varepsilon_{T}\right)}_{\varepsilon_{{\rm BT},k}}$
$\quad+\underbrace{C_{3}(2C_{\mathcal{F}}+\log(10d/\alpha))\left\{\frac{1}{\sqrt{n_{k1}}}+\frac{1}{\sqrt{n_{k0}}}+\frac{1}{\sqrt{n_{k0,c}}}+\frac{1}{\sqrt{\tilde{n}_{k0}}}+\frac{1}{\sqrt{\tilde{n}_{k1}}}\right\}}_{\varepsilon_{{\rm sampling},k}},$
where $C_{3}$ depends on $C_{0k},\pi_{1}$ and $\pi_{0}$. Consequently, with probability at least $1-\alpha$,
$\|\widehat{\bm{\beta}}_{k}^{\rm bc}-\bm{\beta}_{k}^{*}\|_{2}\leq\frac{1}{\underline{\kappa}_{k}}(\varepsilon_{{\rm BT},k}+\varepsilon_{{\rm sampling},k}).$
This completes the proof of the third statement.
Let $\bm{M}=(\bm{\beta}_{1},\ldots,\bm{\beta}_{K})\in\mathbb{R}^{d\times K}$ and let $\bm{U}\in\mathbb{R}^{d\times r}$ be the matrix of the leading $r$ left singular vectors of $\bm{M}$. For any estimator $(\widehat{\bm{\beta}}_{k})_{k=1}^{K}$, define $\widehat{\bm{M}}=(\widehat{\bm{\beta}}_{1},\ldots,\widehat{\bm{\beta}}_{K})$ and let $\widehat{\bm{U}}$ be the matrix of the leading $r$ left singular vectors of $\widehat{\bm{M}}$. Note that the spectral gap $D=\sigma_{r}(\bm{M})-\sigma_{r+1}(\bm{M})>0$. By Davis–Kahan Theorem, we have
$\|\sin\Theta(\widehat{\bm{U}},\bm{U})\|_{F}\leq\frac{\|\widehat{\bm{M}}-\bm{M}\|_{F}}{\sigma_{r}(\bm{M})-\sigma_{r+1}(\bm{M})}=\frac{1}{D}\|\widehat{\bm{M}}-\bm{M}\|_{F}.$
Moreover,
$\|\widehat{\bm{M}}-\bm{M}\|_{F}^{2}=\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}\|_{2}^{2}.$
Therefore,
$\|\sin\Theta(\widehat{\bm{U}},\bm{U})\|_{F}\leq\frac{1}{D}\left(\sum_{k=1}^{K}\|\widehat{\bm{\beta}}_{k}-\bm{\beta}_{k}\|_{2}^{2}\right)^{1/2}.$
Applying this result to $\widehat{\bm{\beta}}_{k}^{\rm raw}$, $\widehat{\bm{\beta}}_{k}^{\rm syn}$ and $\widehat{\bm{\beta}}_{k}^{\rm bc}$ yields statements (iv)–(vi), respectively. This completes the proof of Theorem 3.2.
A.5 Proof of Theorem 3.3
Proof.
Denote
$\varphi(W)=\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})-\tau.$
Thus it suffices to show that $\mathbb{E}[\varphi(W)]=0.$ Decompose the error of the AIPW estimator as
$\widehat{\tau}^{\rm AIPW}-\tau=$ $\mathbb{P}_{n}[\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})]-\mathbb{P}[\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})]$
$=$ $(\mathbb{P}_{n}-\mathbb{P})[\varphi(W)]+\mathbb{P}\big(\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)$
$+(\mathbb{P}_{n}-\mathbb{P})[\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})].$
We first focus on the influence function fluctuation term $(\mathbb{P}_{n}-\mathbb{P})[\varphi(W)]$. By Assumption 3.7,
$|\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})-\tau|\leq|\mu_{1}^{*}(\bm{X})-\mu_{0}^{*}(\bm{X})|+\frac{|Y-\mu_{1}^{*}(\bm{X})|}{\eta}+\frac{|Y-\mu_{0}^{*}(\bm{X})|}{\eta}\leq C(\eta,M).$
By Hoeffding’s inequality, with probability at least $1-\alpha/2$,
$|(\mathbb{P}_{n}-\mathbb{P})[\varphi(W)]|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}.$ (A.18)
Next, we derive the upper bound for the population bias term $\mathbb{P}\big(\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)$. For simplicity, let $\delta_{e}=\widehat{e}-e^{*}$ and $\delta_{a}=\widehat{\mu}_{a}-\mu_{a}^{*}$ for $a\in\{0,1\}$. Note that
$\left\{\widehat{\mu}_{1}(\bm{X})-\frac{Z(Y-\widehat{\mu}_{1}(\bm{X}))}{\widehat{e}(\bm{X})}\right\}-\left\{\mu_{1}^{*}(\bm{X})-\frac{Z(Y-\mu_{1}^{*}(\bm{X}))}{e^{*}(\bm{X})}\right\}$
$=$ $\delta_{1}(\bm{X})\left(1-\frac{Z}{\widehat{e}(\bm{X})}\right)+Z\left(\frac{e^{*}(\bm{X})}{\widehat{e}(\bm{X})}-1\right)\cdot\frac{Y-\mu_{1}^{*}(\bm{X})}{e^{*}(\bm{X})}.$
Taking the expectation conditional on $\bm{X}$ on the right-hand side, by Assumption 3.7 (C1), we have
$\mathbb{E}\left[\left\{\widehat{\mu}_{1}(\bm{X})-\frac{Z(Y-\widehat{\mu}_{1}(\bm{X}))}{\widehat{e}(\bm{X})}\right\}-\left\{\mu_{1}^{*}(\bm{X})-\frac{Z(Y-\mu_{1}^{*}(\bm{X}))}{e^{*}(\bm{X})}\right\}\bigg|\bm{X}\right]=\delta_{1}(\bm{X})\left(1-\frac{e^{*}(\bm{X})}{\widehat{e}(\bm{X})}\right).$
We can use a similar way to derive that
$\mathbb{E}\left[\left\{\widehat{\mu}_{0}(\bm{X})-\frac{(1-Z)(Y-\widehat{\mu}_{0}(\bm{X}))}{1-\widehat{e}(\bm{X})}\right\}-\left\{\mu_{0}^{*}(\bm{X})-\frac{(1-Z)(Y-\mu_{0}^{*}(\bm{X}))}{1-e^{*}(\bm{X})}\right\}\bigg|\bm{X}\right]$
$=$ $\delta_{0}(\bm{X})\left(1-\frac{1-e^{*}(\bm{X})}{1-\widehat{e}(\bm{X})}\right).$
Expanding the population bias term, we have
$\mathbb{P}\big(\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)$
$=$ $\mathbb{E}\left\{\delta_{1}(\bm{X})\left(1-\frac{e^{*}(\bm{X})}{\widehat{e}(\bm{X})}\right)\right\}-\mathbb{E}\left\{\delta_{0}(\bm{X})\left(1-\frac{1-e^{*}(\bm{X})}{1-\widehat{e}(\bm{X})}\right)\right\}.$
Using the overlap condition that $\widehat{e},e^{*}\in[\eta,1-\eta]$ and the Cauchy-Schwarz inequality, we have
$|\mathbb{P}\big(\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})\big)|\leq$ $\frac{1}{\eta}\left(\mathbb{E}[|\delta_{1}(\bm{X})\delta_{e}(\bm{X})|]+\mathbb{E}[|\delta_{0}(\bm{X})\delta_{e}(\bm{X})|]\right)$
$\leq$ $\frac{C_{1}}{\eta}(r_{1}+r_{0})r_{e}.$ (A.19)
Finally, we consider the second-order empirical remainder $(\mathbb{P}_{n}-\mathbb{P})[\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})]$. Denote $\Delta\psi(\bm{W})=\psi(W;\widehat{\mu}_{1},\widehat{\mu}_{0},\widehat{e})-\psi(W;\mu_{1}^{*},\mu_{0}^{*},e^{*})$. By the Cauchy-Schwarz inequality, we have
$\|\Delta\psi\|_{L_{2}(\mathcal{P}_{X})}\leq C(\eta,M)[(r_{1}+r_{0})r_{e}+r_{1}r_{0}].$
Hoeffding’s inequality gives that with probability at least $1-\alpha/2$,
$|(\mathbb{P}_{n}-\mathbb{P})\Delta\psi(W)|\leq C_{0}\sqrt{\frac{\log(4/\alpha)}{2n}}[(r_{1}+r_{0})r_{e}+r_{1}r_{0}].$ (A.20)
Consequently, combining (A.18), (A.19) and (A.20), the proof of Theorem 3.3 is completed.
Appendix B Simulation Studies
In this section, we present simulation results under a range of experimental settings.
B.1 Mean Shift Model
The simulation investigates the performance of imbalanced classification on synthetic augmented data with and without bias correction. We consider an extremely imbalanced binary classification problem with sample size n=10,000n=10{,}000 and feature dimension $d=10$. For each replicate, labels $Y\in\{0,1\}$ are generated i.i.d. from $\mathrm{Bernoulli}(\pi_{1})$ with minority prevalence $\pi_{1}=0.005$. Conditional on $Y$, covariates are generated with independent coordinates and a class-dependent location shift: $\bm{X}\mid(Y=0)$ is centered at the origin, whereas $\bm{X}\mid(Y=1)$ is shifted by $\bm{\mu}$, where $\bm{\mu}=\mathbf{1}_{d}$ for the $t(2)$ and logistic settings, and $\mu=0.5\,\mathbf{1}_{d}$ for the normal setting. We examine three distributional settings for the feature noise: (i) heavy-tailed Student-$t$ with $2$ degrees of freedom, i.e., $\bm{X}=\bm{\xi}+\bm{\mu}\,I(Y=1)$ with elements $\xi_{j}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}t(2)$; (ii) Gaussian, with $\xi_{j}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathcal{N}(0,1)$; and (iii) logistic, with $\xi_{j}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Logistic}(0,1)$. Each replicate is split into training, validation and test sets with proportions $60\%/20\%/20\%$. We fit logistic regression models using gradient descent for $100$ epochs with learning rate $0.05$. For methods using augmentation, we generate $n_{s}=\lfloor(n_{0}-n_{1})\cdot\texttt{syn\_rate}\rfloor$ additional minority samples, where $n_{0}$ and $n_{1}$ are the majority/minority counts in the training set and $\texttt{syn\_rate}=2\pi_{1}(1-2\pi_{1})$. We use $k=15$ nearest neighbors for SMOTE. As a non-synthetic baseline for imbalance, we also consider sample-size reweighting via a weighted logistic loss that upweights minority observations.
Consider the loss function $\ell(\bm{x},y;\bm{\beta})=-y\cdot\bm{x}^{\top}\bm{\beta}+\log(1+\exp(\bm{x}^{\top}\bm{\beta}))$. For the three methods, using raw data, synthetic-augmented data, and synthetic-augmented data with bias correction, we train models by minimizing the respective loss functions given in Equations (2.1), (2.2), and (2.8) for 100 epochs.
Due to the imbalanced nature of the simulated data, overall accuracy is not an appropriate performance metric and can be misleading, particularly when class proportions are modified through upsampling. Therefore, we evaluate all methods using metrics that are more informative for minority-class performance, including recall, precision, and $F_{\beta}$-score with $\beta=1/\pi_{1}$. The $F_{\beta}$ score is defined as
$F_{\beta}=\frac{(1+\beta^{2}){\rm TP}}{(1+\beta^{2}){\rm TP}+{\rm FP}+\beta^{2}{\rm FN}},$
where ${\rm TP},{\rm FP}$ and $\rm FN$} represent the number of true positives, false positives and false negatives, respectively. In this case we choose $\beta=1/\pi_{1}$, to account for class imbalance, so that false negatives are penalized more heavily when the minority class is rare. Importantly, all upsampling procedures are applied exclusively to the training samples. Performance is evaluated on test sets that preserve the original class distribution, so that the reported results reflect performance under the true data-generating mechanism rather than the resampled training space.
Table 1: Performance metrics (recall, precision, F1-score) evaluated on raw and synthetic-augmented data, with/without bias correction (denoted as SMOTE and BC, respectively), across varying distributions based on the mean shift model. The results are based on 100 simulations and bold values indicate the top-performing method per metric.
$t(2)$ $\mathcal{N}(\cdot,1)$ $Logistic(\cdot,1)$
Recall Precision $F_{\beta}$ Score Recall Precision $F_{\beta}$ Score Recall Precision $F_{\beta}$ Score
Raw 0.5942 0.0059 0.5927 0.6404 0.0065 0.6388 0.6384 0.0065 0.6368
SMOTE 0.7240 0.0071 0.7221 0.7916 0.0080 0.7896 0.7927 0.0081 0.7907
Bootstrap 0.7163 0.0071 0.7145 0.7825 0.0079 0.7806 0.7931 0.0081 0.7912
Reweight 0.6703 0.0066 0.6686 0.7271 0.0073 0.7253 0.7196 0.0073 0.7178
SMOTE-BC 0.7317 0.0072 0.7298 0.7962 0.0080 0.7942 0.7972 0.0081 0.7952
The resulting evaluation metrics for the three methods are summarized in Table 1. Based on an analysis of 100 simulations across varying data distributions in Table 1, the application of the SMOTE synthetic data generator produced a marked performance improvement over using raw data alone, as measured by recall, precision, and $F_{\beta}$-score. Moreover, incorporating the bias correction technique on top of SMOTE leads to an additional improvement, indicating that the combined approach enhances model robustness and generalization. These results demonstrate that while SMOTE remains an effective foundation for handling data imbalance, integrating bias correction further refines the model’s predictive performance, yielding consistent gains across multiple evaluation metrics.
B.2 Non-linear Classification
To further evaluate the effectiveness and robustness of the proposed bias correction technique under diverse non-linear decision boundaries, we consider four controlled two-dimensional classification settings, each representing a distinct geometric relationship between the majority and minority classes as illustrated in Figure 3. In each replicate, we generate $n=1,000$ observations with $d=2$ features and labels $Y\in\{0,1\}$ sampled i.i.d. from $\mathrm{Bernoulli}(p_{1})$ with minority prevalence $\pi_{1}=0.05$. The four configurations vary in class overlap, non-convexity, and variance structure: one setting produces two elongated clouds with opposing correlation directions, yielding intersecting diagonal structures, another places the majority class on a noisy ring surrounding a minority strip, a third yields overlapping clusters with unequal spread and a location shift, and a fourth forms a clearly non-convex majority distribution with two lobes while the minority class occupies a horizontal band. For evaluation, we generate an independent test set of size $0.5n$ with balanced labels $\mathrm{Bernoulli}(0.5)$ to assess performance without class-prior effects. For each configuration, we generate synthetic samples using SMOTE with $k=5$ nearest neighbors and $n_{s}=\lfloor(n_{0}-n_{1})\cdot\texttt{syn\_rate}\rfloor$ (with $\texttt{syn\_rate}=0.5$), and then compare standard synthetic augmentation against the bias-corrected objective. We train the classifier using a two-hidden-layer ReLU network with 12 units per layer, optimized by SGD for 500 epochs (learning rate $0.05$ and weight decay $10^{-4}$). Performance is summarized by the $F_{\beta}$ score with $\beta=1/\pi_{1}$, averaged over 100 independent runs.
For each configuration, we compare standard synthetic augmentation with the bias-corrected objective using the same classifier. Performance is evaluated on an independent balanced test set using the $F_{\beta}$ score, with results averaged over $100$ replicates.
Figure 3: Distribution and $F_{\beta}$ score for four non-linear classification settings.
Figure 3 illustrates the data distributions and corresponding $F_{\beta}$ scores for four non-linear classification settings. Across all configurations, applying the bias correction technique consistently improves the performance of synthetic data augmentation, indicating its effectiveness in refining the representativeness of SMOTE-generated samples. Notably, in the right two cases, the support of the majority class distribution is clearly non-convex. Since SMOTE generates new samples through convex combinations of existing data points, this result is particularly striking—it demonstrates that the bias correction term can effectively enhance model performance even when the synthetic generator fails to capture the true underlying distribution. Furthermore, in most settings, the minority distributions exhibit large variance, a condition that typically poses a challenge in imbalanced classification. The improved performance under these high-variance conditions further highlights the robustness and adaptability of the bias correction approach in complementing synthetic oversampling methods.
B.3 Sigmoid Bernoulli Model
In this section, a simulation study on the sigmoid Bernoulli model is conducted to further evaluate the effectiveness and robustness of the proposed bias correction technique under a broad range of data-generating distributions and synthetic generators. Each replicate uses sample size n=1,000n=1{,}000 and dimension $d=10$. Covariates $\bm{X}\in\mathbb{R}^{d}$ are generated from one of several parametric families with a location vector $\mu$ and a common scale factor, and responses follow a sigmoid Bernoulli model
$\mathbb{P}(Y=1\mid\bm{X})=\sigma(\bm{X}^{\top}\bm{\beta}^{\star}),\qquad\bm{\beta}^{\star}=-\mathbf{1}_{d},$
where $\sigma(\cdot)$ denotes the logistic sigmoid function. We mainly study four exponential-family settings for elements of $\bm{X}$: Gaussian, Gumbel, location-scale $t$ (with $t_{4}$ noise), and HSD, and we additionally test two distributions outside the exponential family (Laplace and logistic) to assess robustness. For each distribution, $\bm{\mu}$ is chosen to induce a highly imbalanced label distribution (with a minority rate around $\pi_{1}=0.05$), and the data are split into training/validation/test sets with proportions $60\%/20\%/20\%$. To address imbalance, use SMOTE with $k=5$ nearest neighbors and compare it with the proposed bias correction method. The number of synthetic samples is set to $\tilde{n}=\lfloor(n_{0}-n_{1})\cdot\texttt{syn\_rate}\rfloor$ with $\texttt{syn\_rate}=10\pi_{1}(1-2\pi_{1})$. Performance is evaluated by the estimation error of $\bm{\beta}$, averaged over $100$ independent simulation runs.
Table 2: Estimation error of parameter $\bm{\beta}$ evaluated on raw and synthetic data (with/without bias correction) across varying distributions based on sigmoid Bernoulli model. The results are based on 100 simulations and bold values indicate the top-performing method.
Gaussian Gumbel Loc-Scale t HSD Laplace Logistic
Raw 2.441 2.426 2.431 2.496 2.401 2.384
SMOTE 2.310 2.311 2.347 2.380 2.330 2.362
Bias Corr 2.299 2.308 2.332 2.369 2.318 2.350
The evaluation of parameter estimation error for $\beta$ under the sigmoid Bernoulli model, shown in Table 2, reveals that employing SMOTE-generated synthetic data consistently enhances estimation accuracy across all tested data distributions when compared with models trained solely on raw data. Moreover, incorporating the bias correction technique yields an additional and notable reduction in estimation error, indicating that the correction effectively compensates for residual bias in the synthetic samples. This improvement highlights that, although SMOTE alone serves as a strong baseline for generating high-quality synthetic data, the bias correction step further refines the fidelity of parameter estimation, leading to more accurate recovery of the true underlying model. Overall, these results demonstrate that the proposed bias correction approach provides a meaningful and reliable performance gain beyond standard synthetic augmentation.
B.4 Average Treatment Effect Estimation
In this section, we conduct a simulation study to evaluate the performance of three methods for estimating the average treatment effect (ATE) under covariate imbalance and distributional heterogeneity. For each run, we generate $n=2000$ observations with $d=2$ covariates $\bm{X}$, and assign a binary treatment $Z$ according to a logistic propensity model
$\mathbb{P}(Z=1\mid\bm{X})=\sigma(\beta_{z,0}+\bm{\beta}_{z}^{\top}\bm{X}),$
with $\beta_{z,0}=-1.6$ and $\bm{\beta}_{z}=\mathbf{1}_{d}$, which induces an imbalanced treated proportion. Potential outcomes follow linear models,
$Y(1)=a_{1}+\bm{X}^{\top}\bm{\mu}_{1}+\varepsilon_{1},\qquad Y(0)=a_{0}+\bm{X}^{\top}\bm{\mu}_{0}+\varepsilon_{0},$
with $a_{1}=a_{0}=1$, $\bm{\mu}_{1}=(2,1)^{\top}$, $\bm{\mu}_{0}=\bm{\mu}_{1}$, and independent Gaussian noise with standard deviation $\sigma_{y}=1$. The observed outcome is Y=ZY(1)+(1Z)Y(0)Y=ZY(1)+(1-Z)Y(0), and the true ATE is $\tau=\mathbb{E}[Y(1)-Y(0)]$. We compare (i) the standard AIPW estimator computed on the raw data, (ii) an AIPW estimator that fits the propensity model after augmenting the treated group with SMOTE-generated synthetic covariates, and (iii) the same SMOTE augmentation combined with the proposed bias-corrected propensity fitting. The number of synthetic treated samples is chosen to target a treated share of $0.5$, and SMOTE uses $k=5$ nearest neighbors. Covariates $X$ are generated from four distributions: $t(4)$, $t(6)$, logistic, and Laplace. Performance is summarized by the RMSE of $\widehat{\tau}$ over $100$ independent runs, and the results are reported in Figure 4.
Figure 4: Square Root MSE for ATE Estimation of Three Methods across Four Distributions of Covariates $\bm{X}$.
The simulation results, presented in Figure 4, lead to two main conclusions. First, incorporating synthetic data, either through SMOTE or through our proposed bias correction method, improves ATE estimation relative to using raw data alone. Second, the bias correction procedure plays a crucial role, as it substantially reduces estimation error and consistently delivers superior performance across all examined distributional settings. These findings also indicate that the SMOTE generator introduces a non-negligible bias in this specific task, suggesting that it may not be universally effective across all applications.
Appendix C Synthetic Generators
We briefly review some synthetic generating methods in this section.
Reweighting and Bootstrap.
Reweighting is an intuitive oversampling technique used to address imbalanced data in machine learning. This approach works by assigning a higher weight to samples from the minority class so that the training process focuses more on learning from the underrepresented group. For instance, consider a dataset with $n_{1}$ minority samples and $n_{0}$ majority samples with $n_{1}\ll n_{0}$. A common reweighting approach assigns a weight of $w_{1}=\lfloor n_{0}/n_{1}\rfloor$ to each minority sample and a weight of $w_{0}=1$ to each majority sample (Breiman et al., 2017). This approach is equivalent to oversampling the minority class by replicating each minority sample $\lfloor n_{0}/n_{1}\rfloor-1$ times and training on the resulting augmented dataset with equal weight.
In contrast, bootstrap methods (Efron and Tibshirani, 1994) for imbalanced classification generate synthetic samples by randomly drawing with replacement from the minority samples. Bootstrap can be regarded as a generalization of the fixed-weight reweighting approach as it effectively assigns random weights to the minority samples in each resampling step. While both reweighting and bootstrap are intuitive and straightforward to implement, they are sensitive to outliers in the minority class. By heavily emphasizing or replicating the outliers, these approaches can potentially lead to overfitting to the noise present in the minority group.
Gaussian Mixture Model (GMM).
Gaussian mixture model (McLachlan and Peel, 2000) is an oversampling technique that assumes the minority samples follow a mixture of multivariate Gaussian distributions with unknown means and covariance matrices. This technique typically fits a single Gaussian component to the minority class. Given the minority samples $\bm{X}_{1},\dots,\bm{X}_{n_{1}}$, this approach first estimates the distributional parameters, including the empirical mean $\widehat{\bm{\mu}}_{1}=\frac{1}{n_{1}}\sum_{i=1}^{n_{1}}\bm{X}_{i}$ and the sample covariance matrix $\widehat{\bm{\Sigma}}_{1}=\frac{1}{n_{1}-1}\sum_{i=1}^{n_{1}}(\bm{X}_{i}-\widehat{\bm{\mu}}_{1})(\bm{X}_{i}-\widehat{\bm{\mu}}_{1})^{\top}$. Next, the synthetic samples are generated by randomly drawing from the estimated Gaussian distribution with mean $\widehat{\bm{\mu}}_{1}$ and covariance matrix $\widehat{\bm{\Sigma}}_{1}$. This approach effectively captures the first two moments of the minority distribution. However, the strong underlying Gaussian distribution assumption imposes significant constraints. It might generate poorly representative synthetic samples when the true minority distribution is not unimodal, particularly when the distribution is heavy tailed. For example, it involves a non-convex support or is highly skewed. In such cases, the synthetic data fails to accurately reflect the manifold of the minority class. Consequently, the introduction of the synthetic noise can potentially degrade the performance and robustness of the following training step.
Synthetic Minority Oversampling TEchnique (SMOTE).
SMOTE, introduced by Chawla et al. (2002), is a widely used oversampling method that generates synthetic minority samples in imbalanced datasets. SMOTE generates new synthetic samples by linearly interpolating between pairs of minority samples. It works as follows: for a randomly selected minority class sample, first find its $K$ nearest neighbors in the minority group. Then randomly select one of these $K$ nearest neighbors and create a new point along the line segment between the original point and the chosen neighbor. This procedure is repeated until the desired number of synthetic samples is reached. SMOTE requires a hyperparameter $K$, the number of nearest neighbors considered for each minority sample. Algorithm 2 provides a step-wise description on how SMOTE generates $\tilde{n}_{1}$ synthetic samples based on input data $\bm{X}_{1},\ldots,\bm{X}_{n_{1}}$.
Algorithm 2 Synthetic Minority Oversampling TEchnique (SMOTE)
0: Samples $(\bm{X}_{i})_{i=1}^{n_{1}}$, the number of nearest neighbors $K$, synthetic sample size $\tilde{n}_{1}$.
1:for each $i$ in $1:n_{1}$ do
2:  Find the $K$ nearest neighbors of $\bm{X}_{i}$, denoted as $\bm{X}_{i(1)},\ldots,\bm{X}_{i(K)}$.
3:end for
4:for each $i$ in $1:\tilde{n}_{1}$ do
5:  Sample index $t$ uniformly from $\{1,2,\dots,n_{1}\}$.
6:  Sample $U_{i}$ from $\mathcal{U}(0,1)$, i.e., from the uniform distribution on the interval [0, 1].
7:  Sample $k$ uniformly from $\{1,\ldots,K\}$.
8:  Generate the SMOTE sample $\tilde{\bm{X}}_{i}^{(1)}\leftarrow\bm{X}_{t}+U_{i}(\bm{X}_{t(k)}-\bm{X}_{t})$.
9:end for
9: Synthetic samples $(\tilde{\bm{X}}_{i}^{(1)})_{i=1}^{\tilde{n}_{1}}$.
Diffusion Model.
Diffusion models (Ho et al., 2020; Song et al., 2020) form one of the most popular classes of generative models for data synthesis. A diffusion model learns the distribution of observed samples by simulating and statistically revising a Markovian diffusion process that maps the data to standard Gaussian noise and then reconstructs data from noise. The framework consists of two phases: a fixed forward process, which maps a data example to Gaussian noise, and a learned backward process, which iteratively maps random noise back to a realistic data sample.
The forward process is a fixed Markov chain that progressively corrupts a sample with Gaussian noise over $T$ time steps, parameterized by a schedule of variance terms $\beta_{t}\in(0,1)$ for $t=1,\dots,T$. Starting with an original data sample $\bm{x}\in\mathbb{R}^{d}$ and letting $\bm{z}_{0}=\bm{x}$, a series of intermediate latent variables $\bm{z}_{1},\dots,\bm{z}_{T}\in\mathbb{R}^{d}$ are generated according to the following iterative equation,
$\bm{z}_{t}=$ $\>\sqrt{1-\beta_{t}}\cdot\bm{z}_{t-1}+\sqrt{\beta_{t}}\cdot\bm{\epsilon}_{t},\quad t=1,\dots,T,$
where $\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ is noise added at time $t$. Denoting $\alpha_{t}=\prod_{s=1}^{t}(1-\beta_{t})$ for $t=1,\dots,T$, this process allows for a direct-sampling property, which makes it possible to obtain $\bm{z}_{t}$ from $\bm{x}$ in one step:
$\bm{z}_{t}=\sqrt{\alpha_{t}}\cdot\bm{x}+\sqrt{1-\alpha_{t}}\cdot\bm{\epsilon},$
where $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$. Since $\beta_{t}<1$ is chosen such that $\alpha_{T}\approx 0$ for large $T$, the final latent variable $\bm{z}_{T}$ is guaranteed to be close to the standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}_{d})$.
The backward process defines the generative model. It defines a learned Markov chain that attempts to reverse the diffusion process, starting from pure noise $\bm{z}_{T}\sim\mathcal{N}(\bm{0},\bm{I}_{d})$ and iteratively denoising it back to a data sample $\bm{z}_{0}$:
$\bm{z}_{t-1}\mid(\bm{z}_{t},\phi_{t})\sim\mathcal{N}(\bm{f}_{t}(\bm{z}_{t},\phi_{t}),\sigma_{t}^{2}\bm{I}),\quad t=T,T-1,\dots,1.$
The function $\bm{f}_{t}(\bm{z}_{t},\phi_{t})$ is a neural network that is trained to estimate the mean of the approximate Gaussian distribution for the mapping from $\bm{z}_{t}$ to $\bm{z}_{t-1}$, and $\sigma_{t}$ is predetermined by the variance parameter $\beta_{t}$. By chaining these steps, diffusion models can synthesize high-fidelity data by gradually transforming Gaussian noise to structured samples.
Flow matching.
Flow matching (Lipman et al., 2022) aims to learn a smooth and invertible map from a simple base distribution, say, the standard Gaussian distribution, to the target data distribution. For observation $\bm{x}\in\mathbb{R}^{d}$, consider the probability density path $p:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}_{+}$ such that $\int p_{t}(\bm{x}){\rm d}\bm{x}=1$ for any $t\in[0,1]$. Let $p_{0}$ be the simple base distribution, and $p_{1}$ be the target data distribution. Define the flow $\bm{\phi}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ as a time-dependent differomorphic map satisfying that if $\bm{x}\sim p_{0}$, then $\bm{\phi}_{t}(\bm{x})\sim p_{t}$. Without loss of generality, let $\bm{\phi}_{0}(\bm{x})=\bm{x}$. The flow can be generated by a continuous normalizing flow vector field $\bm{v}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that satisfies $\frac{\rm d}{{\rm d}t}\bm{\phi}_{t}(\bm{x})=\bm{v}_{t}(\bm{\phi}_{t}(\bm{x}))$ (Chen et al., 2018). Flow matching simplifies the learning problem by utilizing conditional flows that define a straight path between the noise and a data point. For a given data point $\bm{x}_{1}\sim p_{1}$ and a noise sample $\bm{x}_{0}\sim p_{0}$, the optimal conditional vector field is the straight line path $\bm{u}_{t}(\bm{x})=\bm{x}_{1}-\bm{x}_{0}$.
The goal of flow matching is to train a neural network field $\bm{v}_{t}(\bm{x};\theta)$ parameterized by $\theta$ to match the ideal conditional vector field $\bm{u}_{t}$ in expectation. Suppose the target probability density path $p_{t}$ is generated by the vector field $\bm{u}_{t}$, flow matching aims to minimize the objective function
$L_{\rm FM}(\theta)=\mathbb{E}_{t,p_{0},p_{1}}\left[\|v_{t}(\bm{x}_{0}+t(\bm{x}_{1}-\bm{x}_{0});\theta)-(\bm{x}_{1}-\bm{x}_{0})\|^{2}\right],$
where $t\sim U(0,1)$, $\bm{x}_{0}\sim p_{0}$ and $\bm{x}_{1}\sim p_{1}$. With the learned vector field $\bm{v}_{t}(\bm{x};\theta)$ and a random noise sample $\bm{z}\sim p_{0}$, synthetic samples are generated by
$\tilde{\bm{x}}=\bm{\phi}_{1}(\bm{z}),\quad\text{where }\frac{\rm d}{{\rm d}t}\bm{\phi}_{t}(\bm{z})=\bm{v}_{t}(\bm{\phi}_{t}(\bm{z});\theta).$
There are many other synthetic generators, such as generative adversarial networks (GANs) (Goodfellow et al., 2014, 2020), normalizing flows (Rezende and Mohamed, 2015), and variational autoencoders (VAE) (Kingma and Welling, 2013). Please see Figueira and Vaz (2022); Lu et al. (2023) for a comprehensive survey.